### v1

# Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation

Abstract
Most existing methods for adapting models to out-of-distribution (OOD) domains rely on invariant representation learning to eliminate the influence of biased features. However, should bias always be eliminated—and if not, when should it be retained, and how can it be leveraged? To address these questions, we first present a theoretical analysis that explores the conditions under which biased features can be identified and effectively utilized. Building on this theoretical foundation, we introduce a novel framework that strategically leverages bias to complement invariant representations during inference. The framework comprises two key components that leverage bias in both direct and indirect ways: (1) using invariance as guidance to extract predictive ingredients from bias, and (2) exploiting identified bias to estimate the environmental condition and then use it to explore appropriate bias-aware predictors to alleviate environment gaps. We validate our approach through experiments on both synthetic datasets and standard domain generalization benchmarks. Results consistently demonstrate that our method outperforms existing approaches, underscoring its robustness and adaptability.
footnotetext: Correspondence to: Guangyi Chen <guangyichen1994@gmail.com>
1 Introduction
A widely used example of data bias is the “cow vs. camel classification” problem. As illustrated in Figure 1(a), models trained on imbalanced datasets often rely on the biased background features for prediction. Specifically, cows are typically photographed in grassy fields, while camels appear in desert settings. Consequently, the model may mistakenly learn to associate the background context, rather than the animal itself, with the class label. From a causal perspective, this issue arises due to spurious correlations, where the background acts as a confounding variable influencing both the images and the associated label, as shown in Figure 1(b).
Figure 1: Illustration of data bias in the “Cow vs. Camel Classification” problem. (a) An example of the Cow vs. Camel classification task, where both cows and camels are observed in distinct backgrounds, such as grasslands and deserts. (b) A causal graph illustrating spurious correlations introduced by confounders. (c) An intuitive example demonstrating how humans classify an image with ambiguous content as either a cow or a camel.
Most existing methods identify data bias as the key challenge in predicting unseen out-of-distribution (OOD) images and thus learn invariant representations to eliminate its influence. For instance, Invariant Risk Minimization (IRM) irm aims to learn a representation that supports a single classifier across all domains. Domain Adversarial Neural Networks (DANN) dann employ adversarial learning to encourage domain-invariant representations. Similarly, Self-Training lee2013pseudo improves generalization by generating pseudo-labels for target domain data based on learned invariant predictions. Some de-bias methods qi2020two leverage causal inference techniques, such as the back-door and front-door rules, to discover true causal relationships from spurious correlations. More discussions on related work can be found in Appendix A.
The core assumption of these methods is that biased features are uninformative, and only features invariant across different domains should be transferred. In this paper, we challenge this assumption and raise a critical question: Should bias always be eliminated? We provide an intuitive motivation for this question in part (c) of Figure 1. As humans, when classifying an image with ambiguous content as either a cow or a camel, we often rely on contextual background cues to make an educated guess. This intuitively suggests that, rather than being entirely detrimental, in some scenarios, the bias derived from background information can sometimes serve as a useful reference in prediction. Mathematically, it indicates that the domain-specific (biased) features and the prediction target are conditionally dependent given the invariant content, which conflicts with the assumption of uninformative bias in invariant representation learning.
Although bias can contain useful information, using it for prediction is challenging and potentially harmful. Recent studies chang2019domain ; wei2018general ; kong2023partial ; erm ; eastwood2024spuriosity have explored the use of biased features, with the most relevant being Stable Feature Boosting (SFB) eastwood2024spuriosity , which adapts conditionally independent biased features using pseudo-labels from stable ones at test time. However, these methods lack theoretical analysis of the conditions under which data bias can be reliably identified and effectively leveraged. Overlooking this can be risky, as bias often contains both “good” components (e.g., context that aids prediction) and “bad” components (e.g., spurious noise that harms prediction), making it essential to carefully disentangle and isolate the useful information.
To address this challenge, we present a theoretical analysis for the identification and utilization conditions of the data bias. Our results show that bias can indeed be leveraged, provided there is sufficient distributional variation across domains/labels and the relationship between the biased features and labels remains unblocked. Notably, even in more general scenarios where domain shifts affect the distributions of both the input data and the label, our analysis demonstrates that biased features can still offer valuable information for improving model performance in the inference stage.
Building on our theoretical analysis, we propose a principled generative framework to identify data bias and effectively leverage it for prediction. Specifically, we employ a Variational Autoencoder (VAE) to learn latent representations and disentangle content features from biased ones in an unsupervised manner. We then leverage the bias for prediction through both direct and indirect strategies. For the direct approach, we extract the beneficial components of the bias by aligning them with invariant predictions. It is achieved by taking the invariant predictions as pseudo-labels to guide the refinement of biased features that filter out harmful bias components, thereby retaining only the useful context for prediction. For the indirect approach, we utilize the biased feature to estimate the underlying environment context and adapt the prediction accordingly. In particular, we decompose the predictor into two components: an invariant predictor that takes content features as input, and a set of domain experts conditioned on biased features. Assuming that the target domain can be represented as a linear combination of source domains, we obtain a bias-aware predictor by the weighted mixture of experts using the inferred bias, enabling better alignment with the target distribution. The main contributions of this paper are summarized as follows:
  • We revisit the role of bias in OOD generalization. Unlike prior work that aims to eliminate bias, we highlight scenarios where bias can be a useful signal for improving prediction.
  • We provide a theoretical analysis to explore conditions under which data bias can be reliably identified and effectively leveraged.
  • We introduce a principled framework to leverage bias in both direct and indirect ways and validate our methods through experiments on both synthetic and real-world datasets.
2 Theoretical analysis
2.1 Problem setup and data generation Process
Figure 2: A graphical representation of the data generation process. The content variables ${\mathbf{c}}$ are invariant to environment changes, while the bias variables ${\mathbf{b}}$ vary across environments. Observed data ${\mathbf{x}}$ is generated by $g({\mathbf{c}},{\mathbf{b}})$, with ${\mathbf{c}}$ and ${\mathbf{b}}$ forming the latent variable ${\mathbf{z}}$. The environment $e$ affects ${\mathbf{b}}$ but not ${\mathbf{c}}$, and the target $y$ may also depend on $e$, reflecting complex data-environment interactions.
Notation and problem setup. We formulate the task as a multi-source domain generalization problem, aiming to learn a predictor that remains robust in an unseen target domain using only data from multiple source domains. Let ${\mathbf{x}}_{k}$ denote a high-dimension image observation ${\mathbf{x}}_{k}:=[x_{1},\cdots,x_{n_{x}}]\in{\mathcal{X}}\subset\mathbb{R}^{n_{x}}$, and $y$ denotes the corresponding label. With access to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$, with each domain represented as $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})=({\mathbf{x}}^{\mathcal{S}_{i}}_{k},y^{\mathcal{S}i}_{k})^{m_{i}}_{k=1}$, our goal is to learn a predictor $f({\mathbf{x}})$ that can generalize well to the new domain $\mathcal{T}$ where $\mathcal{T}\nsubseteq\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
Data generation Process. To illustrate the origin and impact of bias, we model the data generation process using a latent variable model. Let ${\mathbf{c}}:=[c_{1},\cdots,c_{n_{c}}]\in\mathcal{C}\subset\mathbb{R}^{n_{c}}$ denotes content variables and ${\mathbf{b}}:=[b_{1},\cdots,b_{n_{b}}]\in\mathcal{B}\subset\mathbb{R}^{n_{b}}$ denotes the bias. As shown in Figure 2, the generation process of the observation ${\mathbf{x}}$ is defined as ${\mathbf{x}}:=g({\mathbf{c}},{\mathbf{b}})$, where ${\mathbf{z}}:=[{\mathbf{c}},{\mathbf{b}}]$ with $n_{z}=n_{c}+n_{b}$ dimension denotes the latent variables and $g:[{\mathbf{c}},{\mathbf{b}}]\mapsto{\mathbf{x}}$ denotes the generating function. To distinguish ${\mathbf{c}}$ and ${\mathbf{b}}$, we introduce an environment variable $e$, where the content ${\mathbf{c}}$ remains invariant despite changes of the environment, whereas the bias ${\mathbf{b}}$ varies as the environment changes. Taking the “cow vs. camel classification” problem for illustration, ${\mathbf{c}}$ denotes the characteristics of the object, like the color/shape of cow; while ${\mathbf{b}}$ represents the background information. Note that we consider a more general case where $y$ may also be influenced by the environment, allowing for a more complex relationship modeling between the data and the environment. This assumption is reasonable, as the probability of encountering a cow versus a camel is strongly influenced by the environment (e.g., grasslands vs. deserts) in real-world scenarios. Here we highlight two causal paths between bias and label target, $y\rightarrow{\mathbf{b}}$ and $y\leftarrow e\rightarrow{\mathbf{b}}$, which motivate the direct and indirect ways to leverage the bias for prediction, respectively.
2.2 Theoretical conditions for identifiability and inference
Bias identification. We demonstrate that, with the advancement of representation learning, content and bias variables can be effectively disentangled with an identifiability guarantee. Since disentangling content and bias is fundamental for leveraging biased information, we briefly summarize the mild conditions under which such disentanglement can be achieved. It is important to note that this identifiability result has already been established in previous work and is not claimed as a contribution of this paper. However, despite the identification results, existing methods rely on invariant prediction (including invariant content variable and invariant relation between bias variable and label). Different from them, building on this identifiability framework, we go further by exploring how to effectively utilize biased data to enhance OOD performance.
Lemma 2.1.
(Block-wise Identification of ${\mathbf{c}}$ and ${\mathbf{b}}$ kong2023partial .) Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
  • A1 (Smooth and Positive Density): The probability density function of latent variables is smooth and positive, i.e. $P_{\mathbf{z}|\mathbf{e}}>0$.
  • A2 (Conditional Independence): Conditioned on $\mathbf{e}$, each $z_{i}$ is independent of any other $z_{j}$ for $i.j\in\{1,\cdots,n\},i\neq j$, i.e., $\log P_{\mathbf{z}|\mathbf{e},\mathbf{y}}=\sum_{i=1}^{n}q(z_{i},\mathbf{e},\mathbf{y})$, where $q_{i}(z_{i},\mathbf{e},\mathbf{y})$ is the log density of the conditional distribution, i.e., $q_{i}:\log P_{z_{i}|\mathbf{e},\mathbf{y}}$.
  • A3 (Linear Independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$, there exists $n_{b}+1$ values of $\mathbf{e}$, i.e., $e_{j}$ with $j=0,1,.\cdots,n_{b}$ such that these $n_{b}$ vectors $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})-\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{0})$ with $j=1,\cdots,n_{b}$ are linearly independent, where vector $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})$ is defined as follows:
    $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}})=\left(\frac{\partial q_{1}({\mathbf{b}}_{1},{\mathbf{e}})}{\partial{\mathbf{b}}_{1}},\cdots\frac{\partial q_{n_{b}}({\mathbf{b}}_{n_{b}},{\mathbf{e}})}{\partial{\mathbf{b}}_{n_{b}}}\right),$ (1)
  • A4 (Domain Variability:) There exist two values of $\mathbf{e}$, i.e., $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, s.t., for any set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability measure and cannot be expressed as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$, we have
[…]
Then the learned $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$ are block-wise identifiable.
The detailed proof can be found in Appendix B.1 . Such block-wise identification ensures that the learned content and bias representations, $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$, are functions of the ground truth content ${\mathbf{c}}$ and bias ${\mathbf{b}}$, respectively.
Use bias for inference. After identifying bias, we further provide theoretical insights into the conditions under which it can be leveraged to enhance predictive performance. Specifically, we show that the identified bias can contribute to prediction as long as there exists an active (i.e., unblocked) causal path from the bias to the prediction target, meaning the good information carried by the bias is not fully mediated or blocked by the content.
[…]
Lemma 2.2.
The bias ${\mathbf{b}}$ can be effectively utilized for prediction if Assumptions A1 through A5 hold.
As shown in Lemma 2.1, under Assumptions A1–A4, both content and bias information ${\mathbf{c}},{\mathbf{b}}$ can be well identified. It indicates that we can extract ${\mathbf{c}},{\mathbf{b}}$ containing all the information as the ground-truth variables. Assumption A5 states that
$P(y\mid{\mathbf{c}},{\mathbf{b}})\;\neq\;P(y\mid{\mathbf{c}})\quad\text{on a set of nonzero measure.}$.
In Bayesian decision theory (for any proper loss, e.g. log-loss or squared error), the Bayes-optimal predictor using $({\mathbf{c}},{\mathbf{b}})$ is
$f^{*}({\mathbf{c}},{\mathbf{b}})=\arg\min_{f}\;\mathbb{E}\bigl{[}\ell\bigl{(}y,f({\mathbf{c}},{\mathbf{b}})\bigr{)}\bigr{]},$
which depends on the full conditional $P(y\mid{\mathbf{c}},{\mathbf{b}})$. The best predictor using only ${\mathbf{c}}$ is
$\arg\min_{f}\;\mathbb{E}\bigl{[}\ell(y,f({\mathbf{c}}))\bigr{]},$
which depends only on $P(y\mid{\mathbf{c}})$. Since Assumption A5 ensures that the conditional distributions $p(y\mid{\mathbf{c}})$ and $p(y\mid{\mathbf{c}},{\mathbf{b}})$ differ, the joint predictor $f^{*}({\mathbf{c}},{\mathbf{b}})$ achieves a strictly lower expected loss compared to any predictor that relies solely on ${\mathbf{c}}$. This confirms that the bias variable ${\mathbf{b}}$ can indeed be exploited to improve predictive accuracy. A detailed proof, demonstrating that incorporating ${\mathbf{b}}$ leads to a better predictor under log-loss or squared error, is provided in Appendix B.2.
2.3 Understand bias’ contribution in prediction
Furthermore, we explore how bias contributes to the predictions. To illustrate the core insights of our approach, we begin with a binary classification setting. The extension to the multi-class scenario is in Appendix C. The core intuition is that $\mathbf{c}$ and $\mathbf{b}$ are conditionally independent given $y$, which indicates that we can decompose $p(y\mid\mathbf{c},\mathbf{b})$ into three parts, including a bias-aware predictor, an invariant predictor, and a label prior.
Theorem 2.3.
[Decomposition of Marginal Probability] Given the data generation process illustrated in Figure 2, we have:
$\ p(y=1\mid\mathbf{c},\mathbf{b})=\sigma\biggl{(}\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{b})\bigr{)}}_{\text{Bias-aware Predictor}}+\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c})\bigr{)}}_{\text{Invariant Predictor}}-\underbrace{\operatorname{logit}\bigl{(}p(y=1)\bigr{)}}_{\text{Label Prior}}\biggr{)},$
where $\sigma$ is the sigmoid function. The last two terms combined can be viewed as an invariant part of the model across all domains.
Proof sketch of Theorem 2.3
We begin by reformulating $p(y\mid\mathbf{c},\mathbf{b})$ to make its structure more explicit. Specifically, we have:
$\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c},\mathbf{b})\bigr{)}=\frac{p(\mathbf{c},\mathbf{b}\mid y=1)\,p(y=1)}{p(\mathbf{c},\mathbf{b}\mid y=0)\,p(y=0)}\quad\quad\text{consider }\mathbf{c}\,\perp\!\!\!\perp\,(\mathbf{b})\,\mid\,y:$
$\quad=\frac{p(\mathbf{b}\mid y=1)\,p(\mathbf{c}\mid y=1)\,p(y=1)}{p(\mathbf{b}\mid y=0)\,p(\mathbf{c}\mid y=0)\,p(y=0)}=\frac{p\bigl{(}y=1\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=1\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=1\bigr{)}}}{p\bigl{(}y=0\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=0\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=0\bigr{)}}}.$
A more detailed proof is provided in Appendix D.1.
Discussion on invariant prediction.
We observe that $p(\mathbf{c}\mid y)$ remains invariant across domains, since $y$ d-separate ${\mathbf{c}}$ and the environment variable $e$. Formally, $\forall\,e_{i},e_{j}\in E_{all}$, since $\mathbf{c}$ only depends on $y$, it follows that
$p(\mathbf{c}\mid y=1,\,e=e_{i})\;=\;p(\mathbf{c}\mid y=1,\,e=e_{j})\Rightarrow\frac{p(y=1,\,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}\;=\;\frac{p(y=1,\,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
which implies $\tfrac{p(y\mid\mathbf{c})}{p(y)}$ is an invariant quantity across all domains.
2.4 Leverage bias for better prediction
As shown in Theorem 2.3, we have established that the bias affects prediction solely through the bias-dependent component $p(y=1\mid\mathbf{b})$. Consequently, ensuring the reliability of this bias-aware prediction becomes crucial for effectively leveraging bias. However, relying on this term without additional assumptions poses challenges, as the bias-aware prediction component can vary significantly across environments. In the absence of proper constraints, incorporating it naively would revert the model to standard empirical risk minimization (ERM), failing to account for bias robustness. Furthermore, as shown in the causal graph in Figure 2, the term $p(y=1\mid\mathbf{b})$ can be decomposed into two components: a direct path $\mathbf{b}\rightarrow y$ and an indirect path $\mathbf{b}\leftarrow e\rightarrow y$. Motivated by this structure, we propose to leverage bias for prediction in both direct and indirect ways. The direct approach uses invariant features to guide the extraction of the predictive (i.e., positively correlated) component of the bias. In contrast, the indirect approach utilizes the bias to infer the environment $e$, which in turn informs the selection of an appropriate predictor tailored to that environment.
2.4.1 Using bias directly
The core idea of directly leveraging bias for the prediction is to extract its beneficial components, guided by the information encoded in invariant features.
Assumption 2.4 (Sufficient Information of ${\mathbf{c}}$).
[…]
Theorem 2.5 (Correction with Pseudo Labels).
Consider the predictor $f_{c}:{\mathbf{c}}\rightarrow y$. Define pseudo-label prediction as $\hat{y}=f_{c}({\mathbf{c}})$. Let $h_{0}\;=\;p(\hat{y}=0\mid y=0),\quad h_{1}\;=\;p(\hat{y}=1\mid y=1).$ Then it holds that
[…]
Justification about Theorem 2.5.
Equation 3 depends and only depends on the 2.4. The detailed proof can be found in Appendix D.2. Intuitively, if $c$ helps distinguish between $y=0$ and $y=1$, then a predictor based on $c$ should be corrected more often than chance, implying $h_{0}+h_{1}>1$.
Proof of Equation 4.
The key observation is that because $\hat{y}$ only depends on $y$ (through ${\mathbf{c}}$), we can relate the conditional distribution of the pseudo-label $\hat{y}$ given ${\mathbf{b}}$ to the distribution of the true label $y$. Formally,
$p(\hat{y}=1|{\mathbf{b}})=p(\hat{y}=1|y=1,{\mathbf{b}})p(y=1|{\mathbf{b}})+p(\hat{y}=1|y=0,{\mathbf{b}})p(y=0|{\mathbf{b}})$
$\textbf{Consider }\hat{y}\textbf{ only depends on }y,\hat{y}\perp\!\!\!\perp{\mathbf{b}}|y$
$=p(\hat{y}=1|y=1)p(y=1|{\mathbf{b}})+p(\hat{y}=1|y=0)p(y=0|{\mathbf{b}})$
$=h_{1}p(y=1|{\mathbf{b}})+(1-h_{0})(1-p(y=1|{\mathbf{b}}))$
Solving for $p(y=1\mid{\mathbf{b}})$ gives Equation 4.
2.4.2 Using bias indirectly
An alternative approach is to improve prediction by more accurately estimating the underlying environmental context. Let’s start from the objective, which can be formulated with the following optimization task:
$q^{*}(y\!\mid\!{\mathbf{b}})=\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}{\textstyle\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\bigl{[}D_{\mathrm{KL}}(p(y\!\mid\!{\mathbf{b}},e)\,\|\,q(y\!\mid\!{\mathbf{b}}))\bigr{]}}.$ (5)
This suggests identifying a predictor that minimizes the expected Kullback-Leibler (KL) divergence from $p(y\mid{\mathbf{b}},e)$ across all environments. The Bayes-optimal solution ohn2023optimal is
$q^{*}(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)$
This implies that we can leverage the identified bias to infer the environmental context and then develop a suitable bias-aware predictor $p(y\mid{\mathbf{b}})$ with a set of domain experts $p(y\mid{\mathbf{b}},e)$ to bridge the gaps between environments. The detailed derivation can be found in Appendix D.3.
Figure 3: Overall framework of the BAG method. The framework consists of three main modules: representation learning, predictor training, and test-time adaptation. In the representation learning stage, we employ a VAE model to disentangle content and bias variables. For prediction, we construct an invariant predictor, a label prior, and a bias-aware predictor that reweights multiple domain experts using a domain estimator. These three predictor components work together for the final prediction. During test-time adaptation, the bias-aware predictor is further refined by invariant guidance, using pseudo-labels generated by the invariant predictor, enhancing its adaptation to new domains.
3 Method
In this section, we build on our theoretical findings in  Section 2 to propose the Bias-Aware Generalization (BAG, in short) model which uses data bias for generalization. The whole pipeline can be found in Appendix E.1 . The multi-class version can be found in Appendix E.2 .
3.1 Overall framework
Representation learning and disentanglement.
Following Section 2.2, we first take the observed data ${\mathbf{x}}$ from the source domains as the inputs of the image encoder, which maps the image into the latent space, such as ${\mathbf{z}}=\text{Encoder}({\mathbf{x}})$. Then, we take ${\mathbf{z}}$ into a decode, $\text{Decoder}(\cdot)$, to generate the reconstruction as $\hat{{\mathbf{x}}}$.
We apply the VAE loss to train both encoder and decoder, where the objectives can be written as
$\mathcal{L}_{\text{vae}}\;=\;\mathbb{E}_{{\mathbf{x}}}\bigl{[}\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|^{2}\;+\;\beta\mathrm{KL}\bigl{(}q_{\phi}(\mathbf{z}\mid{\mathbf{x}})\,\|\,p(\mathbf{z})\bigr{)}\bigr{]},$ (6)
Where $\beta$ is a hyper-parameter to balance the reconstruction loss and the KL divergence, which is set to be 1 in our approach. The prior distribution $p({\mathbf{z}})$ is chosen as an isotropic Gaussian. Minimizing equation 6 encourages $(\mathbf{c},\mathbf{b})$ to be identifiable in a block-wise sense and captures both invariant and domain-specific factors for further modeling.
Following the method jiang2022invariant , we leverage a necessary but not sufficient condition for $\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y$, that is $\mathbb{E}\bigl{[}\mathbf{c}\cdot(\mathbf{b}-\mathbb{E}[\mathbf{b}\mid y])\bigr{]}=0$. Specifically, we add a regular term to constrain the condition independence, which can be written as:
$\mathcal{L}_{ind}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{c}_{i}\cdot\bigr{(}\mathbf{b}_{i}-\frac{1}{\bigl{|}\{j:y_{j}=y_{i}\}\bigr{|}}\sum_{j:y_{j}=y_{i}}\mathbf{b}_{j}\bigr{)}$
Predictors
Based on the learned content and bias variables, ${\mathbf{c}}$ and ${\mathbf{b}}$, we can make predictions from the latent space to the label target. Guided By Theorem 2.3, we implement the label predictor into three components:
$f({\mathbf{z}})\;=\;f_{b}(\mathbf{b})\;+\;f_{c}(\mathbf{c})\;-\;Pr,$
where $f_{b}$ is the bias-aware predictor (i.e. $p(y\mid{\mathbf{b}})$ ), whose input is the bias ${\mathbf{b}}$, $f_{c}$ is the invariant predictor to leverage the content variable ${\mathbf{c}}$. $Pr\in\mathbb{R}^{2}$ is a learnable bias term representing the logits prior $\operatorname{logit}(p(y))$. In particular, $f_{c}({\mathbf{c}})$ and $Pr$ together form an “invariant” module in all domains and can be used directly in the target domain without further modification. Since $\mathbf{b}$ depends on the domain $e$, using the $f_{s}$ only trained in the source domain naively at test time may introduce unwanted biases if the target domain differs significantly from the source domains.
3.2 Bias-Aware predictors
To exploit the bias $\mathbf{b}$ in the indirect manner, we introduce a set of $M$ learnable domain embeddings $\{e_{1},e_{2},\ldots,e_{M}\}$, which are corresponding to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$. Each embedding $e_{k}$ captures a distinct “mode” or style that may emerge in the source training data. Intuitively, this design follows the mixture-of-experts principle, allowing each expert to specialize in a subset of the training distribution.
In terms of model design, each $e_{k}$ is a learnable vector (or a small neural component) trained to represent a particular domain bias. We implement a domain classifier $p(e=e_{i}\mid{\mathbf{b}})$ by a softmax layer over linear outputs, to produce the scalar weights These weights indicate how well each expert $e_{i}$ aligns with the domain-specific bias $\mathbf{b}$. For each expert, we parametrize $p\bigl{(}y\mid\mathbf{b},e=e_{i}\bigr{)}$ as an MLP network, whose input is the combination of ${\mathbf{b}}$ and $e_{i}$. It models the expert specializing in domain $e_{i}$, making specific predictions based on the latent representation ${\mathbf{b}}$. Formally, we calculate the bias-aware predictor as a re-weighting process of different domain-specific predictors as
$f_{b}(\mathbf{b})\ =\sum_{i=0}^{M}p\bigl{(}y\mid\mathbf{b},e=e_{i}\bigr{)}p(e=e_{i}\mid{\mathbf{b}}).$
In training, we jointly learn all bias-aware predictions, the domain classifier, the invariant predictor, and the logits prior with a classification loss $\mathcal{L}_{\text{cls}}$.
3.3 Test time adaptation with invariance guidance
During test time, when accessing test data, we can further refine the model by exploiting the invariant component $\mathbf{c}$ as guided by Theorem 2.5. Concretely, when we receive test-domain inputs $\{{\mathbf{x}}^{\mathcal{T}}\}$, we first extract their latent representations $[\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}]$ using the pre-trained image encoder. We then leverage the learned invariant predictor $f_{c}(\mathbf{c}^{\mathcal{T}})$ to generate pseudo-labels $\hat{y}^{\mathcal{T}}$ on unlabeled test data.
In source domains, we can also have pseudo-labels, so we can collect $h_{0}$ and $h_{1}$ in source domains as shown in Theorem 2.5. It’s part of the probability confusion matrix between pseudo-labels and ground truth. Intuitively, these parameters adjust how the domain-invariant information can map the latent style factor $\mathbf{b}$ to valid probability estimates in the new domain.
According to Equation 4, recovering the corrected prediction $p(y=1\mid\mathbf{b})$ require knowing $p(\hat{y}=1\mid\mathbf{b})$ and the calibration terms $h_{0},h_{1}$. To estimate $p(\hat{y}=1\mid\mathbf{b})$, we update the bias-aware predictor $f_{b}$ with the test data $\mathbf{b}^{\mathcal{T}}$ and the generated pseudo-labels $\hat{y}^{\mathcal{T}}$. This process can be viewed as a lightweight test-time adaptation step:
$\min_{f_{b}}\;\mathbb{E}_{\mathbf{b}^{\mathcal{T}}}\Bigl{[}\mathcal{L}_{ada}=\ell\Bigl{(}\sigma\bigl{(}f_{b}(\mathbf{b}^{\mathcal{T}})\bigr{)},\;\hat{y}\Bigr{)}\Bigr{]},$ (7)
where the adaption loss $\mathcal{L}_{ada}$ is characterized as the cross-entropy loss with pseudo labels. We record this post-trained predictor as $\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}})$. Let $\phi(\cdot)=\operatorname{logit}((\frac{\sigma(\cdot)+h_{0}-1}{h_{0}+h_{1}-1}))$ to represent the test-time correction. We denote the corrected bias-aware predictor as $q(y|b^{\mathcal{T}})$. The final prediction is
$f([\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}])=f(\mathbf{c}^{\mathcal{T}})+\phi(\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}}))-Pr.$
Hence, the losses of the framework in both the training and test-time adaptation stages are as follows:
[…]
where $\lambda_{0}$, $\lambda_{1}$ denote the hyper-parameters.
4 Experiments
In this section, we test our BAG method on both synthetic data and real-world datasets that require adaptation in the target domain. To make a fair comparison, we follow the setting in jiang2022invariant .
Baseline methods
We compare our method with ERM and IRM irm , a classic approach that leverages invariant information for making predictions. Additionally, we compare with ACTIR jiang2022invariant , which disentangles the invariant feature and bias by an independent regularization. And with SFB eastwood2024spuriosity , which aims to exploit spurious features but in a weaker setting. We briefly summarized the difference between SFB and our method in the last part of the related work. SFB can serve as a special case of our method when 1) the environment doesn’t affect the label, and 2) don’t bias aware for prediction. For Office-Home dataset, since SFB doesn’t have results on it, we add MMD and GMDG as more baselines to show our performance.
4.1 Simulation experiments
Data simulation
To examine whether BAG can effectively identify and leverage domain shift information for better generalization, we devised a series of simulation experiments. Our data generation process strictly adheres to the causal structure depicted in Figure 2. First, we sample an environmental variable $e$ following a categorical distribution, $e\sim\mathrm{Categorical}(\{\pi_{1},\dots,\pi_{M}\})$, which subsequently induces variations across different domains.
Then, a binary label $y$ is generated in an environment-dependent manner: $y=\mathbf{1}\{\,w^{\mathsf{T}}E+b_{0}>0\}.$ Next, we assign stable content $c$ based on the value of $y$: if $y=0$, then $c=c_{0}+\epsilon$; if $y=1$, then $c=c_{1}+\epsilon$. Here, $\epsilon\sim\mathcal{N}(0,\sigma_{c}^{2}I)$. Simultaneously, the bias variable $b$ is determined as: $b=E+\mathbf{C}_{e,y},$ ensuring that the background shift is influenced jointly by the environment and the label. This setup emulates the common real-world scenario where false associations often arise. Finally, stable content and bias jointly generate the observed data: $x=M\begin{bmatrix}c\\ b\end{bmatrix}+\eta,$ where $M$ is an invertible matrix, and $\eta\sim\mathcal{N}(0,\sigma_{x}^{2}I)$ represents measurement noise.
Results and Discussions
This study evaluates several methods on a synthetic dataset of 5000 samples. To ensure result reliability, each method was repeated five times, and the average performance was recorded. As shown in the right side of Table 1, both ERM and IRM performed well on the training set but achieved test accuracies just above 50%. We attribute this to the significant distributional disparity of environmental information $e$ across domains in the dataset, which limits the effectiveness of the class label $y$ in supporting accurate predictions. In other words, an excessive focus on invariant information may hinder model optimization, leading to marginal performance gains. In contrast, our method leverages domain-specific information (${\mathbf{b}}$) more effectively, significantly enhancing prediction performance by reducing noise compared to the class label $y$.
Table 1: The comparison of the PACS dataset and the Synthetic dataset. All BAG results on PACS are obtained by averaging over 3 seeds. Baseline results are taken from eastwood2024spuriosity
Algorithm PACS Synthetic
P A C S Avg Acc
ERM 93.0$\pm$0.7 79.3$\pm$0.5 74.3$\pm$0.7 65.4$\pm$1.5 78.0 54.62$\pm$3.70
ERM + PL 93.7$\pm$0.4 79.6$\pm$1.5 74.1$\pm$1.2 63.1$\pm$3.1 77.6
IRM (irm, ) 93.3$\pm$0.3 78.7$\pm$0.7 75.4$\pm$1.5 65.6$\pm$2.5 78.3 54.73$\pm$2.72
IRM + PL 94.1$\pm$0.7 78.9$\pm$2.9 75.1$\pm$4.6 62.9$\pm$4.9 77.8
ACTIR (jiang2022invariant, ) 94.8$\pm$0.1 82.5$\pm$0.4 76.6$\pm$0.6 62.1$\pm$1.3 79.0 70.84$\pm$6.37
SFB eastwood2024spuriosity 95.8$\pm$0.6 80.4$\pm$1.3 76.6$\pm$0.6 71.8$\pm$2.0 81.2
BAG 96.6$\pm$0.8 86.0$\pm$0.4 77.9$\pm$1.0 73.0$\pm$0.4 83.4 97.48$\pm$1.82
4.2 Real-world experiments
Experimental Settings
We test our approach in commonly used PACS li2017deeper and Office-Homeoffice datasets. In the PACS dataset, we let one domain become the target and other domains as source domains. We use ResNet18 for PACS as the backbone and ResNet50 for Office-Home with MLP-based VAEs as the backbone and classifiers. We show our results over 3 random seeds. Our framework employs a single-layer and a two-layer linear layer as the encoder and decoder, respectively, and utilizes a single-layer linear layer for each specific classifier. We define three learnable parameters for the environment embedding. The experiment details can be found in Appendix F.
Table 2: Results on Office-Home with ResNet-50. BAG results are averaged over 3 seeds.
Alg. Office-Home Art Clipart Product Real Avg ERM 63.1 51.9 77.2 78.1 67.6 IRMirm 62.4 53.4 75.5 77.7 67.2 SFBeastwood2024spuriosity MMDli2018domain 62.4 53.6 75.8 76.4 67.1 GMDGtan2024rethinking 68.9 56.2 79.9 82.0 71.7 BAG 68.8 57.2 80.0 82.1 72.0
Results and Discussions
For the PACS data, in the left side of Table 1, we can find that our model is far superior to other models, with an average improvement of $2.2\%$ compared with best baseline, SFB method. Our model shows consistent improvement across all target domains. Among all target domains, we achieved better improvement on the Art Painting test, which may be because of the related bias in the domains. According to the experimental results of the Office-Home dataset on the Table 2, our BAG model outperforms all other baselines on three transfer tasks. It should be noted that we achieve much better performance in Clipart, which is the harder task; it shows BAG can achieve better generalization even when the domain gap is larger than the baselines. Considering that BAG involves two steps of optimization, we demonstrate the effects of each individual optimization strategy, namely BAG-BA, and BAG-TTA. Additionally, we explore the learning of latent variables by designing a model variant, BAG-VAE, which is a version of the model without the VAE component. We show that all the models performance better compared with baselines both in PACS and Synthetic data, but worse than BAG. The details of ablation study can be found in Appendix F.
5 Conclusion
In this paper, we challenge the conventional methods for OOD generation that treat bias solely as an obstacle. Instead, we theoretically demonstrate that bias can be strategically leveraged to enhance prediction, particularly when it retains useful dependencies on the target domain labels. Our framework demonstrates that the bias can be utilized in both direct and indirect ways. Through the empirical validation on synthetic and real-world benchmarks, we establish the effectiveness of this approach. The results consistently show that leveraging bias in a structured manner leads to more robust and adaptable models, offering a new perspective on OOD generation. Limitation: This work primarily aims to question the conventional treatment of bias and provide a theoretical foundation for understanding its effects. The experiments presented in this paper are intended to validate our proposed method, which don’t include large-scale experiments on extensive datasets. We acknowledge this limitation and leave systematic benchmarking on large datasets as an important direction for future work.
Appendix for
"Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation"
Appendix A Related Work
Domain generalization. Domain Generalization (DG) aims to develop models that generalize to unseen OOD target domains using only observed source domains for training. A common approach is learning invariant representations across all source domains, as seen in Invariant Risk Minimization (IRM) [1] and its extensions [16, 17, 18, 19, 20]. Other methods leverage adversarial learning [21, 22, 23, 24, 25] to extract domain-invariant features or use disentangled representation learning [26, 27, 11] to separate domain-invariant from domain-specific information. Causal inference techniques [28, 4], such as back-door and front-door rules, help mitigate bias. Furthermore techniques such as data augmentation [29, 30, 31, 32], feature augmentation [33, 34, 35], and network regularization [36, 37] can reduce overfitting and improve generalization.
Unsupervised domain adaptation Unlike DG, unsupervised domain adaptation utilizes unlabeled target data to guide model adaptation. Due to the observation of target distribution, many methods focus on distribution alignment by minimizing domain divergence [38, 39, 40, 41], using importance reweighting [42, 43, 44], and learning with adversarial discriminators, such as DANN [2], ADDA [45], and WDDA [46]. Beyond distribution alignment, some methods focus on invariant information-based adaptation, including self-training [3, 47, 48, 49, 50, 51], which uses pseudo-labels from a source-trained model to guide target domain learning, and batch normalization-based regularization [52, 53], which adapts only BN layers while keeping other parameters invariant. Recently, disentangled representation learning [54, 55] has shown strong potential for adaptation by separating domain-invariant content from domain-specific bias.
Test-Time domain adaptation In real-world scenarios, the target domain distribution is often unknown in advance, making it challenging to train an adaptive model beforehand, which motivates test time adaptation. The typical solution is Test-Time Training (TTT) [56, 57, 58, 59], which adapts models during inference by optimizing the same self-supervised learning objective used during pre-training. Besides, even without access to the target distribution, adaptation remains feasible with invariant information learned from the source domain, such as through self-training[60, 61, 62, 63] or network regularization [64, 65]. The most closely related work to ours is Stable Feature Boosting (SFB), which leverages unstable features alongside invariant representations to generate pseudo-labels when only the image is affected by environmental factors. Our approach differs in three key ways. First, we delve deeper into latent variable modeling to provide a more comprehensive explanation of bias. Second, our model accommodates more complex bias scenarios, allowing the environment variable to influence both the image and the label. Third, we integrate invariant correction and expert reweighting to effectively utilize biased data.
Appendix B Proof of identification and better prediction
B.1 Proof of the identification of latent variables
Lemma B.1.
(Block-wise Identification of ${\mathbf{c}}$ and ${\mathbf{b}}$ [7].) Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
  • A1 (Smooth and Positive Density): The probability density function of latent variables is smooth and positive, i.e. $P_{\mathbf{z}|\mathbf{e}}>0$.
  • A2 (Conditional Independence): Conditioned on $\mathbf{e}$, each $z_{i}$ is independent of any other $z_{j}$ for $i.j\in\{1,\cdots,n\},i\neq j$, i.e., $\log P_{\mathbf{z}|\mathbf{e},\mathbf{y}}=\sum_{i=1}^{n}q(z_{i},\mathbf{e},\mathbf{y})$, where $q_{i}(z_{i},\mathbf{e},\mathbf{y})$ is the log density of the conditional distribution, i.e., $q_{i}:\log P_{z_{i}|\mathbf{e},\mathbf{y}}$.
  • A3 (Linear Independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$, there exists $n_{b}+1$ values of $\mathbf{e}$, i.e., $e_{j}$ with $j=0,1,.\cdots,n_{b}$ such that these $n_{b}$ vectors $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})-\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{0})$ with $j=1,\cdots,n_{b}$ are linearly independent, where vector $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})$ is defined as follows:
    $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}})=\left(\frac{\partial q_{1}({\mathbf{b}}_{1},{\mathbf{e}})}{\partial{\mathbf{b}}_{1}},\cdots\frac{\partial q_{n_{b}}({\mathbf{b}}_{n_{b}},{\mathbf{e}})}{\partial{\mathbf{b}}_{n_{b}}}\right),$ (8)
  • A4 (Domain Variability:) There exist two values of $\mathbf{e}$, i.e., $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, s.t., for any set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability measure and cannot be expressed as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$, we have
[…]
Then the learned $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$ are block-wise identifiable.
Proof.
We begin with the matched marginal distribution $p_{\mathbf{x}|\mathbf{e},\mathbf{y}}$ to bridge the relation between $\mathbf{z}$ and $\hat{\mathbf{z}}$. Suppose that $\hat{g}:\mathbf{z}\rightarrow\mathcal{X}$ is an invertible estimated generating function, we have Equation (10).
$\forall\mathbf{e}\in\mathcal{(}E),\quad p_{\hat{\mathbf{x}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{x}|\mathbf{e},\mathbf{y}}\Longleftrightarrow p_{\hat{g}(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{g(\mathbf{z})|\mathbf{e},\mathbf{y}}.$ (10)
Sequentially, by using the change of variables formula, we can further obtain Equation (11)
$p_{\hat{g}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}=p_{g(\mathbf{z}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{g^{-1}\circ g(\hat{\mathbf{z}})|\mathbf{e}}|\mathbf{J}_{g^{-1}}|=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{g^{-1}}|\Longleftrightarrow p_{h(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}},$ (11)
where $h:=g^{-1}\circ g$ is the transformation between the ground-true and the estimated latent variables, respectively. $\mathbf{J}_{g^{-1}}$ denotes the absolute value of Jacobian matrix determinant of $g^{-1}$. Since we assume that $g$ and $\hat{g}$ are invertible, $|\mathbf{J}_{g^{-1}}|\neq 0$ and $h$ is also invertible.
According to A2 (conditional independent assumption), we can have Equation (12).
$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{z_{i}|\mathbf{e},\mathbf{y}}(z_{i}|\mathbf{e},\mathbf{y});\quad p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}(\hat{z}_{i}|\mathbf{e},\mathbf{y}).$ (12)
For convenience, we take logarithm on both sides of Equation (12) and further let $q_{i}:=\log p_{z_{i}|\mathbf{e},\mathbf{y}},\hat{q}_{i}:=\log p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}$. Hence we have:
$\log p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i},\mathbf{e},\mathbf{y});\quad\log p_{\hat{\mathbf{z},\mathbf{y}}|\mathbf{e}}=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}).$ (13)
By combining Equation (13) and Equation (11), we have:
$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}=p_{h(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{h^{-1}}|\Longleftrightarrow\sum_{i=1}^{n}q_{i}({z_{i},\mathbf{e}},\mathbf{y})+\log|\mathbf{J}_{h^{-1}}|=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}),$ (14)
where $\mathbf{J}_{h^{-1}}$ are the Jacobian matrix of $h^{-1}$.
Sequentially, we take the first-order derivative with $\hat{c}_{j}$ on Equation (14), where $j\in\{1,\cdots,n_{c}\}$, and have
$\sum_{i=1}^{n}\frac{\partial q_{i}(z_{i},\mathbf{u},\mathbf{y})}{\partial z_{i}}\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}+\frac{\partial\log|\mathbf{J}_{h^{-1}}|}{\partial\hat{c}_{j}}=\frac{\partial q_{j}(\hat{c}_{j},\mathbf{u},\mathbf{y})}{\partial\hat{c}_{j}}.$ (15)
Suppose $\mathbf{e}=e_{0},e_{1},\cdots,e_{n_{b}}$, we subtract the Equation (15) corresponding to $e_{k}$ with that corresponds to $e_{0}$, and we have:
$\sum_{i=1}^{n}\left(\frac{\partial q_{i}(z_{i},e_{k},\mathbf{y})}{\partial z_{i}}-\frac{\partial q_{i}(z_{i},e_{0},\mathbf{y})}{\partial z_{i}}\right)\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}=\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}.$ (16)
Since the distribution of estimated $\hat{\mathbf{c}}_{j}$ does not change across different domains, $\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}=0$. Since $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial z_{i}}$ does not change across different domains, $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial c_{i}}=\frac{\partial q_{i}(c_{i},e_{0},\mathbf{y})}{\partial c_{i}}$ for $i\in\{1,\cdots,n_{c}\}$. So we have
$\sum_{i=1}^{n_{b}}\left(\frac{\partial q_{i}(b_{i},e_{k},\mathbf{y})}{\partial b_{i}}-\frac{\partial q_{i}(b_{i},e_{0},\mathbf{y})}{\partial b_{i}}\right)\cdot\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0.$ (17)
Based on the linear independence assumption (A3), the linear system is a full-rank system. Therefore, the only solution is $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$.
Since $h(\cdot)$ is smooth over $\mathbf{z}$, its Jacobian can be formalized as follows
$\begin{gathered}\mathbf{J}_{h}=\begin{bmatrix}\begin{array}[]{c|c}\textbf{A}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{b}}}&\textbf{B}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{c}}}\\ \hline\cr\textbf{C}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{b}}}&\textbf{D}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{c}}}.\end{array}\end{bmatrix}\end{gathered}$ (18)
Note that $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$ means that $\textbf{B}=0$. Since $h(\cdot)$ is invertible, $\mathbf{J}_{h}$ is a full-rank matrix. Therefore, for each $b_{i},i\in\{1,\cdots,n_{b}\}$, there exists a $h_{i}$ such that $b_{i}=h_{i}(\hat{\mathbf{b}})$.
Besides, based on A4, one can show that all entries in the submatrix C zero according to part of the proof of Theorem 4.2 in [7](Steps 1, 2, and 3). Therefore, $\mathbf{b}$ and $\mathbf{c}$ are block-wise identifiable. ∎
B.2 Better predictor with adding bias
Proof.
Under Assumptions A1–A5, we compare the Bayes-optimal risks for predictors based on $c$ versus $(c,b)$.
Log-loss.
The Bayes-optimal log-loss risks satisfy
$R_{\log}\bigl{(}q^{*}(\cdot\mid c)\bigr{)}=H(Y\mid C),\qquad R_{\log}\bigl{(}q^{*}(\cdot\mid c,b)\bigr{)}=H(Y\mid C,B),$
where
$H(Y\mid C)=-\mathbb{E}_{C}\bigl{[}\mathbb{E}_{Y\mid C}[\log p(Y\mid C)]\bigr{]},\quad H(Y\mid C,B)=-\mathbb{E}_{C,B}\bigl{[}\mathbb{E}_{Y\mid C,B}[\log p(Y\mid C,B)]\bigr{]}.$
By A5 there is a positive-measure set of $(c,b)$ on which $p(y\mid c,b)\neq p(y\mid c)$. Hence
$I(Y;B\mid C)=H(Y\mid C)-H(Y\mid C,B)=\mathbb{E}_{C,B}\!\bigl{[}D_{\mathrm{KL}}(p(y\mid c,b)\,\|\,p(y\mid c))\bigr{]}>0,$
so $H(Y\mid C,B)<H(Y\mid C)$ and therefore
$R_{\log}\bigl{(}q^{*}(\cdot\mid c,b)\bigr{)}<R_{\log}\bigl{(}q^{*}(\cdot\mid c)\bigr{)}.$
Squared-error loss.
The Bayes-optimal squared-error risks satisfy
$R_{\mathrm{sq}}\bigl{(}f^{*}(c)\bigr{)}=\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C)\bigr{]},\qquad R_{\mathrm{sq}}\bigl{(}f^{*}(c,b)\bigr{)}=\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}.$
By the law of total variance,
$\mathrm{Var}(Y\mid C)=\mathbb{E}_{B}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}\;+\;\mathrm{Var}_{B}\!\bigl{(}\mathbb{E}[Y\mid C,B]\bigm{|}C\bigr{)}.$
Since A5 ensures that $B$ influences the conditional mean of $Y$ given $C$, the second term is strictly positive on a set of positive probability. Thus
$\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}<\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C)\bigr{]},$
and consequently
$R_{\mathrm{sq}}\bigl{(}f^{*}(c,b)\bigr{)}<R_{\mathrm{sq}}\bigl{(}f^{*}(c)\bigr{)}.$
This completes the proof that incorporating $b$ strictly reduces both log-loss and squared-error Bayes risks. ∎
Appendix C Multi-class Method
Suppose we have $K\geq 2$ classes. We "one-hot encode" these classes, so that $y$ takes values in the set $\mathcal{Y}=\{(1,0,...,0),(0,1,0,...,0),...,(0,...,0,1)\}\subseteq\{0,1\}^{K}.$ In the training stage, let $\hat{y}=f_{\mathbf{c}}(\mathbf{c})$ $\epsilon\in[0,1]^{\mathcal{Y}\times\mathcal{Y}}$ with
$\epsilon_{y_{i},y_{j}}=p[\hat{y}=y_{i}\mid y=y_{j}]$
denote the class-conditional confusion matrix of the pseudo-labels. Then, we have
$\frac{p(y=y_{i}\mid\mathbf{c},\mathbf{b})}{p(y\neq y_{i}\mid\mathbf{c},\mathbf{b})}$ (19)
$=\frac{p(\mathbf{c},\mathbf{b},y=y_{i})}{\sum_{y_{j}\neq y_{i}}p(\mathbf{c},\mathbf{b},y=y_{j})}$
$=\frac{p(\mathbf{c},\mathbf{b}\mid y=y_{i})p(y=y_{i})}{\sum_{y_{j}\neq y_{i}}p(\mathbf{c},\mathbf{b}\mid y=y_{j})p(y=y_{j})}$
$\quad\text{consider }\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y:$
$=\frac{p(\mathbf{c}\mid y=y_{i})p(\mathbf{b}\mid y=y_{i})p(y=y_{i})}{\sum_{y_{j}\neq y_{i}}p(\mathbf{c}\mid y=y_{j})p(\mathbf{b}\mid y=y_{j})p(y=y_{j})}$
$=\frac{p(y=y_{i}\mid\mathbf{b})p(\mathbf{c}\mid y=y_{i})}{\sum_{y_{j}\neq y_{i}}p(y=y_{j}\mid\mathbf{b})p(\mathbf{c}\mid y=y_{j})}$
$=\frac{p(y=y_{i}\mid\mathbf{b})\frac{p(y=y_{i}\mid\mathbf{c})}{p(y=y_{i})}}{\sum_{y_{j}\neq y_{i}}p(y=y_{j}\mid\mathbf{b})\frac{p(y=y_{j}\mid\mathbf{c})}{p(y=y_{j})}}$
for $P\in R^{\mathcal{Y}}$ defined by
$P_{y_{i}}=\frac{p(y=y_{i}\mid\mathbf{b})p(y=y_{i}\mid\mathbf{c})}{p(y=y_{i})}\quad\text{ for each }y_{i}\in\mathcal{Y}.$
By using the logit method, we obtain:
$\operatorname{logit}(p(y=y_{i}\mid\mathbf{c},\mathbf{b}))$ $=\log\left(\frac{P_{y_{i}}}{\sum_{y_{j}\neq y_{i}}P_{y_{j}}}\right)$ (20)
$=\operatorname{logit}\left(\frac{P_{y_{i}}}{\|P\|_{1}}\right)$
Applying the sigmoid function to each side, we have
$p(y=y_{i}\mid\mathbf{c},\mathbf{b})=\frac{P}{\|P\|_{1}}.$
Like binary classification, we only need to estimate $p(y=y_{i}\mid\mathbf{b},e)$ since we assume that $p(y=y_{i}\mid\mathbf{c})$ and $p(y=y_{i})$ can be directly used.
$E[\hat{y}\mid\mathbf{b}]$ (21)
$=\sum_{y_{i}\in\mathcal{Y}}E[\hat{y}\mid y=y_{i},\mathbf{b}]p[y=y_{i}\mid\mathbf{b}]$
$\hat{y}\textbf{ only depends on }y$
$=\sum_{y_{i}\in\mathcal{Y}}E[\hat{y}\mid y=y_{i}]p[y=y_{i}\mid\mathbf{b}]$
$=\epsilon E[y\mid\mathbf{b}]$
When $\epsilon$ is non-singular, this has the unique solution $E[y\mid\mathbf{b}]=\epsilon^{-1}E[\hat{y}\mid\mathbf{b}]$, giving a multiclass equivalent. However, it is numerically more stable to estimate $E[y\mid\mathbf{b}]$ by the least-squares solution
$\operatorname{argmin}_{p\in\Delta^{\mathcal{Y}}}\left\|\epsilon p-E[\hat{y}\mid\mathbf{b}]\right\|_{2}.$
Appendix D Proof of section 2
D.1 Decomposition and invariant
Assume by causal representation learning, we have obtained $\mathbf{c}$ (stable latent variables) and $\mathbf{b}$ (unstable latent variables). Our goal is to estimate $p(y\mid\mathbf{c},\mathbf{b})$.
$\frac{p(y=1\mid\mathbf{c},\mathbf{b})}{p(y=0\mid\mathbf{c},\mathbf{b})}$ (22)
$=\frac{p(\mathbf{c},\mathbf{b},y=1)}{p(\mathbf{c},\mathbf{b},y=0)}$
$=\frac{p(\mathbf{b}\mid y=1)p(\mathbf{c}\mid y=1)p(y=1)}{p(\mathbf{b}\mid y=0)p(\mathbf{c}\mid y=0)p(y=0)}$
$=\frac{p(y=1\mid\mathbf{b})p(\mathbf{c}\mid y=1)}{p(y=0\mid\mathbf{b})p(\mathbf{c}\mid y=0)}$
$=\frac{p(y=1\mid\mathbf{b})\frac{p(y=1\mid\mathbf{c})}{p(y=1)}}{p(y=0\mid\mathbf{b})\frac{p(y=0\mid\mathbf{c})}{p(y=0)}}$
Assume we have $e_{all}=\{e_{0},e_{1},...,e_{n}\}$, representing different domains. We assume the source domains are $e_{s}=\{e_{0},e_{1},e_{s}\}$, and target domains are $e_{t}=\{e_{s+1},...,e_{n}\}$.
$\forall e_{i},e_{j}\in e_{all}$ (23)
[…]
${p(\mathbf{c}\mid y=1,e=e_{i})}=p(\mathbf{c}\mid y=1,e=e_{j})$
   Use Bayes rule
$\frac{p(y=1,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}=\frac{p(y=1,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
$p(y=1\mid\mathbf{c})=\sum_{e\in e_{s}}{\frac{p(y=1,e=e\mid\mathbf{c})}{p(y=1,e=e)}p(e\mid e_{s})}$
This result allows us to estimate $p(y=1\mid\mathbf{c})$ using data from source domains and apply it in the target domain.
D.2 Proof of theorem 2.5
In this section, we would like to prove $h_{0}+h_{1}=1$ if and only if $\mathbf{c}\perp\!\!\!\perp y$.
Forward Implication: Suppose $h_{0}+h_{1}=1$.
$h_{0}+h_{1}=1\quad\text{if and only if}\quad\mathbf{c}\perp y.$
Note that the entire result also holds, with almost identical proof, in the multi-environment setting of Sections 3 and 5, conditioned on a particular environment $E$.
Proof. We first prove the forward implication.
Suppose $h_{0}+h_{1}=1$.
If $\Pr[y=1]\in\{0,1\}$, then $\mathbf{c}$ and $y$ are trivially independent, because the value of $y$ is constant. In this case, the result holds trivially. So, we assume $\Pr[y=1]\in(0,1)$.
We now compute the expected value of the pseudo-label $\hat{y}$ using the Law of Total Expectation:
$E[\hat{y}]=\Pr[\hat{y}=1\mid y=1]\cdot\Pr[y=1]+\Pr[\hat{y}=0\mid y=0]\cdot\Pr[y=0].$
That is,
$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p),$
where $p=\Pr[y=1]$.
Next, we simplify this expression:
$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p)=(h_{0}+h_{1}-1)\cdot p+1-h_{0}.$
Using the assumption $h_{0}+h_{1}=1$, we get:
$E[\hat{y}]=1-h_{0}.$
Thus, $E[\hat{y}]=1-h_{0}$.
Now, since the expected value of $\hat{y}$ is independent of the actual value of $y$, this suggests that $\hat{y}$ provides no information about $y$, implying that $\hat{y}$ and $y$ are independent.
Since $\hat{y}$ is a deterministic function of $\mathbf{c}$, this implies that $\mathbf{c}$ and $y$ must also be independent. Therefore, we have shown that if $h_{0}+h_{1}=1$, then $\mathbf{c}\perp y$.
To prove the reverse implication:
Suppose $\mathbf{c}\perp y$.
If $\mathbf{c}$ and $y$ are independent, then the pseudo-label $\hat{y}$ is also independent of $y$. Hence, we have:
$\Pr[\hat{y}=1\mid y=1]=\Pr[\hat{y}=1\mid y=0].$
This implies that:
$h_{1}=1-h_{0}.$
Therefore, we have:
$h_{0}+h_{1}=h_{0}+(1-h_{0})=1.$
Thus, we have shown that if $\mathbf{c}\perp y$, then $h_{0}+h_{1}=1$.
Putting the forward and reverse implications together, we conclude that:
$h_{0}+h_{1}=1\quad\Longleftrightarrow\quad\mathbf{c}\perp y.$
This completes the proof.
D.3 The derivation of the KL optimization
We want to solve the following problem:
$q^{*}(y\mid{\mathbf{b}})\;=\;\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}\Bigr{]}.$
Let
$\mathcal{L}[q(\cdot\mid{\mathbf{b}})]\;=\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},E)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}\Bigr{]}.$
By the definition of KL divergence, for discrete $y$:
$D_{\mathrm{KL}}\bigl{(}P(y)\,\|\,Q(y)\bigr{)}=\sum_{y}P(y)\,\log\frac{P(y)}{Q(y)}.$
Hence, for each $e$,
$D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}=\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
Therefore,
$\mathcal{L}[q]=\sum_{e}p(e\mid{\mathbf{b}})\,\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
We want to choose $q(y\mid{\mathbf{b}})$ to minimize $\mathcal{L}[q]$, which needs $q(y\mid{\mathbf{b}})$ to satisfy $\sum_{y}q(y\mid{\mathbf{b}})\;=\;1,\text{for each fixed }{\mathbf{b}}.$ To incorporate this constraint, we introduce a Lagrange multiplier $\lambda$ and consider the Lagrangian:
$\mathcal{J}[q,\lambda]=\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\!\biggl{[}\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\biggr{]}\;+\;\lambda\Bigl{(}\sum_{y}q(y\mid{\mathbf{b}})-1\Bigr{)}.$
To find the minimizing $q$, we set
[…]
First, note
$\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\;=\;\log p(y\mid{\mathbf{b}},e)-\log q(y\mid{\mathbf{b}}).$
Only the term $-\log q(y\mid{\mathbf{b}})$ depends on $q$. For a specific $y$,
$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\bigl{[}-\log q(y^{\prime}\mid{\mathbf{b}})\bigr{]}=\begin{cases}-\frac{1}{q(y\mid{\mathbf{b}})},&\text{if }y^{\prime}=y,\\ 0,&\text{if }y^{\prime}\neq y.\end{cases}$
Hence, for each $e$:
$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\biggl{[}\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}\biggr{]}=-\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
Summing over $e$,
$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
Also,
$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\Bigl{[}\lambda\bigl{(}\sum_{y^{\prime}}q(y^{\prime}\mid{\mathbf{b}})-1\bigr{)}\Bigr{]}=\lambda.$
Hence,
$\frac{\partial\mathcal{J}}{\partial q(y\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda.$
Setting this to zero:
$-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda=0,$
which gives
$\frac{1}{q(y\mid{\mathbf{b}})}\,\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\lambda.$
Thus
$q(y\mid{\mathbf{b}})=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
We require $\sum_{y}q(y\mid{\mathbf{b}})=1$. Then:
$1=\sum_{y}q(y\mid{\mathbf{b}})=\sum_{y}\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e).$
Since $\sum_{y}p(y\mid{\mathbf{b}},e)=1$ for each $e$, we get
$\sum_{e}p(e\mid{\mathbf{b}})\,\underbrace{\sum_{y}p(y\mid{\mathbf{b}},e)}_{=1}=\sum_{e}p(e\mid{\mathbf{b}})=1.$
Hence
$11=\frac{1}{\lambda}\cdot 1\quad\Longrightarrow\quad\lambda=$ 1 .
Therefore,
$q(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
Appendix E Algorithms
E.1 Binary Bias-Aware Generalization
Algorithm 1 Bias-Aware Generalization
1:  Input: Source data $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})_{i=1}^{M}$; hyper-parameters $\lambda_{0},\lambda_{1}$,
2:  : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding, $Pr$ means logits prior, $\operatorname{logit}(p(y))$.
3:  Stage 1: Source Training
4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
5:  Reconstruct $\hat{{\mathbf{x}}}$.
6:  Compute VAE loss $\mathcal{L}_{\text{vae}}$.
7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
8:  Calculate the bias predictor, obtain the final prediction.
9:  Compute the classification loss $\mathcal{L}_{\text{cls}}$.
10:  Add $\mathcal{L}_{\text{ind}}$ to enforce $\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y$.
11:  Compute $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{cls}}+\lambda_{0}\mathcal{L}_{\text{vae}}+\lambda_{1}\mathcal{L}_{\text{ind}}$ and optimize parameters.
12:  For source data, calculate $h_{0}=P(\hat{y}=0\mid y=0),h_{1}=P(\hat{y}=1\mid y=1)$ using $\hat{y}=f_{c}(\mathbf{c})$.
13:  Stage 2: Test-Time Adaptation
14:  Obtain $(\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}})$ with pre-trained encoders.
15:  Generate pseudo-label as $\hat{y}^{\mathcal{T}}=f_{c}(\mathbf{c}^{\mathcal{T}})$.
16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
17:  Correct the the results of bias predictor as $\phi(\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}}))=\operatorname{logit}((\frac{\sigma((\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}})))+h_{0}-1}{h_{0}+h_{1}-1}))$ ,
18:  Make the final prediction.
E.2 Multi-Class Bias-Aware Generalization
Algorithm 2 Multi-Class Bias-Aware Generalization
1:  Input: Source data $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})_{i=1}^{M}$; hyper-parameters $\lambda_{0},\lambda_{1}$,
2:  : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding
3:  Stage 1: Source Training
4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
5:  Reconstruct $\hat{{\mathbf{x}}}$.
6:  Compute VAE loss $\mathcal{L}_{\text{vae}}$.
7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
8:  Calculate the bias predictor, obtain the final prediction as equation 20.
9:  Compute the classification loss $\mathcal{L}_{\text{cls}}$.
10:  Add $\mathcal{L}_{\text{ind}}$ to enforce $\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y$.
11:  Compute $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{cls}}+\lambda_{0}\mathcal{L}_{\text{vae}}+\lambda_{1}\mathcal{L}_{\text{ind}}$ and optimize parameters.
12:  For source data, calculate Probability Confusion Matrix $\epsilon$
13:  Stage 2: Test-Time Adaptation
14:  Obtain $(\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}})$ with pre-trained encoders.
15:  Generate pseudo-label as $\hat{y}^{\mathcal{T}}=f_{c}(\mathbf{c}^{\mathcal{T}})$.
16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
17:  Correct the the results of bias predictor by least squares using $\epsilon$ according Appendix C
18:  Make the final prediction as equation 20.
Appendix F Experiment details and ablation Study
F.1 Experiment details
PACS
We evaluate our BAG model on the PACS dataset, which comprises four distinct domains (Photo, Art painting, Cartoon, Sketch) and seven object categories. Following the standard leave-one-domain-out protocol, we train on three domains and test on the held-out domain. Our backbone is a ResNet-18 pretrained on ImageNet, used to extract visual features. On top of this, we attach a variational autoencoder that separates causal factors from spurious factors, and we introduce learnable domain embeddings to encourage independence across environments. All key hyperparameters—such as the dimensions of the latent representations, the relative weighting between reconstruction and regularization losses, and the learning rates—are chosen via Bayesian optimization, with test-domain accuracy serving as the selection criterion. We train with the Adam optimizer for 70 epochs and select the final model based on its performance under test-time augmentation on the held-out domain. We used the code based on [11].
Office-Home
We evaluate our BAG model on the Office-Home dataset [13], which comprises four domains (Art, Clipart, Product, Real-World) and 65 object categories. Following the leave-one-domain-out protocol, we train on three domains and test on the held-out one. Hyperparameters for each held-out domain are optimized via Bayesian optimization; the selected values for learning rate, regularization weight, number of epochs for the $z_{s}$ classifier, and VAE weight are given in Table F.1. The full code are provided in the supplementary material.
Table F.1: Optimal hyperparameters on Office-Home
Domain Learning Rate $\lambda_{1}$ epochs (test time) $\lambda_{0}$
Art (Ar) $5.7\times 10^{-5}$ 0.030 3 $1.2\times 10^{-4}$
Clipart (Cl) $6.6\times 10^{-5}$ 0.012 3 $2.1\times 10^{-5}$
Product (Pr) $5.1\times 10^{-5}$ 0.030 4 $2.2\times 10^{-4}$
Real-World (Rw) $5.0\times 10^{-5}$ 0.046 6 $2.5\times 10^{-5}$
F.2 Ablation Study
Table F.2 presents the performance of various BAG variants on the PACS dataset and in simulation experiments. All variants outperform ERM, demonstrating the effectiveness of our two bias-utilization methods. Notably, BAG-RE and BAG-TTA each exhibit their own strengths and weaknesses, which may be due to data-specific effects. BAG-VAE performs somewhat worse than BAG, underscoring the importance of accurately identifying and disentangling $\mathbf{b}$ and $\mathbf{c}$.
Table F.2: Accuracy of different variants including BAG, BAG-VAE, BAG-RE, and BAG-TTA on four PACS domains and Synthetic data.
Algorithm PACS Synthetic
P A C S Avg Acc
BAG 96.6 86.0 77.9 73.0 83.4 97.7
BAG-VAE 95.8 84.5 78.9 66.8 81.5 94.5
BAG-RE 96.1 82.8 78.4 69.7 81.8 77.0
BAG-TTA 96.3 80.0 78.8 73.0 82.0 85.5
Appendix G Statement
This work contributes to the advancement of machine learning by challenging conventional out-of-distribution (OOD) approaches and introducing a framework that strategically leverages bias rather than eliminating it. On the positive side, our methods can improve model robustness when deployed in naturally biased environments—e.g., medical imaging systems that must generalize across hospitals with different demographic distributions. By revealing how bias can be harnessed as a constructive signal, we hope to inspire new lines of research into principled bias-aware generalization across diverse domains. Because this is foundational research evaluated on controlled benchmarks and synthetic datasets, we do not identify any immediate negative societal impacts.

### v2

@@
-Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation
+Should Bias be Eliminated? A General Framework to Use Bias for OOD Generalization
 Abstract
-Most existing methods for adapting models to out-of-distribution (OOD) domains rely on invariant representation learning to eliminate the influence of biased features.
-However, should bias always be eliminated—and if not, when should it be retained, and how can it be leveraged?
-To address these questions, we first present a theoretical analysis that explores the conditions under which biased features can be identified and effectively utilized.
-Building on this theoretical foundation, we introduce a novel framework that strategically leverages bias to complement invariant representations during inference.
-The framework comprises two key components that leverage bias in both direct and indirect ways: (1) using invariance as guidance to extract predictive ingredients from bias, and (2) exploiting identified bias to estimate the environmental condition and then use it to explore appropriate bias-aware predictors to alleviate environment gaps.
-We validate our approach through experiments on both synthetic datasets and standard domain generalization benchmarks.
-Results consistently demonstrate that our method outperforms existing approaches, underscoring its robustness and adaptability.
-footnotetext: Correspondence to: Guangyi Chen <guangyichen1994@gmail.com>
+Most approaches to out-of-distribution (OOD) generalization learn domain-invariant representations by discarding contextual bias.
+In this paper, we raise a critical question: Should bias be eliminated?
+If not, is there a general way to leverage bias for better OOD generalization?
+To answer these questions, we first provide a theoretical analysis that characterizes the circumstances in which biased features contribute positively.
+Although theoretical results show that bias may sometimes play a positive role, leveraging it effectively is non-trivial, since its harmful and beneficial components are often entangled.
+Recent advances have sought to refine the prediction of bias by presuming reliable predictions from invariant features.
+However, such assumptions may be too strong in the real world, especially when the target also shifts from training to testing domains.
+Motivated by this challenge, we introduce a framework to leverage bias in a more general scenario.
+Specifically, we employ a generative model to capture the data generation process and identify the underlying bias factors, which are then used to construct a bias-aware predictor.
+Since the bias-aware predictor may shift across environments, we first estimate the environment state to train predictors under different environments, combining them as a mixture of domain experts for the final prediction.
+Then, we build a general invariant predictor, which can be invariant under label shift to guide the adaptation of the bias-aware predictor.
+Evaluations on synthetic data and standard domain generalization benchmarks demonstrate that our method consistently outperforms both invariance only baselines, recent bias utilization approaches and advanced baselines, yielding improved robustness and adaptability.
+Machine Learning, ICML
 1 Introduction
@@
 camel classification” problem.
-As illustrated in Figure 1(a), models trained on imbalanced datasets often rely on the biased background features for prediction.
+As illustrated in Figure 1(a), models trained on imbalanced datasets may rely on biased background features for prediction.
 Specifically, cows are typically photographed in grassy fields, while camels appear in desert settings.
 Consequently, the model may mistakenly learn to associate the background context, rather than the animal itself, with the class label.
-From a causal perspective, this issue arises due to spurious correlations, where the background acts as a confounding variable influencing both the images and the associated label, as shown in Figure 1(b).
+From a causal perspective, this issue arises from spurious correlations, where the background acts as a confounding variable influencing both the images and the associated label, as shown in Figure 1(b).
 Figure 1: Illustration of data bias in the “Cow vs.
@@
 (c) An intuitive example demonstrating how humans classify an image with ambiguous content as either a cow or a camel.
-Most existing methods identify data bias as the key challenge in predicting unseen out-of-distribution (OOD) images and thus learn invariant representations to eliminate its influence.
-For instance, Invariant Risk Minimization (IRM) irm aims to learn a representation that supports a single classifier across all domains.
-Domain Adversarial Neural Networks (DANN) dann employ adversarial learning to encourage domain-invariant representations.
-Similarly, Self-Training lee2013pseudo improves generalization by generating pseudo-labels for target domain data based on learned invariant predictions.
-Some de-bias methods qi2020two leverage causal inference techniques, such as the back-door and front-door rules, to discover true causal relationships from spurious correlations.
-More discussions on related work can be found in Appendix A.
-The core assumption of these methods is that biased features are uninformative, and only features invariant across different domains should be transferred.
-In this paper, we challenge this assumption and raise a critical question: Should bias always be eliminated?
-We provide an intuitive motivation for this question in part (c) of Figure 1.
-As humans, when classifying an image with ambiguous content as either a cow or a camel, we often rely on contextual background cues to make an educated guess.
-This intuitively suggests that, rather than being entirely detrimental, in some scenarios, the bias derived from background information can sometimes serve as a useful reference in prediction.
-Mathematically, it indicates that the domain-specific (biased) features and the prediction target are conditionally dependent given the invariant content, which conflicts with the assumption of uninformative bias in invariant representation learning.
-Although bias can contain useful information, using it for prediction is challenging and potentially harmful.
-Recent studies chang2019domain ; wei2018general ; kong2023partial ; erm ; eastwood2024spuriosity have explored the use of biased features, with the most relevant being Stable Feature Boosting (SFB) eastwood2024spuriosity , which adapts conditionally independent biased features using pseudo-labels from stable ones at test time.
-However, these methods lack theoretical analysis of the conditions under which data bias can be reliably identified and effectively leveraged.
-Overlooking this can be risky, as bias often contains both “good” components (e.g., context that aids prediction) and “bad” components (e.g., spurious noise that harms prediction), making it essential to carefully disentangle and isolate the useful information.
-To address this challenge, we present a theoretical analysis for the identification and utilization conditions of the data bias.
-Our results show that bias can indeed be leveraged, provided there is sufficient distributional variation across domains/labels and the relationship between the biased features and labels remains unblocked.
-Notably, even in more general scenarios where domain shifts affect the distributions of both the input data and the label, our analysis demonstrates that biased features can still offer valuable information for improving model performance in the inference stage.
-Building on our theoretical analysis, we propose a principled generative framework to identify data bias and effectively leverage it for prediction.
-Specifically, we employ a Variational Autoencoder (VAE) to learn latent representations and disentangle content features from biased ones in an unsupervised manner.
-We then leverage the bias for prediction through both direct and indirect strategies.
-For the direct approach, we extract the beneficial components of the bias by aligning them with invariant predictions.
-It is achieved by taking the invariant predictions as pseudo-labels to guide the refinement of biased features that filter out harmful bias components, thereby retaining only the useful context for prediction.
-For the indirect approach, we utilize the biased feature to estimate the underlying environment context and adapt the prediction accordingly.
-In particular, we decompose the predictor into two components: an invariant predictor that takes content features as input, and a set of domain experts conditioned on biased features.
-Assuming that the target domain can be represented as a linear combination of source domains, we obtain a bias-aware predictor by the weighted mixture of experts using the inferred bias, enabling better alignment with the target distribution.
-The main contributions of this paper are summarized as follows:
-  • We revisit the role of bias in OOD generalization. Unlike prior work that aims to eliminate bias, we highlight scenarios where bias can be a useful signal for improving prediction.
-  • We provide a theoretical analysis to explore conditions under which data bias can be reliably identified and effectively leveraged.
-  • We introduce a principled framework to leverage bias in both direct and indirect ways and validate our methods through experiments on both synthetic and real-world datasets.
+Most OOD generalization methods, such as IRM (Arjovsky et al., 2019), DANN (Ganin et al., 2016), and Self-Training (Lee and others, 2013), regard data bias as the primary obstacle and therefore aim to learn domain-invariant representations (see Appendix A for further discussion).
+The core assumption of these methods is that biased features are uninformative and that only invariant features should be transferred.
+As illustrated in Figure 1(c), when an image is ambiguous (e.g., cow vs.
+camel), humans often rely on background context to make an educated guess.
+This intuitively suggests that, rather than being entirely detrimental, in some scenarios, the bias derived from background information can serve as a useful reference in the prediction process.
+To analyze the effect of bias, we provide a theoretical framework that characterizes when data bias is identifiable and beneficial for prediction.
+Our results show that, in general, biased features can be exploited when they can be disentangled from invariant content, and the dependence between biased features and labels is not screened off by invariant content.
+It indicates that we can disentangle the bias component and study its predictive contribution, enabling bias features to provide complementary information that can be effectively exploited at inference.
+Although bias contains useful information, using it for prediction is challenging.
+The most relevant prior work that leverages bias for OOD generalization is Stable Feature Boosting (SFB)  (Eastwood et al., 2023).
+SFB splits the representation into invariant and biased components, using the predictions based on the invariant features as pseudo-labels for test-time adaptation of the biased component.
+However, this method relies on the assumption that invariant features alone can yield accurate and robust predictions.
+In particular, it presumes that (1) invariant features contain sufficient information to support prediction, and (2) the prediction target itself remains invariant, i.e., no label shift (Lipton et al., 2018) occurs.
+In real-world applications, the invariant content may be under-learned due to limited domain heterogeneity, insufficient supervision, or an intrinsic lack of information.
+Meanwhile, the prediction using the invariant feature may shift in the testing domain, as shown in the common label shift setting (Li et al., 2023b; Wu et al., 2021; Koh et al., 2021).
+To leverage bias in more general and realistic settings, we adopt a generative model that factorizes representations into invariant content and contextual bias in an identifiable manner.
+Building on the identified bias features, we uncover two pathways through which bias can contribute to prediction.
+Firstly, we use the bias features to estimate environmental states and design an environment routing mechanism that leverages these states to gate a mixture of predictors, each conditioned on a specific environment.
+This environment estimation enables the model to exploit bias as a useful signal, particularly in cases where the invariant content alone is insufficient for accurate prediction.
+Secondly, we extend the bias correction in SFB (Eastwood et al., 2023) with an adaptive label prior.
+We compose predictors using bias and content together by Bayes’ Rule into two parts: bias predictor and invariant predictor which consists of content predictor and label prior.
+In addition to correcting the bias predictor using content predictions as pseudo-labels with a fixed label prior, we further learn the label prior in the training stage.
+This modification can effectively address the problem when both label shift and covariate shift co-exist.
+The contributions of this paper are summarized as follows:
+  • We revisit the role of bias in OOD generalization, highlighting circumstances in which bias offers a constructive signal rather than a nuisance to be eliminated.
+  • We provide a theoretical analysis that specifies identification and utility conditions for data bias, including cases with concurrent shifts in input and label distributions.
+  • We introduce a general bias-aware framework that makes predictions with an environment routing mechanism and conducts adaptive bias correction.
+  • We evaluate the proposed method on synthetic and real-world datasets and show that it consistently outperforms baseline approaches across both settings.
 2 Theoretical Analysis
 2.1 Problem Setup and Data Generation Process
+Notation and problem setup.
+We formulate the task as a multi-source domain generalization problem, aiming to learn a predictor that remains robust in an unseen target domain using only data from multiple source domains.
+Let ${\mathbf{x}}_{k}$ denote a high-dimensional image observation ${\mathbf{x}}_{k}:=[x_{1},\cdots,x_{n_{x}}]\in{\mathcal{X}}\subset\mathbb{R}^{n_{x}}$, and $y_{k}$ denote the corresponding label.
+With access to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$, with each domain represented as $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})=({\mathbf{x}}^{\mathcal{S}_{i}}_{k},y^{\mathcal{S}_{i}}_{k})^{m_{i}}_{k=1}$, our goal is to learn a predictor $f({\mathbf{x}})$ that can generalize well to the new domain $\mathcal{T}$ where $\mathcal{T}\notin\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
 Figure 2: A graphical representation of the data generation process.
-The content variables ${\mathbf{c}}$ are invariant to environment changes, while the bias variables ${\mathbf{b}}$ vary across environments.
+The content variables $\mathbf{c}$} are invariant to environment changes once label $y$ is given, while the bias variables $\mathbf{b}$} vary across environments.
 Observed data $\mathbf{x}$} is generated by $g({\mathbf{c}},{\mathbf{b}})$, with $\mathbf{c}$} and $\mathbf{b}$} forming the latent variable $\mathbf{z}$}.
 The environment $e$ affects $\mathbf{b}$} but not $\mathbf{c}$}, and the target $y$ may also depend on $e$, reflecting complex data-environment interactions.
-Notation and problem setup.
-We formulate the task as a multi-source domain generalization problem, aiming to learn a predictor that remains robust in an unseen target domain using only data from multiple source domains.
-Let ${\mathbf{x}}_{k}$ denote a high-dimension image observation ${\mathbf{x}}_{k}:=[x_{1},\cdots,x_{n_{x}}]\in{\mathcal{X}}\subset\mathbb{R}^{n_{x}}$, and $y$ denotes the corresponding label.
-With access to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$, with each domain represented as $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})=({\mathbf{x}}^{\mathcal{S}_{i}}_{k},y^{\mathcal{S}i}_{k})^{m_{i}}_{k=1}$, our goal is to learn a predictor $f({\mathbf{x}})$ that can generalize well to the new domain $\mathcal{T}$ where $\mathcal{T}\nsubseteq\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
 Data generation process.
 To illustrate the origin and impact of bias, we model the data generation process using a latent variable model.
-Let ${\mathbf{c}}:=[c_{1},\cdots,c_{n_{c}}]\in\mathcal{C}\subset\mathbb{R}^{n_{c}}$ denotes content variables and ${\mathbf{b}}:=[b_{1},\cdots,b_{n_{b}}]\in\mathcal{B}\subset\mathbb{R}^{n_{b}}$ denotes the bias.
-As shown in Figure 2, the generation process of the observation ${\mathbf{x}}$ is defined as ${\mathbf{x}}:=g({\mathbf{c}},{\mathbf{b}})$, where ${\mathbf{z}}:=[{\mathbf{c}},{\mathbf{b}}]$ with $n_{z}=n_{c}+n_{b}$ dimension denotes the latent variables and $g:[{\mathbf{c}},{\mathbf{b}}]\mapsto{\mathbf{x}}$ denotes the generating function.
-To distinguish ${\mathbf{c}}$ and ${\mathbf{b}}$, we introduce an environment variable $e$, where the content ${\mathbf{c}}$ remains invariant despite changes of the environment, whereas the bias ${\mathbf{b}}$ varies as the environment changes.
+Let ${\mathbf{c}}:=[c_{1},\cdots,c_{n_{c}}]\in\mathcal{C}\subset\mathbb{R}^{n_{c}}$ denote content variables and ${\mathbf{b}}:=[b_{1},\cdots,b_{n_{b}}]\in\mathcal{B}\subset\mathbb{R}^{n_{b}}$ denote the bias.
+As shown in Figure 2, the generation process of the observation $\mathbf{x}$} is defined as ${\mathbf{x}}:=g({\mathbf{c}},{\mathbf{b}})$, where ${\mathbf{z}}:=[{\mathbf{c}},{\mathbf{b}}]$ with $n_{z}=n_{c}+n_{b}$ dimensions denotes the latent variables and $g:[{\mathbf{c}},{\mathbf{b}}]\mapsto{\mathbf{x}}$ denotes the generating function.
+To distinguish $\mathbf{c}$} and $\mathbf{b}$}, we introduce an environment variable $e$, where the content $\mathbf{c}$} remains invariant despite changes in the environment once $y$ is given, whereas the bias $\mathbf{b}$} varies as the environment changes.
 Taking the “cow vs.
-camel classification” problem for illustration, ${\mathbf{c}}$ denotes the characteristics of the object, like the color/shape of cow; while ${\mathbf{b}}$ represents the background information.
+camel classification” problem for illustration, $\mathbf{c}$} denotes the characteristics of the object, like the color/shape of the animal, while $\mathbf{b}$} represents the background information.
 Note that we consider a more general case where $y$ may also be influenced by the environment, allowing for a more complex relationship modeling between the data and the environment.
@@
 deserts) in real-world scenarios.
-Here we highlight two causal paths between bias and label target, $y\rightarrow{\mathbf{b}}$ and $y\leftarrow e\rightarrow{\mathbf{b}}$, which motivate the direct and indirect ways to leverage the bias for prediction, respectively.
-2.2 Theoretical conditions for identifiability and inference
-Bias identification.
-We demonstrate that, with the advancement of representation learning, content and bias variables can be effectively disentangled with an identifiability guarantee.
-Since disentangling content and bias is fundamental for leveraging biased information, we briefly summarize the mild conditions under which such disentanglement can be achieved.
+Here we highlight two causal paths between bias and label target, $y\leftarrow e\rightarrow{\mathbf{b}}$ and $y\rightarrow{\mathbf{b}}$, which motivate two ways to leverage the bias for prediction.
+First, we can estimate the environment state from $\mathbf{b}$} and use such information to adjust the prediction.
+Second, we can leverage content to guide the direct prediction from $\mathbf{b}$} to $y$.
+2.2 Theoretical Analysis For Identifiability and Usage
+Bias and Content identification.
+We demonstrate that content and bias variables can be effectively disentangled with an identifiability guarantee.
+We extend the theoretical results of the previous causal representation learning methods to align with our formulation, with the detailed modification, proof and discussions that can be found in Sections B.1 and B.2.
+Such block-wise identification enforces invertible mappings between learned and ground-truth factors: $\hat{\mathbf{c}}=f({\mathbf{c}})$, $\hat{\mathbf{b}}=g({\mathbf{b}})$, with $f,g$ being invertible.
 It is important to note that this identifiability result has already been established in previous work and is not claimed as a contribution of this paper.
-However, despite the identification results, existing methods rely on invariant prediction (including invariant content variable and invariant relation between bias variable and label).
-Different from them, building on this identifiability framework, we go further by exploring how to effectively utilize biased data to enhance OOD performance.
+However, despite the identification results, existing methods still rely on invariant prediction (including invariant content variable and invariant relation between bias variable and label).
+Unlikely, building on this identifiability framework, we go further by exploring how to effectively utilize biased data to enhance OOD performance.
 Lemma 2.1.
-(Block-wise Identification of ${\mathbf{c}}$ and ${\mathbf{b}}$ kong2023partial .) Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
-  • A1 (Smooth and Positive Density): The probability density function of latent variables is smooth and positive, i.e. $P_{\mathbf{z}|\mathbf{e}}>0$.
-  • A2 (Conditional Independence): Conditioned on $\mathbf{e}$, each $z_{i}$ is independent of any other $z_{j}$ for $i.j\in\{1,\cdots,n\},i\neq j$, i.e., $\log P_{\mathbf{z}|\mathbf{e},\mathbf{y}}=\sum_{i=1}^{n}q(z_{i},\mathbf{e},\mathbf{y})$, where $q_{i}(z_{i},\mathbf{e},\mathbf{y})$ is the log density of the conditional distribution, i.e., $q_{i}:\log P_{z_{i}|\mathbf{e},\mathbf{y}}$.
-  • A3 (Linear Independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$, there exists $n_{b}+1$ values of $\mathbf{e}$, i.e., $e_{j}$ with $j=0,1,.\cdots,n_{b}$ such that these $n_{b}$ vectors $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})-\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{0})$ with $j=1,\cdots,n_{b}$ are linearly independent, where vector $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})$ is defined as follows:
-    $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}})=\left(\frac{\partial q_{1}({\mathbf{b}}_{1},{\mathbf{e}})}{\partial{\mathbf{b}}_{1}},\cdots\frac{\partial q_{n_{b}}({\mathbf{b}}_{n_{b}},{\mathbf{e}})}{\partial{\mathbf{b}}_{n_{b}}}\right),$ (1)
-  • A4 (Domain Variability:) There exist two values of $\mathbf{e}$, i.e., $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, s.t., for any set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability measure and cannot be expressed as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$, we have
-[…]
+(Block-wise Identification of $\mathbf{c}$} and $\mathbf{b}$} (Kong et al., 2022)). Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
+  • A1 (Smooth and positive density): $p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})$ is smooth with strictly positive density on its support.
+  • A2 (Conditional independence and content invariance): Given $(\mathbf{e},\mathbf{y})$, the coordinates of $\mathbf{z}$ are independent, $\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n_{z}}q_{i}(z_{i};\mathbf{e},\mathbf{y}),$
+  • A3 (Linear independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$ (and the relevant $\mathbf{y}$), there exist $n_{b}+1$ environments $e_{0},e_{1},\ldots,e_{n_{b}}$ such that the $n_{b}$ vectors $\bm{w}(\mathbf{b},e_{j},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y}),\quad j=1,\ldots,n_{b},$ are linearly independent, where $\bm{w}(\mathbf{b},e_{i},\mathbf{y}):=\Big(\tfrac{\partial q_{1}(b_{1};e_{i},\mathbf{y})}{\partial b_{1}},\ldots,\tfrac{\partial q_{n_{b}}(b_{n_{b}};e_{i},\mathbf{y})}{\partial b_{n_{b}}}\Big)^{\top}.$
+  • A4 (Domain variability): There exist $e_{i}\neq e_{j}$ such that, for the relevant $\mathbf{y}$ and for any measurable set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability that cannot be written as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$, $\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{i},\mathbf{y})\,d\mathbf{z}\;\neq\;\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{j},\mathbf{y})\,d\mathbf{z}.$
 Then the learned $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$ are block-wise identifiable.
-The detailed proof can be found in Appendix B.1 .
-Such block-wise identification ensures that the learned content and bias representations, $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$, are functions of the ground truth content ${\mathbf{c}}$ and bias ${\mathbf{b}}$, respectively.
-Use bias for inference.
+Definition 2.2.
+(Unblocked Influence): Let ${\bm{z}}_{1}\in\mathbb{R}^{d_{1}}$, ${\bm{z}}_{2}\in\mathbb{R}^{d_{2}}$, and ${\bm{z}}_{3}\in\mathbb{R}^{d_{3}}$ be random vectors.
+We say that unblocked influence for ${\bm{z}}_{1}$, ${\bm{z}}_{2}$ given ${\bm{z}}_{3}$ holds if there exists a causal path from ${\bm{z}}_{1}$ to ${\bm{z}}_{3}$ that is not blocked by ${\bm{z}}_{2}$.
+Formally, this condition requires ${\bm{z}}_{1}\not\!\perp\!\!\!\perp{\bm{z}}_{3}\;\mid\;{\bm{z}}_{2}.$
+Lemma 2.3.
+The bias $\mathbf{b}$} can be effectively utilized for prediction if Assumptions A1 through A4 hold and there is unblocked influence for $\mathbf{b}$},$y$ given $\mathbf{c}$}.
+When can bias be leveraged?
 After identifying bias, we further provide theoretical insights into the conditions under which it can be leveraged to enhance predictive performance.
 Specifically, we show that the identified bias can contribute to prediction as long as there exists an active (i.e., unblocked) causal path from the bias to the prediction target, meaning the good information carried by the bias is not fully mediated or blocked by the content.
-[…]
-Lemma 2.2.
-The bias ${\mathbf{b}}$ can be effectively utilized for prediction if Assumptions A1 through A5 hold.
+As shown in Figure 2, there exists a path $y\rightarrow{\mathbf{b}}$ which indicates that $\mathbf{b}$} is not blocked with $y$ by $\mathbf{c}$}.
+And it should be noted that even without this path, since $\mathbf{b}\leftarrow e\rightarrow y$, $\mathbf{b}$} is not blocked with $y$ by $\mathbf{c}$}.
 As shown in Lemma 2.1, under Assumptions A1–A4, both content and bias information ${\mathbf{c}},{\mathbf{b}}$ can be well identified.
 It indicates that we can extract ${\mathbf{c}},{\mathbf{b}}$ containing all the information as the ground-truth variables.
-Assumption A5 states that
-$P(y\mid{\mathbf{c}},{\mathbf{b}})\;\neq\;P(y\mid{\mathbf{c}})\quad\text{on a set of nonzero measure.}$.
-In Bayesian decision theory (for any proper loss, e.g. log-loss or squared error), the Bayes-optimal predictor using $({\mathbf{c}},{\mathbf{b}})$ is
-$f^{*}({\mathbf{c}},{\mathbf{b}})=\arg\min_{f}\;\mathbb{E}\bigl{[}\ell\bigl{(}y,f({\mathbf{c}},{\mathbf{b}})\bigr{)}\bigr{]},$
-which depends on the full conditional $P(y\mid{\mathbf{c}},{\mathbf{b}})$. The best predictor using only ${\mathbf{c}}$ is
-$\arg\min_{f}\;\mathbb{E}\bigl{[}\ell(y,f({\mathbf{c}}))\bigr{]},$
-which depends only on $P(y\mid{\mathbf{c}})$.
-Since Assumption A5 ensures that the conditional distributions $p(y\mid{\mathbf{c}})$ and $p(y\mid{\mathbf{c}},{\mathbf{b}})$ differ, the joint predictor $f^{*}({\mathbf{c}},{\mathbf{b}})$ achieves a strictly lower expected loss compared to any predictor that relies solely on ${\mathbf{c}}$.
+Unblocked influence states that
+$P(y\mid{\mathbf{c}},{\mathbf{b}})\;\neq\;p(y\mid{\mathbf{c}})\quad\text{on a set of nonzero measure.}$
+In Bayesian decision theory (for any proper loss, e.g.
+log‐loss or squared error), the Bayes‐optimal predictor using $({\mathbf{c}},{\mathbf{b}})$ is $f^{*}({\mathbf{c}},{\mathbf{b}})=\arg\min_{f}\;\mathbb{E}\bigl[\ell\bigl(y,f({\mathbf{c}},{\mathbf{b}})\bigr)\bigr],$ which depends on the full conditional $p(y\mid{\mathbf{c}},{\mathbf{b}})$.
+The best predictor using only $\mathbf{c}$} is $\arg\min_{f}\;\mathbb{E}\bigl[\ell(y,f({\mathbf{c}}))\bigr],$ which depends only on $p(y\mid{\mathbf{c}})$.
+Since unblocked influence ensures that the conditional distributions $p(y\mid{\mathbf{c}})$ and $p(y\mid{\mathbf{c}},{\mathbf{b}})$ differ, the joint predictor $f^{*}({\mathbf{c}},{\mathbf{b}})$ achieves a strictly lower expected loss compared to any predictor that relies solely on $\mathbf{c}$}.
 This confirms that the bias variable $\mathbf{b}$} can indeed be exploited to improve predictive accuracy.
-A detailed proof, demonstrating that incorporating ${\mathbf{b}}$ leads to a better predictor under log-loss or squared error, is provided in Appendix B.2.
-2.3 Understand bias’ contribution in prediction
-Furthermore, we explore how bias contributes to the predictions.
-To illustrate the core insights of our approach, we begin with a binary classification setting.
-The extension to the multi-class scenario is in Appendix C.
-The core intuition is that $\mathbf{c}$ and $\mathbf{b}$ are conditionally independent given $y$, which indicates that we can decompose $p(y\mid\mathbf{c},\mathbf{b})$ into three parts, including a bias-aware predictor, an invariant predictor, and a label prior.
-Theorem 2.3.
-[Decomposition of Marginal Probability] Given the data generation process illustrated in Figure 2, we have:
-$\ p(y=1\mid\mathbf{c},\mathbf{b})=\sigma\biggl{(}\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{b})\bigr{)}}_{\text{Bias-aware Predictor}}+\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c})\bigr{)}}_{\text{Invariant Predictor}}-\underbrace{\operatorname{logit}\bigl{(}p(y=1)\bigr{)}}_{\text{Label Prior}}\biggr{)},$
-where $\sigma$ is the sigmoid function. The last two terms combined can be viewed as an invariant part of the model across all domains.
-Proof sketch of Theorem 2.3
-We begin by reformulating $p(y\mid\mathbf{c},\mathbf{b})$ to make its structure more explicit. Specifically, we have:
-$\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c},\mathbf{b})\bigr{)}=\frac{p(\mathbf{c},\mathbf{b}\mid y=1)\,p(y=1)}{p(\mathbf{c},\mathbf{b}\mid y=0)\,p(y=0)}\quad\quad\text{consider }\mathbf{c}\,\perp\!\!\!\perp\,(\mathbf{b})\,\mid\,y:$
-$\quad=\frac{p(\mathbf{b}\mid y=1)\,p(\mathbf{c}\mid y=1)\,p(y=1)}{p(\mathbf{b}\mid y=0)\,p(\mathbf{c}\mid y=0)\,p(y=0)}=\frac{p\bigl{(}y=1\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=1\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=1\bigr{)}}}{p\bigl{(}y=0\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=0\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=0\bigr{)}}}.$
-A more detailed proof is provided in Appendix D.1.
+A detailed proof, demonstrating that incorporating $\mathbf{b}$} leads to a better predictor under log-loss or squared error, is provided in Section B.3.
+3 Preliminary Work on Bias Correction
+Recently, Stable Feature Boosting (SFB) (Eastwood et al., 2023) proposed a bias correction method that explores how bias can be leveraged for prediction.
+The core intuition is that predictions derived from invariant features are reliable across environments, and thus can serve as a reference signal for correcting the predictions from biased features.
+Concretely, SFB treats the predictions from invariant features as pseudo-labels and uses them to adjust the bias predictor, thereby reducing spurious correlations while still exploiting useful signals embedded in bias.
+Reliable invariant prediction assumption.
+This assumption states that $p(y=1\mid{\mathbf{c}})$ remains stable across all domains, which indicates that $y\perp\!\!\!\perp e\mid{\mathbf{c}}$.
+Decomposition of Marginal Probability.
+Specifically, according to Theorem 4.4 in (Eastwood et al., 2023), consider ${\mathbf{b}}\perp\!\!\!\perp{\mathbf{c}}\mid y$, the prediction process can be decomposed as follows:
+$\ p(y=1\mid\mathbf{c},\mathbf{b})=\sigma\biggl(\operatorname{logit}\bigl(p(y=1\mid\mathbf{b})\bigr)$
+$+\operatorname{logit}\bigl(p(y=1\mid\mathbf{c})\bigr)-\operatorname{logit}\bigl(p(y=1)\bigr)\biggr)$
+where $\sigma$ is the sigmoid function.
+By the reliable invariant prediction assumption, the predicted outputs using invariant features can serve as the pseudo labels in the correction process, and the label prior $p(y=1)$ can be regarded as a constant in all domains.
+Correction with Pseudo Labels
+Consider the predictor $f_{c}:{\mathbf{c}}\rightarrow y$.
+Define pseudo label prediction as $\hat{y}=f_{c}({\mathbf{c}})$.
+Let $h_{0}\;=\;p(\hat{y}=0\mid y=0),\quad h_{1}\;=\;p(\hat{y}=1\mid y=1).$ In the case that $\mathbf{c}$} contains the information for predicting $y$, it directly follows $h_{0}+h_{1}\;>\;1.$ Intuitively, if $c$ helps distinguish between $y=0$ and $y=1$, then a predictor based on $\mathbf{c}$} should be corrected more often than chance.
+Then, the correction of the bias predictor can be formulated as follows:
+$p\bigl(y=1\mid{\mathbf{b}}\bigr)\;=\;\frac{\Pr\bigl(\hat{y}=1\mid{\mathbf{b}}\bigr)\;+\;h_{0}\;-\;1}{\,h_{0}+h_{1}\;-\;1\,}.$ (1)
+Since $\hat{y}$ is a function of $\mathbf{c}$}, this correction relies on the assumption that $\mathbf{c}$} depends only on $y$, and the quantities $h_{0}$ and $h_{1}$ remain constant across all domains.
+Due to space constraints, we present only the core idea and method here, and refer readers to (Eastwood et al., 2023) and Appendix B for detailed derivations.
+Challenges.
+The method relies on the assumption of reliable invariant prediction, i.e., $y\perp\!\!\!\perp e\mid{\mathbf{c}}$.
+However, as illustrated in Figure 2, this assumption no longer holds under label shift.
+In such cases, the prediction $p(y=1\mid{\mathbf{b}})$ is also affected by the confounding influence of $\mathbf{e}$}.
+Ignoring $\mathbf{e}$} can therefore lead to suboptimal predictors.
+4 Method
+Figure 3: Overall framework of the BAG method.
+The framework consists of three main modules: representation learning, predictor training, and adaptation.
+In the representation learning stage, we employ a VAE to disentangle content and bias variables.
+For prediction, we construct a content predictor, a label prior, and a bias-aware predictor that reweights multiple domain experts using a domain estimator and also retrain under labels from content predictor in the test stage.
+Three predictor components work together for the final prediction.
+In this section, we introduce the Bias-Aware Generalization (BAG) framework, which systematically incorporates bias into the prediction process to improve OOD performance.
+The whole pipeline can be found in Figure 3.
+We start by learning disentangled representations so that $\mathbf{b}$} and $\mathbf{c}$} align with the data-generating process, producing factorized bias and content.
+Accordingly, we propose (i) Bias-Aware Prediction with an Environment Routing Mechanism and (ii) Bias Correction with an Adaptive Label Prior.
+For each component, we provide a theoretical analysis with a binary classification case and corresponding method details, while the multi-class extension and the full algorithm are deferred to Appendices C, D.1 and D.2.
+4.1 Representation Learning and Disentanglement
+Generative modeling.
+Following Lemma 2.1, we take the observed data $\mathbf{x}$} from the source domains as input to the encoder, which maps it into the latent variable $\mathbf{z}$}.
+We then decode $\mathbf{z}$} to reconstruct $\hat{{\mathbf{x}}}$ and train the encoder–decoder using the standard VAE loss:
+$\mathcal{L}_{\text{vae}}\;=\;\mathbb{E}_{{\mathbf{x}}}\bigl[\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|^{2}\;+\;\beta\mathrm{KL}\bigl(q_{\phi}(\mathbf{z}\mid{\mathbf{x}})\,\|\,p(\mathbf{z})\bigr)\bigr],$ (2)
+where $\beta$ is a hyper-parameter to balance the reconstruction loss and the KL divergence, which is set to 1 in our approach.
+The prior distribution $p({\mathbf{z}})$ is chosen as an isotropic Gaussian.
+Minimizing Equation 2 encourages $(\mathbf{c},\mathbf{b})$ to be identifiable in a block-wise sense and captures both invariant and domain-specific factors for further modeling.
+Regularization.
+Following (Jiang and Veitch, 2022), we leverage a necessary but not sufficient condition for $\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y$, that is $\mathbb{E}\bigl[\mathbf{c}\cdot(\mathbf{b}-\mathbb{E}[\mathbf{b}\mid y])\bigr]=0$.
+Specifically, we add a regularization term to constrain the conditional independence, which can be written as:
+$\mathcal{L}_{ind}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{c}_{i}\cdot\bigr(\mathbf{b}_{i}-\frac{1}{\bigl|\{j:y_{j}=y_{i}\}\bigr|}\sum_{j:y_{j}=y_{i}}\mathbf{b}_{j}\bigr).$
+After learning $\mathbf{b}$} and $\mathbf{c}$}, our goal is to predict $y$ in the target domain via $p(y\mid{\mathbf{b}},{\mathbf{c}})$.
+4.2 Bias-aware Prediction with Environment Routing Mechanism
+Bias-aware predictors.
+Only relying on content to guide the bias predictor has practical limitations: when $\mathbf{c}$} carries little information about $y$, the pseudo-labels $\hat{y}$ may be nearly random, making such guidance ineffective.
+Motivated by the causal path $y\leftarrow e\rightarrow{\mathbf{b}}$, we instead propose further leveraging bias to infer the environment $e$, and then using $e$ to route predictions.
+This improves $p(y\mid{\mathbf{b}})$ by explicitly modeling environmental context, independent of $\mathbf{c}$}, and yields more robust bias utilization.
+The objective can be formulated as the following optimization task:
+$q^{*}(y\!\mid\!{\mathbf{b}})=\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}{\textstyle\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\bigl[D_{\mathrm{KL}}(p(y\!\mid\!{\mathbf{b}},e)\,\|\,q(y\!\mid\!{\mathbf{b}}))\bigr]}.$ (3)
+This suggests identifying a predictor that minimizes the expected Kullback-Leibler (KL) divergence from $p(y\mid{\mathbf{b}},e)$ across all environments.
+The Bayes-optimal solution (Ohn and Lin, 2023) is
+$q^{*}(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)$
+This implies that we can leverage the identified bias to infer the environmental context and then develop a suitable bias-aware predictor $p(y\mid{\mathbf{b}})$ with a set of domain experts $p(y\mid{\mathbf{b}},e)$ to bridge the gaps between environments.
+The detailed derivation can be found in Section B.4.
+Environment routing mechanism.
+To exploit the bias $\mathbf{b}$ in an indirect manner, we introduce a set of $M$ learnable domain embeddings $\{e_{1},e_{2},\ldots,e_{M}\}$, which correspond to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
+Each embedding $e_{k}$ captures a distinct “mode” or style that may emerge in the source training data.
+Intuitively, this design follows the mixture of experts principle, allowing each expert to specialize in a subset of the training distribution.
+In terms of model design, each $e_{k}$ is a learnable vector (or a small neural component) trained to represent a particular domain bias.
+We implement a domain classifier $p(e=e_{i}\mid{\mathbf{b}})$ via a softmax layer over linear outputs to produce scalar weights.
+These weights indicate how well each expert $e_{i}$ aligns with the domain-specific bias $\mathbf{b}$.
+For each expert, we parametrize $p\bigl(y\mid\mathbf{b},e=e_{i}\bigr)$ as an MLP network, whose input is the combination of $\mathbf{b}$} and $e_{i}$.
+It models the expert specializing in domain $e_{i}$, making specific predictions based on the latent representation $\mathbf{b}$}.
+Formally, we calculate the bias-aware predictor $f_{b}$ (i.e., $p(y\mid{\mathbf{b}})$) by reweighting domain-specific predictors:
+$f_{b}(\mathbf{b})\;=\;\sum_{i=1}^{M}p\bigl(y\mid\mathbf{b},e=e_{i}\bigr)\,p(e=e_{i}\mid{\mathbf{b}}).$ (4)
+4.3 Bias Correction with Adaptive Label Prior
+Existing bias correction methods, such as SFB (Eastwood et al., 2023), rely on content predictions as pseudo-labels to adjust the bias predictor under a fixed label prior.
+While effective in certain settings, this assumption breaks down under label shift, where the label distribution varies across environments.
+In such cases, a fixed prior leads to systematic miscalibration and limits the correction’s effectiveness.
+To address this issue, we introduce an adaptive label prior that dynamically accounts for shifts in label distribution and identifies a new invariant component under label shift, thereby enabling more accurate and robust bias correction across domains.
+Since $p(y\mid{\mathbf{c}})$ is not invariant under label shift, we introduce an alternative decomposition to retain the invariant components while leveraging bias.
+Theorem 4.1 (Decomposition under label shift).
+Given the data generation process illustrated in Figure 2, we have
+$p(y=1\mid\mathbf{c},\mathbf{b})=\sigma\Biggl(\underbrace{\operatorname{logit}\bigl(p(y=1\mid\mathbf{b})\bigr)}_{\text{Bias Predictor}}$ (5)
+$+\underbrace{\log\frac{p(y=1\mid\mathbf{c})/p(y=1)}{p(y=0\mid\mathbf{c})/p(y=0)}}_{\text{Invariant Predictor}}\Biggr)$
+Due to the label shift, the content prediction $p(y\mid{\mathbf{c}})$ is not invariant, and the label prior $p(y)$ is not constant. Interestingly, we found that the composition of both is invariant.
 Discussion on invariant prediction.
-We observe that $p(\mathbf{c}\mid y)$ remains invariant across domains, since $y$ d-separate ${\mathbf{c}}$ and the environment variable $e$.
+Although $p(y\mid\mathbf{c})$ is not invariant, we observe that $p(\mathbf{c}\mid y)$ remains invariant across domains, since $y$ d-separates $\mathbf{c}$} from the environment variable $e$.
 Formally, $\forall\,e_{i},e_{j}\in E_{all}$, since $\mathbf{c}$ only depends on $y$, it follows that
-$p(\mathbf{c}\mid y=1,\,e=e_{i})\;=\;p(\mathbf{c}\mid y=1,\,e=e_{j})\Rightarrow\frac{p(y=1,\,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}\;=\;\frac{p(y=1,\,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
+$p(\mathbf{c}\mid y=1,\,e=e_{i})\;=\;p(\mathbf{c}\mid y=1,\,e=e_{j})\Rightarrow$
+$\frac{p(y=1,\,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}\;=\;\frac{p(y=1,\,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
 which implies $\tfrac{p(y\mid\mathbf{c})}{p(y)}$ is an invariant quantity across all domains.
-2.4 Leverage bias for better prediction
-As shown in Theorem 2.3, we have established that the bias affects prediction solely through the bias-dependent component $p(y=1\mid\mathbf{b})$.
-Consequently, ensuring the reliability of this bias-aware prediction becomes crucial for effectively leveraging bias.
-However, relying on this term without additional assumptions poses challenges, as the bias-aware prediction component can vary significantly across environments.
-In the absence of proper constraints, incorporating it naively would revert the model to standard empirical risk minimization (ERM), failing to account for bias robustness.
-Furthermore, as shown in the causal graph in Figure 2, the term $p(y=1\mid\mathbf{b})$ can be decomposed into two components: a direct path $\mathbf{b}\rightarrow y$ and an indirect path $\mathbf{b}\leftarrow e\rightarrow y$.
-Motivated by this structure, we propose to leverage bias for prediction in both direct and indirect ways.
-The direct approach uses invariant features to guide the extraction of the predictive (i.e., positively correlated) component of the bias.
-In contrast, the indirect approach utilizes the bias to infer the environment $e$, which in turn informs the selection of an appropriate predictor tailored to that environment.
-2.4.1 Using bias directly
-The core idea of directly leveraging bias for the prediction is to extract its beneficial components, guided by the information encoded in invariant features.
-Assumption 2.4 (Sufficient Information of ${\mathbf{c}}$).
+It motivates us to jointly learn the content predictor and the learnable label prior.
+The detailed proof can be found in Section B.5.
+Predictors
+Guided by Theorem 4.1, we implement the label predictor into two components:
+$f({\mathbf{z}})\;=\;f_{b}({\mathbf{b}})\;+f_{inv}({\mathbf{c}},Pr),$
+where $f_{b}$ denotes the bias-aware predictor $p(y\mid{\mathbf{b}})$, and $f_{inv}$ is the invariant predictor built on the content variable $\mathbf{c}$} and the learnable prior term $Pr=\operatorname{logit}\!\big(p(y)\big)$.
+In detail, $f_{in}({\mathbf{c}},Pr)=f_{c}({\mathbf{c}})-Pr$ where $f_{c}({\mathbf{c}})=p(y\mid{\mathbf{c}})$.
+In particular, $f_{inv}$ is an “invariant” module across all domains and can be used directly in the target domain without further modification, even under label shift.
+In training, we jointly learn bias predictors, the content predictor, and the logits prior with a cross-entropy classification loss $\mathcal{L}_{\text{cls}}$.
+Theorem 4.2 (Upper bound of Bias Correction).
+Under the data generation process in Figure 2, define the corrected bias predictor $\widehat{f}({\mathbf{b}})\;=\;\frac{\widehat{q}({\mathbf{b}})+\hat{h}_{0}-1}{\,\hat{h}_{0}+\hat{h}_{1}-1\,}$, where $\widehat{q}({\mathbf{b}})$ is the estimator of $q({\mathbf{b}})=p(\hat{y}=1\mid{\mathbf{b}})$, $\hat{h}_{0},\hat{h}_{1}$ are estimators of $h_{0}$ and $h_{1}$.
+Let $f^{\star}({\mathbf{b}})=p(y=1\mid{\mathbf{b}})$ mean using only the $\mathbf{b}$} learned on sources to predict $y$ at the test stage.
+For the target cross-entropy risk on the bias subproblem, define: $\mathcal{R}_{(b)}(g)=\mathbb{E}_{{\mathbf{b}}\sim p({\mathbf{b}})}\Big(\mathrm{KL}\big(\mathrm{Bern}(p(y=1\mid{\mathbf{b}}))\,\|\,\mathrm{Bern}(g({\mathbf{b}}))\big)\Big).$ Then there exist constants $c_{1},c_{2}>0$ (depending only on the margin assumption) such that:
+$\mathcal{R}_{(b)}(\widehat{f})\;\leq\;\frac{c_{1}}{(h_{0}+h_{1}-1)^{2}}\,\mathbb{E}\!\big[(\widehat{q}({\mathbf{b}})-q({\mathbf{b}}))^{2}\big]\;+\;$ (6)
+$\frac{c_{2}}{(h_{0}+h_{1}-1)^{2}}\,\Big((\hat{h}_{0}-h_{0})^{2}+(\hat{h}_{1}-h_{1})^{2}\Big)$
+In particular, if this upper bound is smaller than $\mathcal{R}_{(b)}(f^{\star})$ (equivalently, $\Delta_{\mathrm{DG}}$), then $\widehat{f}$ attains strictly lower risk on the target domain than $f^{\star}$.
+Theorem 4.2 shows that while correction incurs small statistical costs from estimating $q({\mathbf{b}})$ and ($h_{0}$$h_{1}$), these diminish with data.
+So under the data generation process in Figure 2, the corrected predictor is provably more reliable than source-only deployment (see Section B.7 for details).
+Bias adaptation.
+In the training phase, the content predictor $f_{c}$ and the learnable label prior $Pr$ are jointly optimized to form the invariant component, which is then frozen for the test phase.
+During test time, given test-domain inputs $\{{\mathbf{x}}^{\mathcal{T}}\}$, we first extract the latent representations $\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}]$ using the pre-trained encoder.
+We then leverage the learned content predictor $f_{c}(\mathbf{c}^{\mathcal{T}})$ to generate pseudo-labels $\hat{y}^{\mathcal{T}}$ on unlabeled test data.
+In the source domains, we collect $h_{0}$ and $h_{1}$, which are entries of the confusion-matrix probabilities between pseudo-labels and ground truth.
+Intuitively, these parameters adjust how domain-invariant information maps the bias $\mathbf{b}$ to valid probability estimates in the new domain.
+To estimate $p(\hat{y}=1\mid\mathbf{b})$, we update the bias-aware predictor $f_{b}$ with the generated pseudo-labels $\hat{y}^{\mathcal{T}}$ as
 […]
-Theorem 2.5 (Correction with Pseudo Labels).
-Consider the predictor $f_{c}:{\mathbf{c}}\rightarrow y$.
-Define pseudo-label prediction as $\hat{y}=f_{c}({\mathbf{c}})$.
-Let $h_{0}\;=\;p(\hat{y}=0\mid y=0),\quad h_{1}\;=\;p(\hat{y}=1\mid y=1).$ Then it holds that
+where the adaptation loss $\mathcal{L}_{ada}$ is characterized as the cross-entropy loss with pseudo labels.
+We record this post-trained predictor as $\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}})$.
+Let $\phi(\cdot)=\operatorname{logit}((\frac{\sigma(\cdot)+h_{0}-1}{h_{0}+h_{1}-1}))$ represent the correction function.
+The final prediction is
+$f([\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}])=f_{c}(\mathbf{c}^{\mathcal{T}})+\phi(\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}}))-Pr.$
+Objectives.
+Overall, the losses in both the training and bias adaptation stages are as follows:
+$\text{Training Stage: }\mathcal{L}_{all}=\mathcal{L}_{cls}+\lambda_{0}\mathcal{L}_{\text{vae}}+\lambda_{1}\mathcal{L}_{ind}$
+$\text{Adaptation Stage: }\mathcal{L}_{ada}$
+where $\lambda_{0}$, $\lambda_{1}$ are the hyperparameters.
+5 Experiments
+In this section, we test BAG on both synthetic data and real-world datasets, following the setting in the main baseline methods ACTIR (Jiang and Veitch, 2022) and SFB (Eastwood et al., 2023), for a fair comparison.
+Baseline methods.
+We compare our method with ERM and IRM  (Arjovsky et al., 2019), classic approaches that leverage invariant information for making predictions.
+Additionally, we compare with ACTIR  (Jiang and Veitch, 2022), which disentangles the invariant feature and bias through independent regularization.
+We also compare with SFB  (Eastwood et al., 2023), which aims to exploit spurious features but in a weaker setting.
+SFB can serve as a special case of our method when 1) the environment doesn’t affect the label and 2) it is not bias-aware for prediction.
+We implemented the SFB algorithm and used it for subsequent testing.
+We also include RDM (Nguyen et al., 2024) and GMDG (Tan et al., 2024) as more recent and advanced baselines.
+5.1 Simulation Experiments
+Data simulation.
+To examine whether BAG can effectively identify and leverage domain shift information for better generalization, we devised a series of simulation experiments.
+Our data generation process strictly adheres to the causal structure depicted in  Figure 2.
+First, we sample an environmental variable $e$ following a categorical distribution, $e\sim\mathrm{Categorical}(\{\pi_{1},\dots,\pi_{M}\})$, which subsequently induces variations across different domains.
+Then, a binary label $y$ is generated in an environment-dependent manner: $y=\mathbf{1}\{\,w^{\mathsf{T}}E+b_{0}>0\},$ where $E$ denotes an embedding of $e$.
+Next, we assign stable content $c$ based on the value of $y$: if $y=0$, then $c=c_{0}+\epsilon$; if $y=1$, then $c=c_{1}+\epsilon$.
+Here, $\epsilon\sim\mathcal{N}(0,\sigma_{c}^{2}I)$.
+Simultaneously, the bias variable $b$ is determined as: $b=E+\mathbf{C}_{e,y},$ ensuring that the background shift is influenced jointly by the environment and the label.
+This setup emulates the common real-world scenario where false associations often arise.
+Finally, stable content and bias jointly generate the observed data: $x=M\begin{bmatrix}c\\ b\end{bmatrix}+\eta,$ where $M$ is an invertible matrix, and $\eta\sim\mathcal{N}(0,\sigma_{x}^{2}I)$ represents measurement noise.
+Figure 4: Mean and standard deviation across different methods in binary classification simulation experiments.
+Results and discussions.
+This study evaluates several methods on a synthetic dataset of 5000 samples.
+To ensure result reliability, each method was repeated five times, and the average performance was recorded.
+As shown in Figure 4, both ERM and IRM performed well on the training set but achieved test accuracies just above 50%.
+We attribute this to the significant distributional disparity of environmental information $e$ across domains in the dataset, which limits the effectiveness of the class label in supporting accurate predictions.
+In other words, an excessive focus on invariant information may hinder model optimization, leading to marginal performance gains.
+In contrast, our method leverages domain specific information ($\mathbf{b}$}) more effectively than ACTIR or SFB, significantly enhancing prediction performance by reducing noise relative to relying solely on the class label.
+5.2 Real-world Experiments
+Experimental settings.
+We evaluate our method on three widely used domain generalization benchmarks: PACS (Li et al., 2017), Office-Home (Venkateswara et al., 2017), and DomainNet (Peng et al., 2019).
+For each dataset, we adopt the leave one domain out protocol, treating one domain as the target and the remaining domains as sources.
+We use ResNet-18 as the backbone on PACS (following SFB) and ResNet-50 on Office-Home and DomainNet (following RDM).
+All models are equipped with MLP-based VAEs and classifiers.
+We report results averaged over three random seeds.
+Our framework uses a single-layer linear encoder and a two-layer linear decoder.
+Each domain-specific classifier is implemented as a single-layer linear head.
+We parameterize the environment embedding with three learnable vectors.
+Further implementation details are provided in Section E.1.
+Table 1: Results on PACS (ResNet-18) and Office-Home (ResNet-50). BAG results are averaged over 3 seeds. ACTIR and SFB results are taken from (Eastwood et al., 2023).
+[…]
+Table 2: Domain-specific OOD accuracy on DomainNet. All methods use ResNet-50 as backbones.
+Algorithm Clipart Infograph Painting Quickdraw Real Sketch Avg ERM 58.1 18.8 46.7 12.2 59.6 49.8 40.9 IRM(Arjovsky et al., 2019) 48.5 15.0 38.3 10.9 48.2 42.3 33.9 ACTIR (Jiang and Veitch, 2022) 50.0 22.6 43.5 11.7 57.8 46.8 38.7 RDM (Nguyen et al., 2024) 62.1 20.7 49.2 14.1 63.0 51.4 43.4 GMDG (Tan et al., 2024) 63.4 22.4 51.4 13.4 64.4 52.4 44.6 SFB (Eastwood et al., 2023) 60.7 25.3 49.0 12.0 64.0 50.2 43.5 BAG 61.8 25.6 51.8 13.9 65.5 50.3 44.8
+Results and discussions.
+For PACS (left in Table 1), our method achieves the best average accuracy, improving by $1.9\%$ over the strongest baseline, GMDG.
+Compared with the most closely related method, SFB, our model consistently improves performance across all target domains.
+The largest gain is observed on Art Painting, possibly because this domain contains stronger style/background shifts, where modeling bias is particularly beneficial.
+On Office-Home (right in Table 1), BAG outperforms all baselines on average.
+In particular, it achieves improvement on three of the domains, and obtains comparable results on Arts.
+The notably larger improvement on Clipart which is generally regarded as more challenging, suggests that BAG generalizes better under larger domain shifts.
+On DomainNet, BAG achieves the best average accuracy, surpassing GMDG and RDM.
+It attains the best performance on infograph, painting and real, while remaining competitive on clipart , quickdraw, and sketch.
+The largest gain appears on infograph, where background and stylistic artifacts are prominent, indicating that disentangling content from bias helps suppress spurious domain cues and better exploit informative bias factors.
+We also outperform ACTIR and SFB by a substantial margin, which we attribute to more accurate disentanglement and a more principled use of bias.
+5.3 More Ablation Studies, Sensitivity Check, and Visualization.
+Since BAG involves a two-stage optimization procedure and a VAE component, we conduct some ablation studies to quantify the contributions of each design choice.
+We remove the corresponding components, such as VAE, test-time adaptation that uses pseudo-labels to guide the bias branch, and the reweighting-of-experts module while keeping the rest unchanged as ablation studies.
+Full ablation details are provided in Section E.2.
+To verify the stability of BAG, we study its sensitivity to the number of domain experts and the embedding dimensionality in Section E.3.
+We also visualize the learned representations to illustrate the disentanglement of latent variables in Section E.4.
+6 Conclusion
+In this paper, we challenge the conventional methods for OOD generalization that treat bias solely as an obstacle.
+Instead, we theoretically demonstrate that bias can be strategically leveraged to enhance prediction, particularly when it retains useful dependencies on the target domain labels.
+Our framework demonstrates that bias can be utilized in two ways: bias-aware prediction with an environment routing mechanism and bias correction with an adaptive label prior.
+Through empirical validation on a synthetic task and some real-world benchmarks, we establish the effectiveness of the proposed approach.
+The results consistently show that leveraging bias in a structured manner leads to more robust and adaptable models, offering a new perspective on OOD generalization.
+Limitations.
+This paper provides a principled general framework that leverages bias for domain generalization.
+From a theoretical perspective, the success of such a framework relies on assumptions about the disentanglement of bias features and invariant features.
+Although the disentanglement is supported by both real-world and simulated experiments, it typically depends on intrinsic structural properties of the data (e.g., linear independence) that are difficult to verify in practice.
+We therefore recommend users to conduct independence tests between bias and invariant features or the visualizations on the feature space as a practical check metric to help assess their distinguishability in real applications.
+Impact Statement
+We propose a method that strengthens unsupervised domain generalization, enabling models to transfer knowledge more reliably between different data domains.
+The approach is designed to improve learning efficiency when annotations are limited or unavailable, which is common in real world deployments.
+The technique is methodological and does not create new ethical concerns beyond those already associated with standard machine learning systems.
+Overall, our contribution supports more robust and practical domain generalizing models.
+Appendix A Related Work
+Domain Generalization Domain Generalization (DG) aims to develop models that generalize to unseen OOD target domains using only observed source domains for training (Zou et al., 2024; Zhang et al., 2024; Chen et al., 2023; Prashant et al., 2024).
+A common approach is learning invariant representations across all source domains, as seen in Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) and its extensions (Ahuja et al., 2020; Rosenfeld et al., 2020; Ahuja et al., 2021; Kamath et al., 2021; Krueger et al., 2021).
+Other methods leverage adversarial learning (Li et al., 2018a; Shao et al., 2019; Zhao et al., 2020; Deng et al., 2020; Matsuura and Harada, 2020) to extract domain-invariant features or use disentangled representation learning (Ilse et al., 2020; Wang et al., 2020b; Jiang and Veitch, 2022) to separate domain-invariant from domain-specific information.
+Causal inference techniques (Wang et al., 2020c; Qi et al., 2020), such as back-door and front-door rules, help mitigate bias.
+Furthermore, techniques such as data augmentation (Volpi and Murino, 2019; Shi et al., 2020; Huang et al., 2021; Li et al., 2021), feature augmentation (Mancini et al., 2020; Shu et al., 2021; Zhou et al., 2024), and network regularization (Seo et al., 2020; Segu et al., 2023) can reduce overfitting and improve generalization.
+Causal Representation Learning Causal Representation Learning (CRL) aims to recover latent variables that correspond to the causal generative factors of high-dimensional observations and the mechanisms that link them (Schölkopf et al., 2021).
+Building on the invariance principle, Invariant Causal Prediction formalizes causal variable selection via stability across environments (Peters et al., 2016).
+A rich line of identifiability results shows that with auxiliary variables or weak interventions, nonlinear ICA and VAE-based approaches can recover latent sources up to permissible transformations (Hyvarinen et al., 2019; Khemakhem et al., 2020).
+For vision, structured VAEs with causal priors explicitly separate content from style/context (Yang et al., 2021), and temporal–interventional setups such as Citris leverage intervention sequences to learn identifiable factors (Lippe et al., 2022).
+Beyond model proposals, empirical and theoretical analyses caution that purely unsupervised disentanglement is generally impossible without inductive biases or weak supervision, motivating the use of environment cues or metadata (Locatello et al., 2019).
+One line of research leverages distributional shifts in latent variables, where nonstationary or environmental variation facilitates the recovery of nonlinear independent components and the underlying causal structures (Kong et al., 2022; Zhang and Hyvärinen, 2011; Chen et al., 2024).
+Our approach follows this CRL perspective: we infer a latent space with a VAE, factorize it into invariant content and bias/context, then exploit the bias variables for environment-aware prediction while aligning them with invariant predictions.
+Label shift in OOD problems Label shift occurs when the label prior $p(y)$ changes between training and test domains while the class-conditional $p(x\mid y)$ remains stable (Lipton et al., 2018).
+Classical corrections include EM-based prior adjustment (Saerens et al., 2002) and Black-Box Shift Estimation (BBSE), which uses a calibrated source classifier’s confusion matrix to estimate target priors and reweight predictions (Lipton et al., 2018).
+Subsequent work unifies moment-matching and maximum-likelihood estimators and emphasizes the central role of calibration (e.g., bias-corrected temperature scaling, BCTS) for reliable prior estimation (Garg et al., 2020; Azizzadenesheli et al., 2019).
+In streaming or nonstationary settings, online label-shift adaptation updates class-prior estimates on the fly without access to target labels (Wu et al., 2021; Bai et al., 2022).
+Benchmarks such as WILDS and Wild-Time document that label shift frequently co-occurs with input shift in real applications (e.g., medical imaging, wildlife monitoring, and temporally evolving text) (Koh et al., 2021; Yao et al., 2022).
+Consistent with this evidence, we explicitly account for changes in $p(y)$ when adapting to new environments, combining prior correction with environment-aware mixture-of-experts weighting informed by the biased latent variables.
+Unsupervised domain adaptation Unlike DG, unsupervised domain adaptation utilizes unlabeled target data to guide model adaptation.
+Due to the observation of target distribution, many methods focus on distribution alignment by minimizing domain divergence (Rozantsev et al., 2018; Sun and Saenko, 2016; Kang et al., 2019; Liu et al., 2020), using importance reweighting (Jiang and Zhai, 2007; Xu et al., 2019; Fang et al., 2020), and learning with adversarial discriminators, such as DANN (Ganin et al., 2016), ADDA (Tzeng et al., 2017), and WDDA (Shen et al., 2018).
+Beyond distribution alignment, some methods focus on invariant information-based adaptation, including self-training (Lee and others, 2013; Zou et al., 2018; Saito et al., 2020; Chen et al., 2020; Kumar et al., 2020; Liu et al., 2021a), which uses pseudo-labels from a source-trained model to guide target domain learning, and batch normalization-based regularization (Li et al., 2018b; Liu et al., 2021b), which adapts only BN layers while keeping other parameters invariant.
+Recently, disentangled representation learning (Kong et al., 2022; Li et al., 2023b) has shown strong potential for adaptation by separating domain-invariant content from domain-specific bias.
+Test-Time domain adaptation In real-world scenarios, the target domain distribution is often unknown in advance, making it challenging to train an adaptive model beforehand, which motivates test time adaptation.
+The typical solution is Test-Time Training (TTT) (Sun et al., 2020; Liu et al., 2021c; Mirza et al., 2023; Li et al., 2023a), which adapts models during inference by optimizing the same self-supervised learning objective used during pre-training.
+Besides, even without access to the target distribution, adaptation remains feasible with invariant information learned from the source domain, such as through self-training (Chen et al., 2022; Tomar et al., 2023; Su et al., 2022; Goyal et al., 2022) or network regularization (Wang et al., 2020a; Liang et al., 2020).
+The most closely related work to ours is Stable Feature Boosting (SFB), which leverages unstable features alongside invariant representations to generate pseudo-labels when only the image is affected by environmental factors.
+Our approach differs in three key ways.
+First, we delve deeper into latent variable modeling to provide a more comprehensive explanation of bias.
+Second, our model accommodates more complex bias scenarios, allowing the environment variable to influence both the image and the label.
+Third, we integrate invariant correction and expert reweighting to effectively utilize biased data.
+Appendix B Proof and Discussion of section 2
+B.1 Proof of the identification of latent variables
+Lemma B.1.
+[…]
+  • A1 (Smooth and positive density): $p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})$ is smooth with strictly positive density on its support.
+  • A2 (Conditional independence and content invariance): Given $(\mathbf{e},\mathbf{y})$, the coordinates of $\mathbf{z}$ are independent,
+    $\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i};\mathbf{e},\mathbf{y}),$
+    where $n=n_{b}+n_{c}$, $z_{1},\ldots,z_{n_{b}}$ correspond to $\mathbf{b}$ and $z_{n_{b}+1},\ldots,z_{n}$ correspond to $\mathbf{c}$.
+Moreover, content coordinates are environment-invariant: for $i>n_{b}$, $q_{i}(c_{i-n_{b}};\mathbf{e},\mathbf{y})=q_{i}(c_{i-n_{b}};\mathbf{y})$.
+  • A3 (Linear independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$ (and the relevant $\mathbf{y}$), there exist $n_{b}+1$ environments $e_{0},e_{1},\ldots,e_{n_{b}}$ such that the $n_{b}$ vectors
+    $\bm{w}(\mathbf{b},e_{j},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y}),\quad j=1,\ldots,n_{b},$
+    are linearly independent, where
+    $\bm{w}(\mathbf{b},e_{i},\mathbf{y}):=\Big(\tfrac{\partial q_{1}(b_{1};e_{i},\mathbf{y})}{\partial b_{1}},\ldots,\tfrac{\partial q_{n_{b}}(b_{n_{b}};e_{i},\mathbf{y})}{\partial b_{n_{b}}}\Big)^{\top}.$
+  • A4 (Domain variability): There exist $e_{i}\neq e_{j}$ such that, for the relevant $\mathbf{y}$ and for any measurable set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability that cannot be written as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$,
+    $\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{i},\mathbf{y})\,d\mathbf{z}\;\neq\;\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{j},\mathbf{y})\,d\mathbf{z}.$
+[…]
+Proof.
+We start from the matched conditional distribution $p_{\hat{\mathbf{x}}\mid\mathbf{e},\mathbf{y}}=p_{\mathbf{x}\mid\mathbf{e},\mathbf{y}}$ and introduce the diffeomorphism
+$h:=g^{-1}\circ\hat{g}:\;\hat{\mathbf{z}}\mapsto\mathbf{z}.$
+Then, for all $(\mathbf{e},\mathbf{y})$,
+$p_{\hat{g}(\hat{\mathbf{z}})\mid\mathbf{e},\mathbf{y}}=p_{g(\mathbf{z})\mid\mathbf{e},\mathbf{y}}.$ (8)
+By the change-of-variables formula,
 […]
 […]
-Justification about Theorem 2.5.
-Equation 3 depends and only depends on the 2.4.
-The detailed proof can be found in Appendix D.2.
-Intuitively, if $c$ helps distinguish between $y=0$ and $y=1$, then a predictor based on $c$ should be corrected more often than chance, implying $h_{0}+h_{1}>1$.
-Proof of Equation 4.
+By A2 (conditional independence), letting $q_{i}(z_{i};\mathbf{e},\mathbf{y}):=\log p_{z_{i}\mid\mathbf{e},\mathbf{y}}(z_{i})$ and $\hat{q}_{i}(\hat{z}_{i};\mathbf{e},\mathbf{y}):=\log p_{\hat{z}_{i}\mid\mathbf{e},\mathbf{y}}(\hat{z}_{i})$, we have
+$\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i};\mathbf{e},\mathbf{y}),\quad\log p_{\hat{\mathbf{z}}\mid\mathbf{e},\mathbf{y}}(\hat{\mathbf{z}}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i};\mathbf{e},\mathbf{y}).$ (10)
+[…]
+$\sum_{i=1}^{n}q_{i}\big(z_{i};\mathbf{e},\mathbf{y}\big)\;+\;\log\big|\det\mathbf{J}_{h}(\hat{\mathbf{z}})\big|\;=\;\sum_{i=1}^{n}\hat{q}_{i}\big(\hat{z}_{i};\mathbf{e},\mathbf{y}\big).$ (11)
+Step 1: $\mathbf{B}=\partial\mathbf{b}/\partial\hat{\mathbf{c}}=\mathbf{0}$.
+Differentiate equation 11 with respect to $\hat{c}_{j}$ for any $j\in\{1,\ldots,n_{c}\}$:
+$\sum_{i=1}^{n}\frac{\partial q_{i}(z_{i};\mathbf{e},\mathbf{y})}{\partial z_{i}}\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}\;+\;\frac{\partial}{\partial\hat{c}_{j}}\log\big|\det\mathbf{J}_{h}(\hat{\mathbf{z}})\big|\;=\;\frac{\partial\hat{q}_{j}(\hat{c}_{j};\mathbf{e},\mathbf{y})}{\partial\hat{c}_{j}}.$ (12)
+Note that $\mathbf{J}_{h}$ does not depend on $\mathbf{e}$, since $h=g^{-1}\circ\hat{g}$ is a mapping between latent spaces only.
+Fix $e_{0},\ldots,e_{n_{b}}$ as in A3 and subtract the version of equation 12 at $e_{0}$ from that at $e_{k}$ ($k=1,\ldots,n_{b}$); the Jacobian term cancels:
+$\sum_{i=1}^{n}\!\left(\frac{\partial q_{i}(z_{i};e_{k},\mathbf{y})}{\partial z_{i}}-\frac{\partial q_{i}(z_{i};e_{0},\mathbf{y})}{\partial z_{i}}\right)\!\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}\;=\;\frac{\partial\hat{q}_{j}(\hat{c}_{j};e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j};e_{0},\mathbf{y})}{\partial\hat{c}_{j}}.$ (13)
+By A2, for content coordinates ($i>n_{b}$) the difference on the left is zero because $q_{i}$ does not depend on $\mathbf{e}$.
+By the definition of the learned invariant block, $p_{\hat{c}_{j}\mid\mathbf{e},\mathbf{y}}$ is invariant in $\mathbf{e}$, hence the right-hand side of equation 13 is zero.
+Therefore,
+$\sum_{i=1}^{n_{b}}\!\left(\frac{\partial q_{i}(b_{i};e_{k},\mathbf{y})}{\partial b_{i}}-\frac{\partial q_{i}(b_{i};e_{0},\mathbf{y})}{\partial b_{i}}\right)\!\cdot\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0,\quad k=1,\ldots,n_{b}.$
+By A3, the $n_{b}$ vectors $\bm{w}(\mathbf{b},e_{k},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y})$ are linearly independent, so the only solution is
+$\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0,\qquad\forall\,i\in\{1,\ldots,n_{b}\},\;j\in\{1,\ldots,n_{c}\}.$
+Let the Jacobian of $h$ be block-partitioned as
+$\mathbf{J}_{h}=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix}=\begin{bmatrix}\partial\mathbf{b}/\partial\hat{\mathbf{b}}&\partial\mathbf{b}/\partial\hat{\mathbf{c}}\\ \partial\mathbf{c}/\partial\hat{\mathbf{b}}&\partial\mathbf{c}/\partial\hat{\mathbf{c}}\end{bmatrix}$.
+[…]
+Step 2: $\mathbf{C}=\partial\mathbf{c}/\partial\hat{\mathbf{b}}=\mathbf{0}$.
+[…]
+[…]
+B.2 Discussion of Lemma 2.1
+Concepts Explanation
+We would like to detailedly explain some concepts in Lemma 2.1 as follows: (1)Here “Positive support” means the set of all points where a probability distribution assigns strictly positive probability or density (i.e., $p(x)>0$).
+This is emphasized because the probability distribution may have some point in 0 for example the distribution are uniform distribution.
+To maintain academic rigor, this hypothesis was highlighted.
+(2)The $\mathcal{B}$ is introduced in A3 and it refers to the set (or space) of domain-specific latent variables.
+(3)Linear independence means that changes in different environments produce distinct, non-redundant effects on the domain-specific latent variable $b$; none of these variations can be expressed as a combination of others.
+This is a reasonable assumption because, with multiple diverse environments, it ensures that each domain provides unique information about how $b$ varies, making the latent factors identifiable rather than entangled.
+The detailed explanations of this lemma can be found in paper (Kong et al., 2022).
+Discussion of Assumptions.
+(A1) Smooth and positive density.
+This assumption requires that the conditional densities $p({\mathbf{z}}\mid e,y)$ are smooth and strictly positive on their supports.
+Intuitively, it ensures that the latent space is continuous and that every possible latent configuration has a non-zero probability of occurrence.
+In practice, this assumption is easily satisfied: with sufficient data, empirical or neural-network-based density estimations naturally produce smooth and positive densities.
+Since our framework adopts a VAE-based generative model, the latent distribution is modeled by Gaussian families whose densities are positive everywhere, satisfying this assumption automatically.
+However, this assumption may fail when the latent space is discrete or piecewise constant, such as in deterministic autoencoders or models that collapse the latent variance to zero.
+In those degenerate cases, $p({\mathbf{z}}\mid e,y)$ no longer has a smooth positive density, and the identifiability conditions in Lemma 2.1 no longer hold rigorously.
+(A2) Conditional independence and content invariance.
+This assumption states that the coordinates of the latent variables ${\mathbf{z}}=(z_{1},\dots,z_{n})$ are conditionally independent given $e$ and $y$, i.e., $\log p({\mathbf{z}}\mid e,y)=\sum_{i}q_{i}(z_{i};e,y)$.
+Conceptually, this means that each latent dimension corresponds to an independent generative factor once the environment and label are specified.
+It also implies that content variables are invariant across environments—only the bias-related coordinates are affected by domain change.
+This assumption is reasonable in our framework because, when the generative process is disentangled and the noise level is small, the variation of $\mathbf{z}$} can indeed be attributed to independent factors conditioned on $e$ and $y$.
+Moreover, the VAE objective encourages conditional independence through the factorization of the approximate posterior.
+However, this assumption may be violated when latent variables interact strongly (e.g., through nonlinear dependencies) or when noise in the data introduces correlated variations between coordinates.
+In such cases, exact independence cannot be guaranteed, but approximate independence is often sufficient for stable learning and representation disentanglement.
+(A3) Linear independence.
+This assumption imposes a technical requirement on the partial derivatives of the latent log-density functions with respect to environment changes.
+Intuitively, this means that the changes introduced by different environments are sufficiently diverse so that the bias components can be linearly separated from one another.
+In practice, this condition cannot be enforced directly, but it is often satisfied when multiple environments exist and differ in a sufficiently rich manner.
+Empirically, in synthetic datasets, when the number of environments is large, linear independence almost always holds, guaranteeing successful disentanglement.
+In real-world data, this condition may be partially violated because the number of distinct environments if we think this is domain number is limited.
+Nevertheless, prior works (Kong et al., 2022; Li et al., 2023b) and our empirical results show that even approximate linear independence is sufficient to achieve robust representations.
+This is may because the so-called domain in domain generalization are large domain and in this domain that there existing small domains to satisfy the assumptions.
+(A4) Domain variability.
+This assumption requires that, for at least one pair of distinct environments $e_{i}\neq e_{j}$ and the corresponding class $y$, there exists a measurable subset $A_{z}\subset\mathcal{Z}$ with non-zero probability that cannot be expressed as a purely content-aligned region $\Omega_{c}\times B$, such that
+$\int_{A_{z}}p({\mathbf{z}}\mid e_{i},y)\,dz\;\neq\;\int_{A_{z}}p({\mathbf{z}}\mid e_{j},y)\,dz.$
+This ensures that the conditional latent distribution genuinely varies across domains, capturing environment-dependent factors beyond the content component.
+Because the mapping ${\mathbf{x}}=g({\mathbf{z}})$ is deterministic and noiseless, $p({\mathbf{x}}\mid e,y)$ is the push forward of $p({\mathbf{z}}\mid e,y)$; therefore, if $p({\mathbf{z}}\mid e_{i},y)=p({\mathbf{z}}\mid e_{j},y)$, we would also have $p({\mathbf{x}}\mid e_{i},y)=p({\mathbf{x}}\mid e_{j},y)$.
+However, this equality does not hold in domain generalization, where environments are explicitly assumed to induce observable shifts.
+To see this, consider that if the joint distributions were identical across environments,
+$p({\mathbf{x}},y,e_{i})=p({\mathbf{x}},y,e_{j}),$
+then by the chain rule,
+$p({\mathbf{x}},y,e)=p({\mathbf{x}}\mid y,e)\,p(y\mid e)\,p(e).$
+Integrating both sides over a measurable subset $A_{x}$ gives
+$\int_{A_{x}}p({\mathbf{x}}\mid y,e_{i})\,p(y\mid e_{i})\,dx=\int_{A_{x}}p({\mathbf{x}}\mid y,e_{j})\,p(y\mid e_{j})\,dx.$
+If $p({\mathbf{x}}\mid y,e_{i})=p({\mathbf{x}}\mid y,e_{j})$ held for all $A_{x}$, then equality would require $p(y\mid e_{i})=p(y\mid e_{j})$.
+Yet, in our setting, we explicitly consider label shift, meaning $p(y\mid e)$ varies with $e$; hence, this equality cannot generally hold.
+Consequently, A4 guarantees genuine domain variability at the level of $p({\mathbf{z}}\mid e,y)$, which propagates to $p({\mathbf{x}}\mid e,y)$ through the deterministic mapping.
+Nevertheless, A4 may fail under degenerate or unrealistic conditions, such as:
+  1. (i)
+    when environments do not induce any shift in latent factors ($p({\mathbf{z}}\mid e_{i},y)=p({\mathbf{z}}\mid e_{j},y)$ for all $i,j$), making domains statistically indistinguishable;
+  2. (ii)
+    when the mapping $g$ is stochastic or non-injective, collapsing different latent codes into identical observations, which hides latent variability in the observed space;
+  3. (iii)
+    or when label shift is absent ($p(y\mid e_{i})=p(y\mid e_{j})$), eliminating one key source of distributional difference across domains.
+These cases violate the core assumption that domains differ in meaningful, non-content-related ways and thus fall outside our framework.
+B.3 Better predictor with adding bias
+Proof.
+Under Assumptions A1–A4, we compare the Bayes-optimal risks for predictors based on $c$ versus $(c,b)$.
+Log‐loss.
+The Bayes-optimal log-loss risks satisfy
+$R_{\log}\bigl(q^{*}(\cdot\mid c)\bigr)=H(Y\mid C),\qquad R_{\log}\bigl(q^{*}(\cdot\mid c,b)\bigr)=H(Y\mid C,B),$
+where
+$H(Y\mid C)=-\mathbb{E}_{C}\bigl[\mathbb{E}_{Y\mid C}[\log p(Y\mid C)]\bigr],\quad H(Y\mid C,B)=-\mathbb{E}_{C,B}\bigl[\mathbb{E}_{Y\mid C,B}[\log p(Y\mid C,B)]\bigr].$
+There is a positive-measure set of $(c,b)$ on which $p(y\mid c,b)\neq p(y\mid c)$. Hence
+$I(Y;B\mid C)=H(Y\mid C)-H(Y\mid C,B)=\mathbb{E}_{C,B}\!\bigl[D_{\mathrm{KL}}(p(y\mid c,b)\,\|\,p(y\mid c))\bigr]>0,$
+so $H(Y\mid C,B)<H(Y\mid C)$ and therefore
+$R_{\log}\bigl(q^{*}(\cdot\mid c,b)\bigr)<R_{\log}\bigl(q^{*}(\cdot\mid c)\bigr).$
+Squared‐error loss.
+The Bayes‐optimal squared‐error risks satisfy
+$R_{\mathrm{sq}}\bigl(f^{*}(c)\bigr)=\mathbb{E}\bigl[\mathrm{Var}(Y\mid C)\bigr],\qquad R_{\mathrm{sq}}\bigl(f^{*}(c,b)\bigr)=\mathbb{E}\bigl[\mathrm{Var}(Y\mid C,B)\bigr].$
+By the law of total variance,
+$\mathrm{Var}(Y\mid C)=\mathbb{E}_{B}\bigl[\mathrm{Var}(Y\mid C,B)\bigr]\;+\;\mathrm{Var}_{B}\!\bigl(\mathbb{E}[Y\mid C,B]\bigm|C\bigr).$
+Since $B$ influences the conditional mean of $Y$ given $C$, the second term is strictly positive on a set of positive probability. Thus
+$\mathbb{E}\bigl[\mathrm{Var}(Y\mid C,B)\bigr]<\mathbb{E}\bigl[\mathrm{Var}(Y\mid C)\bigr],$
+and consequently
+$R_{\mathrm{sq}}\bigl(f^{*}(c,b)\bigr)<R_{\mathrm{sq}}\bigl(f^{*}(c)\bigr).$
+This completes the proof that incorporating $b$ strictly reduces both log‐loss and squared‐error Bayes risks. ∎
+B.4 The derivation of the KL optimization
+We want to solve the following problem:
+$q^{*}(y\mid{\mathbf{b}})\;=\;\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl[D_{\mathrm{KL}}\bigl(p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr)\Bigr].$
+Let
+$\mathcal{L}[q(\cdot\mid{\mathbf{b}})]\;=\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl[D_{\mathrm{KL}}\bigl(p(y\mid{\mathbf{b}},E)\,\|\,q(y\mid{\mathbf{b}})\bigr)\Bigr].$
+By the definition of KL divergence, for discrete $y$:
+$D_{\mathrm{KL}}\bigl(P(y)\,\|\,Q(y)\bigr)=\sum_{y}P(y)\,\log\frac{P(y)}{Q(y)}.$
+Hence, for each $e$,
+$D_{\mathrm{KL}}\bigl(p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr)=\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
+Therefore,
+$\mathcal{L}[q]=\sum_{e}p(e\mid{\mathbf{b}})\,\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
+We want to choose $q(y\mid{\mathbf{b}})$ to minimize $\mathcal{L}[q]$, which needs $q(y\mid{\mathbf{b}})$ to satisfy $\sum_{y}q(y\mid{\mathbf{b}})\;=\;1,\text{for each fixed }{\mathbf{b}}.$ To incorporate this constraint, we introduce a Lagrange multiplier $\lambda$ and consider the Lagrangian:
+$\mathcal{J}[q,\lambda]=\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\!\biggl[\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\biggr]\;+\;\lambda\Bigl(\sum_{y}q(y\mid{\mathbf{b}})-1\Bigr).$
+To find the minimizing $q$, we set
+$\frac{\partial\mathcal{J}}{\partial q(y\mid{\mathbf{b}})}\;=\;0.$
+First, note
+$\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\;=\;\log p(y\mid{\mathbf{b}},e)-\log q(y\mid{\mathbf{b}}).$
+Only the term $\log q(y\mid{\mathbf{b}})$ depends on $q$. For a specific $y$,
+$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\bigl[-\log q(y^{\prime}\mid{\mathbf{b}})\bigr]=\begin{cases}-\frac{1}{q(y\mid{\mathbf{b}})},&\text{if }y^{\prime}=y,\\ 0,&\text{if }y^{\prime}\neq y.\end{cases}$
+Hence, for each $e$:
+$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\biggl[\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}\biggr]=-\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
+Summing over $e$,
+$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
+Also,
+$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\Bigl[\lambda\bigl(\sum_{y^{\prime}}q(y^{\prime}\mid{\mathbf{b}})-1\bigr)\Bigr]=\lambda.$
+Hence,
+$\frac{\partial\mathcal{J}}{\partial q(y\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda.$
+Setting this to zero:
+$\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda=0,$
+which gives
+$\frac{1}{q(y\mid{\mathbf{b}})}\,\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\lambda.$
+Thus
+$q(y\mid{\mathbf{b}})=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
+We require $\sum_{y}q(y\mid{\mathbf{b}})=1$. Then:
+$1=\sum_{y}q(y\mid{\mathbf{b}})=\sum_{y}\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e).$
+Since $\sum_{y}p(y\mid{\mathbf{b}},e)=1$ for each $e$, we get
+$\sum_{e}p(e\mid{\mathbf{b}})\,\underbrace{\sum_{y}p(y\mid{\mathbf{b}},e)}_{=1}=\sum_{e}p(e\mid{\mathbf{b}})=1.$
+Hence
+$1=\frac{1}{\lambda}\cdot 1\quad\Longrightarrow\quad\lambda=1.$
+Therefore,
+$q(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
+B.5 Decomposition and invariant
+Assume by causal representation learning, we have obtained $\mathbf{c}$ (stable latent variables) and $\mathbf{b}$ (unstable latent variables). Our goal is to estimate $p(y\mid\mathbf{c},\mathbf{b})$.
+$\frac{p(y=1\mid\mathbf{c},\mathbf{b})}{p(y=0\mid\mathbf{c},\mathbf{b})}$ (14)
+$=\frac{p(\mathbf{c},\mathbf{b},y=1)}{p(\mathbf{c},\mathbf{b},y=0)}$
+$=\frac{p(\mathbf{b}\mid y=1)p(\mathbf{c}\mid y=1)p(y=1)}{p(\mathbf{b}\mid y=0)p(\mathbf{c}\mid y=0)p(y=0)}$
+$=\frac{p(y=1\mid\mathbf{b})p(\mathbf{c}\mid y=1)}{p(y=0\mid\mathbf{b})p(\mathbf{c}\mid y=0)}$
+$=\frac{p(y=1\mid\mathbf{b})\frac{p(y=1\mid\mathbf{c})}{p(y=1)}}{p(y=0\mid\mathbf{b})\frac{p(y=0\mid\mathbf{c})}{p(y=0)}}$
+Assume we have $e_{all}=\{e_{0},e_{1},...,e_{n}\}$, representing different domains.
+We assume the source domains are $e_{s}=\{e_{0},e_{1},e_{s}\}$, and target domains are $e_{t}=\{e_{s+1},...,e_{n}\}$.
+$\forall e_{i},e_{j}\in e_{all}$ (15)
+$\quad\mathbf{c}\textbf{ only depends on }y$
+${p(\mathbf{c}\mid y=1,e=e_{i})}=p(\mathbf{c}\mid y=1,e=e_{j})$
+Use Bayes rule
+$\frac{p(y=1,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}=\frac{p(y=1,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
+$p(y=1\mid\mathbf{c})=\sum_{e\in e_{s}}{\frac{p(y=1,e=e\mid\mathbf{c})}{p(y=1,e=e)}p(e\mid e_{s})}$
+This result allows us to estimate $p(y=1\mid\mathbf{c})$ using data from source domains and apply it in the target domain.
+B.6 Proof of Equation 1
 The key observation is that because $\hat{y}$ only depends on $y$ (through $\mathbf{c}$}), we can relate the conditional distribution of the pseudo-label $\hat{y}$ given $\mathbf{b}$} to the distribution of the true label $y$.
@@
 Solving for $p(y=1\mid{\mathbf{b}})$ gives Equation 1.
-2.4.2 Using bias indirectly
-An alternative approach is to improve prediction by more accurately estimating the underlying environmental context.
-Let’s start from the objective, which can be formulated with the following optimization task:
-$q^{*}(y\!\mid\!{\mathbf{b}})=\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}{\textstyle\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\bigl{[}D_{\mathrm{KL}}(p(y\!\mid\!{\mathbf{b}},e)\,\|\,q(y\!\mid\!{\mathbf{b}}))\bigr{]}}.$ (5)
-This suggests identifying a predictor that minimizes the expected Kullback-Leibler (KL) divergence from $p(y\mid{\mathbf{b}},e)$ across all environments. The Bayes-optimal solution ohn2023optimal is
-$q^{*}(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)$
-This implies that we can leverage the identified bias to infer the environmental context and then develop a suitable bias-aware predictor $p(y\mid{\mathbf{b}})$ with a set of domain experts $p(y\mid{\mathbf{b}},e)$ to bridge the gaps between environments.
-The detailed derivation can be found in Appendix D.3.
-Figure 3: Overall framework of the BAG method.
-The framework consists of three main modules: representation learning, predictor training, and test-time adaptation.
-In the representation learning stage, we employ a VAE model to disentangle content and bias variables.
-For prediction, we construct an invariant predictor, a label prior, and a bias-aware predictor that reweights multiple domain experts using a domain estimator.
-These three predictor components work together for the final prediction.
-During test-time adaptation, the bias-aware predictor is further refined by invariant guidance, using pseudo-labels generated by the invariant predictor, enhancing its adaptation to new domains.
-3 Method
-In this section, we build on our theoretical findings in  Section 2 to propose the Bias-Aware Generalization (BAG, in short) model which uses data bias for generalization.
-The whole pipeline can be found in Appendix E.1 .
-The multi-class version can be found in Appendix E.2 .
-3.1 Overall framework
-Representation learning and disentanglement.
-Following Section 2.2, we first take the observed data ${\mathbf{x}}$ from the source domains as the inputs of the image encoder, which maps the image into the latent space, such as ${\mathbf{z}}=\text{Encoder}({\mathbf{x}})$.
-Then, we take ${\mathbf{z}}$ into a decode, $\text{Decoder}(\cdot)$, to generate the reconstruction as $\hat{{\mathbf{x}}}$.
-We apply the VAE loss to train both encoder and decoder, where the objectives can be written as
-$\mathcal{L}_{\text{vae}}\;=\;\mathbb{E}_{{\mathbf{x}}}\bigl{[}\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|^{2}\;+\;\beta\mathrm{KL}\bigl{(}q_{\phi}(\mathbf{z}\mid{\mathbf{x}})\,\|\,p(\mathbf{z})\bigr{)}\bigr{]},$ (6)
-Where $\beta$ is a hyper-parameter to balance the reconstruction loss and the KL divergence, which is set to be 1 in our approach.
-The prior distribution $p({\mathbf{z}})$ is chosen as an isotropic Gaussian.
-Minimizing equation 6 encourages $(\mathbf{c},\mathbf{b})$ to be identifiable in a block-wise sense and captures both invariant and domain-specific factors for further modeling.
-Following the method jiang2022invariant , we leverage a necessary but not sufficient condition for $\mathbf{c}\perp\!\!\!\perp\mathbf{b}\mid y$, that is $\mathbb{E}\bigl{[}\mathbf{c}\cdot(\mathbf{b}-\mathbb{E}[\mathbf{b}\mid y])\bigr{]}=0$.
-Specifically, we add a regular term to constrain the condition independence, which can be written as:
-$\mathcal{L}_{ind}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{c}_{i}\cdot\bigr{(}\mathbf{b}_{i}-\frac{1}{\bigl{|}\{j:y_{j}=y_{i}\}\bigr{|}}\sum_{j:y_{j}=y_{i}}\mathbf{b}_{j}\bigr{)}$
-Predictors
-Based on the learned content and bias variables, ${\mathbf{c}}$ and ${\mathbf{b}}$, we can make predictions from the latent space to the label target.
-Guided By Theorem 2.3, we implement the label predictor into three components:
-$f({\mathbf{z}})\;=\;f_{b}(\mathbf{b})\;+\;f_{c}(\mathbf{c})\;-\;Pr,$
-where $f_{b}$ is the bias-aware predictor (i.e.
-$p(y\mid{\mathbf{b}})$ ), whose input is the bias ${\mathbf{b}}$, $f_{c}$ is the invariant predictor to leverage the content variable ${\mathbf{c}}$.
-$Pr\in\mathbb{R}^{2}$ is a learnable bias term representing the logits prior $\operatorname{logit}(p(y))$.
-In particular, $f_{c}({\mathbf{c}})$ and $Pr$ together form an “invariant” module in all domains and can be used directly in the target domain without further modification.
-Since $\mathbf{b}$ depends on the domain $e$, using the $f_{s}$ only trained in the source domain naively at test time may introduce unwanted biases if the target domain differs significantly from the source domains.
-3.2 Bias-Aware predictors
-To exploit the bias $\mathbf{b}$ in the indirect manner, we introduce a set of $M$ learnable domain embeddings $\{e_{1},e_{2},\ldots,e_{M}\}$, which are corresponding to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
-Each embedding $e_{k}$ captures a distinct “mode” or style that may emerge in the source training data.
-Intuitively, this design follows the mixture-of-experts principle, allowing each expert to specialize in a subset of the training distribution.
-In terms of model design, each $e_{k}$ is a learnable vector (or a small neural component) trained to represent a particular domain bias.
-We implement a domain classifier $p(e=e_{i}\mid{\mathbf{b}})$ by a softmax layer over linear outputs, to produce the scalar weights These weights indicate how well each expert $e_{i}$ aligns with the domain-specific bias $\mathbf{b}$.
-For each expert, we parametrize $p\bigl{(}y\mid\mathbf{b},e=e_{i}\bigr{)}$ as an MLP network, whose input is the combination of ${\mathbf{b}}$ and $e_{i}$.
-It models the expert specializing in domain $e_{i}$, making specific predictions based on the latent representation ${\mathbf{b}}$.
-Formally, we calculate the bias-aware predictor as a re-weighting process of different domain-specific predictors as
-$f_{b}(\mathbf{b})\ =\sum_{i=0}^{M}p\bigl{(}y\mid\mathbf{b},e=e_{i}\bigr{)}p(e=e_{i}\mid{\mathbf{b}}).$
-In training, we jointly learn all bias-aware predictions, the domain classifier, the invariant predictor, and the logits prior with a classification loss $\mathcal{L}_{\text{cls}}$.
-3.3 Test time adaptation with invariance guidance
-During test time, when accessing test data, we can further refine the model by exploiting the invariant component $\mathbf{c}$ as guided by Theorem 2.5.
-Concretely, when we receive test-domain inputs $\{{\mathbf{x}}^{\mathcal{T}}\}$, we first extract their latent representations $[\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}]$ using the pre-trained image encoder.
-We then leverage the learned invariant predictor $f_{c}(\mathbf{c}^{\mathcal{T}})$ to generate pseudo-labels $\hat{y}^{\mathcal{T}}$ on unlabeled test data.
-In source domains, we can also have pseudo-labels, so we can collect $h_{0}$ and $h_{1}$ in source domains as shown in Theorem 2.5.
-It’s part of the probability confusion matrix between pseudo-labels and ground truth.
-Intuitively, these parameters adjust how the domain-invariant information can map the latent style factor $\mathbf{b}$ to valid probability estimates in the new domain.
-According to Equation 4, recovering the corrected prediction $p(y=1\mid\mathbf{b})$ require knowing $p(\hat{y}=1\mid\mathbf{b})$ and the calibration terms $h_{0},h_{1}$.
-To estimate $p(\hat{y}=1\mid\mathbf{b})$, we update the bias-aware predictor $f_{b}$ with the test data $\mathbf{b}^{\mathcal{T}}$ and the generated pseudo-labels $\hat{y}^{\mathcal{T}}$.
-This process can be viewed as a lightweight test-time adaptation step:
-$\min_{f_{b}}\;\mathbb{E}_{\mathbf{b}^{\mathcal{T}}}\Bigl{[}\mathcal{L}_{ada}=\ell\Bigl{(}\sigma\bigl{(}f_{b}(\mathbf{b}^{\mathcal{T}})\bigr{)},\;\hat{y}\Bigr{)}\Bigr{]},$ (7)
-where the adaption loss $\mathcal{L}_{ada}$ is characterized as the cross-entropy loss with pseudo labels.
-We record this post-trained predictor as $\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}})$.
-Let $\phi(\cdot)=\operatorname{logit}((\frac{\sigma(\cdot)+h_{0}-1}{h_{0}+h_{1}-1}))$ to represent the test-time correction.
-We denote the corrected bias-aware predictor as $q(y|b^{\mathcal{T}})$.
-The final prediction is
-$f([\mathbf{c}^{\mathcal{T}},\mathbf{b}^{\mathcal{T}}])=f(\mathbf{c}^{\mathcal{T}})+\phi(\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}}))-Pr.$
-Hence, the losses of the framework in both the training and test-time adaptation stages are as follows:
+B.7 Proof and Discussion of equation 4.2
+Setup and assumptions.
+Write
+$h_{0}=p(\hat{y}=0\mid y=0),\qquad h_{1}=p(\hat{y}=1\mid y=1),\qquad J=h_{0}+h_{1}-1\in(0,1].$
+Let $p(y=1\mid{\mathbf{b}})$ denote the target conditional and $q({\mathbf{b}})=p(\hat{y}=1\mid{\mathbf{b}})$.
+We assume (i) ${\mathbf{c}}\perp e\mid y$ and ${\mathbf{b}}\perp{\mathbf{c}}\mid y$, hence $\hat{y}=f_{c}({\mathbf{c}})$ implies $\hat{y}\perp{\mathbf{b}}\mid y$; (ii) class-conditional noise (CCN): $p(\hat{y}=1\mid y=1,{\mathbf{b}},e)=h_{1}$ and $p(\hat{y}=1\mid y=0,{\mathbf{b}},e)=1-h_{0}$ for all $({\mathbf{b}},e)$; (iii) a margin assumption $p(y=1\mid{\mathbf{b}})\in[\rho,1-\rho]$ almost surely, for some $\rho\in(0,1/2)$.
+Step 1: Identification under CCN.
+By $\hat{y}\perp{\mathbf{b}}\mid y$ and CCN,
+$q({\mathbf{b}})$ $=p(\hat{y}=1\mid{\mathbf{b}})=p(\hat{y}=1\mid y=1)\,p(y=1\mid{\mathbf{b}})+p(\hat{y}=1\mid y=0)\,p(y=0\mid{\mathbf{b}})$
+$=(1-h_{0})+J\,p(y=1\mid{\mathbf{b}}),$
+hence
+$p(y=1\mid{\mathbf{b}})=\frac{q({\mathbf{b}})+h_{0}-1}{J}.$ (16)
+Step 2: Linearization of the corrected estimator and $L^{2}$ bound.
+Define the corrected bias predictor
+$\widehat{f}({\mathbf{b}})=\frac{\widehat{q}({\mathbf{b}})+\hat{h}_{0}-1}{\hat{h}_{0}+\hat{h}_{1}-1},$
+where $\widehat{q}({\mathbf{b}})$ estimates $q({\mathbf{b}})$ and $(\hat{h}_{0},\hat{h}_{1})$ estimate $(h_{0},h_{1})$ from sources.
+Consider the map $F(q,h_{0},h_{1})=(q+h_{0}-1)/(h_{0}+h_{1}-1)$.
+A first-order expansion of $F$ at $(q,h_{0},h_{1})$ yields, pointwise in $\mathbf{b}$},
+$\widehat{f}({\mathbf{b}})-p(y=1\mid{\mathbf{b}})=\frac{\widehat{q}({\mathbf{b}})-q({\mathbf{b}})}{J}+\frac{1-p(y=1\mid{\mathbf{b}})}{J}\,(\hat{h}_{0}-h_{0})-\frac{p(y=1\mid{\mathbf{b}})}{J}\,(\hat{h}_{1}-h_{1})+\mathrm{h.o.t.}$ (17)
+(The identities $\partial F/\partial q=1/J$, $\partial F/\partial h_{0}=(1-p)/J$, and $\partial F/\partial h_{1}=-p/J$ follow from equation 16.) Squaring equation 17, using $0\leq p(y=1\mid{\mathbf{b}})\leq 1$, and taking expectation over ${\mathbf{b}}\sim p({\mathbf{b}})$, we obtain
+$\mathbb{E}\big[(\widehat{f}({\mathbf{b}})-p(y=1\mid{\mathbf{b}}))^{2}\big]\;\leq\;\frac{2}{J^{2}}\,\mathbb{E}\big[(\widehat{q}({\mathbf{b}})-q({\mathbf{b}}))^{2}\big]+\frac{C}{J^{2}}\,\Big((\hat{h}_{0}-h_{0})^{2}+(\hat{h}_{1}-h_{1})^{2}\Big)+o(1),$ (18)
+for an absolute constant $C>0$ (the $o(1)$ term collects higher-order products of estimation errors).
+Step 3: From $L^{2}$ error to KL risk.
+For any $p,q\in[\rho,1-\rho]$,
+$\mathrm{KL}\big(\mathrm{Bern}(p)\,\|\,\mathrm{Bern}(q)\big)\;\leq\;C_{\rho}\,(p-q)^{2},\qquad C_{\rho}=\frac{1}{\rho(1-\rho)}.$ (19)
+Therefore, with the risk $\mathcal{R}_{(b)}(g)=\mathbb{E}_{{\mathbf{b}}}[\mathrm{KL}(\mathrm{Bern}(p(y=1\mid{\mathbf{b}}))\,\|\,\mathrm{Bern}(g({\mathbf{b}})))]$ and $\mathcal{R}_{(b)}(p)=0$,
 […]
 […]
-where $\lambda_{0}$, $\lambda_{1}$ denote the hyper-parameters.
-4 Experiments
-In this section, we test our BAG method on both synthetic data and real-world datasets that require adaptation in the target domain.
-To make a fair comparison, we follow the setting in jiang2022invariant .
-Baseline methods
-We compare our method with ERM and IRM irm , a classic approach that leverages invariant information for making predictions.
-Additionally, we compare with ACTIR jiang2022invariant , which disentangles the invariant feature and bias by an independent regularization.
-And with SFB eastwood2024spuriosity , which aims to exploit spurious features but in a weaker setting.
-We briefly summarized the difference between SFB and our method in the last part of the related work.
-SFB can serve as a special case of our method when 1) the environment doesn’t affect the label, and 2) don’t bias aware for prediction.
-For Office-Home dataset, since SFB doesn’t have results on it, we add MMD and GMDG as more baselines to show our performance.
-4.1 Simulation experiments
-Data simulation
-To examine whether BAG can effectively identify and leverage domain shift information for better generalization, we devised a series of simulation experiments.
-Our data generation process strictly adheres to the causal structure depicted in Figure 2.
-First, we sample an environmental variable $e$ following a categorical distribution, $e\sim\mathrm{Categorical}(\{\pi_{1},\dots,\pi_{M}\})$, which subsequently induces variations across different domains.
-Then, a binary label $y$ is generated in an environment-dependent manner: $y=\mathbf{1}\{\,w^{\mathsf{T}}E+b_{0}>0\}.$ Next, we assign stable content $c$ based on the value of $y$: if $y=0$, then $c=c_{0}+\epsilon$; if $y=1$, then $c=c_{1}+\epsilon$.
-Here, $\epsilon\sim\mathcal{N}(0,\sigma_{c}^{2}I)$.
-Simultaneously, the bias variable $b$ is determined as: $b=E+\mathbf{C}_{e,y},$ ensuring that the background shift is influenced jointly by the environment and the label.
-This setup emulates the common real-world scenario where false associations often arise.
-Finally, stable content and bias jointly generate the observed data: $x=M\begin{bmatrix}c\\ b\end{bmatrix}+\eta,$ where $M$ is an invertible matrix, and $\eta\sim\mathcal{N}(0,\sigma_{x}^{2}I)$ represents measurement noise.
-Results and Discussions
-This study evaluates several methods on a synthetic dataset of 5000 samples.
-To ensure result reliability, each method was repeated five times, and the average performance was recorded.
-As shown in the right side of Table 1, both ERM and IRM performed well on the training set but achieved test accuracies just above 50%.
-We attribute this to the significant distributional disparity of environmental information $e$ across domains in the dataset, which limits the effectiveness of the class label $y$ in supporting accurate predictions.
-In other words, an excessive focus on invariant information may hinder model optimization, leading to marginal performance gains.
-In contrast, our method leverages domain-specific information (${\mathbf{b}}$) more effectively, significantly enhancing prediction performance by reducing noise compared to the class label $y$.
-Table 1: The comparison of the PACS dataset and the Synthetic dataset. All BAG results on PACS are obtained by averaging over 3 seeds. Baseline results are taken from eastwood2024spuriosity
-Algorithm PACS Synthetic
-P A C S Avg Acc
-ERM 93.0$\pm$0.7 79.3$\pm$0.5 74.3$\pm$0.7 65.4$\pm$1.5 78.0 54.62$\pm$3.70
-ERM + PL 93.7$\pm$0.4 79.6$\pm$1.5 74.1$\pm$1.2 63.1$\pm$3.1 77.6
-IRM (irm, ) 93.3$\pm$0.3 78.7$\pm$0.7 75.4$\pm$1.5 65.6$\pm$2.5 78.3 54.73$\pm$2.72
-IRM + PL 94.1$\pm$0.7 78.9$\pm$2.9 75.1$\pm$4.6 62.9$\pm$4.9 77.8
-ACTIR (jiang2022invariant, ) 94.8$\pm$0.1 82.5$\pm$0.4 76.6$\pm$0.6 62.1$\pm$1.3 79.0 70.84$\pm$6.37
-SFB eastwood2024spuriosity 95.8$\pm$0.6 80.4$\pm$1.3 76.6$\pm$0.6 71.8$\pm$2.0 81.2
-BAG 96.6$\pm$0.8 86.0$\pm$0.4 77.9$\pm$1.0 73.0$\pm$0.4 83.4 97.48$\pm$1.82
-4.2 Real-world experiments
-Experimental Settings
-We test our approach in commonly used PACS li2017deeper and Office-Homeoffice datasets.
-In the PACS dataset, we let one domain become the target and other domains as source domains.
-We use ResNet18 for PACS as the backbone and ResNet50 for Office-Home with MLP-based VAEs as the backbone and classifiers.
-We show our results over 3 random seeds.
-Our framework employs a single-layer and a two-layer linear layer as the encoder and decoder, respectively, and utilizes a single-layer linear layer for each specific classifier.
-We define three learnable parameters for the environment embedding.
-The experiment details can be found in Appendix F.
-Table 2: Results on Office-Home with ResNet-50. BAG results are averaged over 3 seeds.
-Alg.
-Office-Home Art Clipart Product Real Avg ERM 63.1 51.9 77.2 78.1 67.6 IRMirm 62.4 53.4 75.5 77.7 67.2 SFBeastwood2024spuriosity MMDli2018domain 62.4 53.6 75.8 76.4 67.1 GMDGtan2024rethinking 68.9 56.2 79.9 82.0 71.7 BAG 68.8 57.2 80.0 82.1 72.0
-Results and Discussions
-For the PACS data, in the left side of Table 1, we can find that our model is far superior to other models, with an average improvement of $2.2\%$ compared with best baseline, SFB method.
-Our model shows consistent improvement across all target domains.
-Among all target domains, we achieved better improvement on the Art Painting test, which may be because of the related bias in the domains.
-According to the experimental results of the Office-Home dataset on the Table 2, our BAG model outperforms all other baselines on three transfer tasks.
-It should be noted that we achieve much better performance in Clipart, which is the harder task; it shows BAG can achieve better generalization even when the domain gap is larger than the baselines.
-Considering that BAG involves two steps of optimization, we demonstrate the effects of each individual optimization strategy, namely BAG-BA, and BAG-TTA.
-Additionally, we explore the learning of latent variables by designing a model variant, BAG-VAE, which is a version of the model without the VAE component.
-We show that all the models performance better compared with baselines both in PACS and Synthetic data, but worse than BAG.
-The details of ablation study can be found in Appendix F.
-5 Conclusion
-In this paper, we challenge the conventional methods for OOD generation that treat bias solely as an obstacle.
-Instead, we theoretically demonstrate that bias can be strategically leveraged to enhance prediction, particularly when it retains useful dependencies on the target domain labels.
-Our framework demonstrates that the bias can be utilized in both direct and indirect ways.
-Through the empirical validation on synthetic and real-world benchmarks, we establish the effectiveness of this approach.
-The results consistently show that leveraging bias in a structured manner leads to more robust and adaptable models, offering a new perspective on OOD generation.
-Limitation: This work primarily aims to question the conventional treatment of bias and provide a theoretical foundation for understanding its effects.
-The experiments presented in this paper are intended to validate our proposed method, which don’t include large-scale experiments on extensive datasets.
-We acknowledge this limitation and leave systematic benchmarking on large datasets as an important direction for future work.
-Appendix for
-"Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation"
-Appendix A Related Work
-Domain generalization.
-Domain Generalization (DG) aims to develop models that generalize to unseen OOD target domains using only observed source domains for training.
-A common approach is learning invariant representations across all source domains, as seen in Invariant Risk Minimization (IRM) [1] and its extensions [16, 17, 18, 19, 20].
-Other methods leverage adversarial learning [21, 22, 23, 24, 25] to extract domain-invariant features or use disentangled representation learning [26, 27, 11] to separate domain-invariant from domain-specific information.
-Causal inference techniques [28, 4], such as back-door and front-door rules, help mitigate bias.
-Furthermore techniques such as data augmentation [29, 30, 31, 32], feature augmentation [33, 34, 35], and network regularization [36, 37] can reduce overfitting and improve generalization.
-Unsupervised domain adaptation Unlike DG, unsupervised domain adaptation utilizes unlabeled target data to guide model adaptation.
-Due to the observation of target distribution, many methods focus on distribution alignment by minimizing domain divergence [38, 39, 40, 41], using importance reweighting [42, 43, 44], and learning with adversarial discriminators, such as DANN [2], ADDA [45], and WDDA [46].
-Beyond distribution alignment, some methods focus on invariant information-based adaptation, including self-training [3, 47, 48, 49, 50, 51], which uses pseudo-labels from a source-trained model to guide target domain learning, and batch normalization-based regularization [52, 53], which adapts only BN layers while keeping other parameters invariant.
-Recently, disentangled representation learning [54, 55] has shown strong potential for adaptation by separating domain-invariant content from domain-specific bias.
-Test-Time domain adaptation In real-world scenarios, the target domain distribution is often unknown in advance, making it challenging to train an adaptive model beforehand, which motivates test time adaptation.
-The typical solution is Test-Time Training (TTT) [56, 57, 58, 59], which adapts models during inference by optimizing the same self-supervised learning objective used during pre-training.
-Besides, even without access to the target distribution, adaptation remains feasible with invariant information learned from the source domain, such as through self-training[60, 61, 62, 63] or network regularization [64, 65].
-The most closely related work to ours is Stable Feature Boosting (SFB), which leverages unstable features alongside invariant representations to generate pseudo-labels when only the image is affected by environmental factors.
-Our approach differs in three key ways.
-First, we delve deeper into latent variable modeling to provide a more comprehensive explanation of bias.
-Second, our model accommodates more complex bias scenarios, allowing the environment variable to influence both the image and the label.
-Third, we integrate invariant correction and expert reweighting to effectively utilize biased data.
-Appendix B Proof of identification and better prediction
-B.1 Proof of the identification of latent variables
-Lemma B.1.
-(Block-wise Identification of ${\mathbf{c}}$ and ${\mathbf{b}}$ [7].) Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
-  • A1 (Smooth and Positive Density): The probability density function of latent variables is smooth and positive, i.e. $P_{\mathbf{z}|\mathbf{e}}>0$.
-  • A2 (Conditional Independence): Conditioned on $\mathbf{e}$, each $z_{i}$ is independent of any other $z_{j}$ for $i.j\in\{1,\cdots,n\},i\neq j$, i.e., $\log P_{\mathbf{z}|\mathbf{e},\mathbf{y}}=\sum_{i=1}^{n}q(z_{i},\mathbf{e},\mathbf{y})$, where $q_{i}(z_{i},\mathbf{e},\mathbf{y})$ is the log density of the conditional distribution, i.e., $q_{i}:\log P_{z_{i}|\mathbf{e},\mathbf{y}}$.
-  • A3 (Linear Independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$, there exists $n_{b}+1$ values of $\mathbf{e}$, i.e., $e_{j}$ with $j=0,1,.\cdots,n_{b}$ such that these $n_{b}$ vectors $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})-\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{0})$ with $j=1,\cdots,n_{b}$ are linearly independent, where vector $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})$ is defined as follows:
-    $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}})=\left(\frac{\partial q_{1}({\mathbf{b}}_{1},{\mathbf{e}})}{\partial{\mathbf{b}}_{1}},\cdots\frac{\partial q_{n_{b}}({\mathbf{b}}_{n_{b}},{\mathbf{e}})}{\partial{\mathbf{b}}_{n_{b}}}\right),$ (8)
-  • A4 (Domain Variability:) There exist two values of $\mathbf{e}$, i.e., $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, s.t., for any set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability measure and cannot be expressed as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$, we have
+where $c_{1}=2C_{\rho}$ and $c_{2}=C_{\rho}C$ depend only on the margin parameter $\rho$. This is exactly the claimed upper bound equation 6. ∎
+Discussion.
 […]
-Then the learned $\hat{\mathbf{c}}$ and $\hat{\mathbf{b}}$ are block-wise identifiable.
-Proof.
-We begin with the matched marginal distribution $p_{\mathbf{x}|\mathbf{e},\mathbf{y}}$ to bridge the relation between $\mathbf{z}$ and $\hat{\mathbf{z}}$.
-Suppose that $\hat{g}:\mathbf{z}\rightarrow\mathcal{X}$ is an invertible estimated generating function, we have Equation (10).
-$\forall\mathbf{e}\in\mathcal{(}E),\quad p_{\hat{\mathbf{x}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{x}|\mathbf{e},\mathbf{y}}\Longleftrightarrow p_{\hat{g}(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{g(\mathbf{z})|\mathbf{e},\mathbf{y}}.$ (10)
-Sequentially, by using the change of variables formula, we can further obtain Equation (11)
-$p_{\hat{g}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}=p_{g(\mathbf{z}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{g^{-1}\circ g(\hat{\mathbf{z}})|\mathbf{e}}|\mathbf{J}_{g^{-1}}|=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{g^{-1}}|\Longleftrightarrow p_{h(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}},$ (11)
-where $h:=g^{-1}\circ g$ is the transformation between the ground-true and the estimated latent variables, respectively.
-$\mathbf{J}_{g^{-1}}$ denotes the absolute value of Jacobian matrix determinant of $g^{-1}$.
-Since we assume that $g$ and $\hat{g}$ are invertible, $|\mathbf{J}_{g^{-1}}|\neq 0$ and $h$ is also invertible.
-According to A2 (conditional independent assumption), we can have Equation (12).
-$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{z_{i}|\mathbf{e},\mathbf{y}}(z_{i}|\mathbf{e},\mathbf{y});\quad p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}(\hat{z}_{i}|\mathbf{e},\mathbf{y}).$ (12)
-For convenience, we take logarithm on both sides of Equation (12) and further let $q_{i}:=\log p_{z_{i}|\mathbf{e},\mathbf{y}},\hat{q}_{i}:=\log p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}$. Hence we have:
-$\log p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i},\mathbf{e},\mathbf{y});\quad\log p_{\hat{\mathbf{z},\mathbf{y}}|\mathbf{e}}=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}).$ (13)
-By combining Equation (13) and Equation (11), we have:
-$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}=p_{h(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{h^{-1}}|\Longleftrightarrow\sum_{i=1}^{n}q_{i}({z_{i},\mathbf{e}},\mathbf{y})+\log|\mathbf{J}_{h^{-1}}|=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}),$ (14)
-where $\mathbf{J}_{h^{-1}}$ are the Jacobian matrix of $h^{-1}$.
-Sequentially, we take the first-order derivative with $\hat{c}_{j}$ on Equation (14), where $j\in\{1,\cdots,n_{c}\}$, and have
-$\sum_{i=1}^{n}\frac{\partial q_{i}(z_{i},\mathbf{u},\mathbf{y})}{\partial z_{i}}\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}+\frac{\partial\log|\mathbf{J}_{h^{-1}}|}{\partial\hat{c}_{j}}=\frac{\partial q_{j}(\hat{c}_{j},\mathbf{u},\mathbf{y})}{\partial\hat{c}_{j}}.$ (15)
-Suppose $\mathbf{e}=e_{0},e_{1},\cdots,e_{n_{b}}$, we subtract the Equation (15) corresponding to $e_{k}$ with that corresponds to $e_{0}$, and we have:
-$\sum_{i=1}^{n}\left(\frac{\partial q_{i}(z_{i},e_{k},\mathbf{y})}{\partial z_{i}}-\frac{\partial q_{i}(z_{i},e_{0},\mathbf{y})}{\partial z_{i}}\right)\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}=\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}.$ (16)
-Since the distribution of estimated $\hat{\mathbf{c}}_{j}$ does not change across different domains, $\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}=0$.
-Since $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial z_{i}}$ does not change across different domains, $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial c_{i}}=\frac{\partial q_{i}(c_{i},e_{0},\mathbf{y})}{\partial c_{i}}$ for $i\in\{1,\cdots,n_{c}\}$.
-So we have
-$\sum_{i=1}^{n_{b}}\left(\frac{\partial q_{i}(b_{i},e_{k},\mathbf{y})}{\partial b_{i}}-\frac{\partial q_{i}(b_{i},e_{0},\mathbf{y})}{\partial b_{i}}\right)\cdot\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0.$ (17)
-Based on the linear independence assumption (A3), the linear system is a full-rank system.
-Therefore, the only solution is $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$.
-Since $h(\cdot)$ is smooth over $\mathbf{z}$, its Jacobian can be formalized as follows
-$\begin{gathered}\mathbf{J}_{h}=\begin{bmatrix}\begin{array}[]{c|c}\textbf{A}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{b}}}&\textbf{B}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{c}}}\\ \hline\cr\textbf{C}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{b}}}&\textbf{D}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{c}}}.\end{array}\end{bmatrix}\end{gathered}$ (18)
-Note that $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$ means that $\textbf{B}=0$.
-Since $h(\cdot)$ is invertible, $\mathbf{J}_{h}$ is a full-rank matrix.
-Therefore, for each $b_{i},i\in\{1,\cdots,n_{b}\}$, there exists a $h_{i}$ such that $b_{i}=h_{i}(\hat{\mathbf{b}})$.
-Besides, based on A4, one can show that all entries in the submatrix C zero according to part of the proof of Theorem 4.2 in [7](Steps 1, 2, and 3).
-Therefore, $\mathbf{b}$ and $\mathbf{c}$ are block-wise identifiable.
-∎
-B.2 Better predictor with adding bias
-Proof.
-Under Assumptions A1–A5, we compare the Bayes-optimal risks for predictors based on $c$ versus $(c,b)$.
-Log-loss.
-The Bayes-optimal log-loss risks satisfy
-$R_{\log}\bigl{(}q^{*}(\cdot\mid c)\bigr{)}=H(Y\mid C),\qquad R_{\log}\bigl{(}q^{*}(\cdot\mid c,b)\bigr{)}=H(Y\mid C,B),$
-where
-$H(Y\mid C)=-\mathbb{E}_{C}\bigl{[}\mathbb{E}_{Y\mid C}[\log p(Y\mid C)]\bigr{]},\quad H(Y\mid C,B)=-\mathbb{E}_{C,B}\bigl{[}\mathbb{E}_{Y\mid C,B}[\log p(Y\mid C,B)]\bigr{]}.$
-By A5 there is a positive-measure set of $(c,b)$ on which $p(y\mid c,b)\neq p(y\mid c)$. Hence
-$I(Y;B\mid C)=H(Y\mid C)-H(Y\mid C,B)=\mathbb{E}_{C,B}\!\bigl{[}D_{\mathrm{KL}}(p(y\mid c,b)\,\|\,p(y\mid c))\bigr{]}>0,$
-so $H(Y\mid C,B)<H(Y\mid C)$ and therefore
-$R_{\log}\bigl{(}q^{*}(\cdot\mid c,b)\bigr{)}<R_{\log}\bigl{(}q^{*}(\cdot\mid c)\bigr{)}.$
-Squared-error loss.
-The Bayes-optimal squared-error risks satisfy
-$R_{\mathrm{sq}}\bigl{(}f^{*}(c)\bigr{)}=\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C)\bigr{]},\qquad R_{\mathrm{sq}}\bigl{(}f^{*}(c,b)\bigr{)}=\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}.$
-By the law of total variance,
-$\mathrm{Var}(Y\mid C)=\mathbb{E}_{B}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}\;+\;\mathrm{Var}_{B}\!\bigl{(}\mathbb{E}[Y\mid C,B]\bigm{|}C\bigr{)}.$
-Since A5 ensures that $B$ influences the conditional mean of $Y$ given $C$, the second term is strictly positive on a set of positive probability. Thus
-$\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C,B)\bigr{]}<\mathbb{E}\bigl{[}\mathrm{Var}(Y\mid C)\bigr{]},$
-and consequently
-$R_{\mathrm{sq}}\bigl{(}f^{*}(c,b)\bigr{)}<R_{\mathrm{sq}}\bigl{(}f^{*}(c)\bigr{)}.$
-This completes the proof that incorporating $b$ strictly reduces both log-loss and squared-error Bayes risks. ∎
 Appendix C Multi-class Method
 Suppose we have $K\geq 2$ classes.
-We "one-hot encode" these classes, so that $y$ takes values in the set $\mathcal{Y}=\{(1,0,...,0),(0,1,0,...,0),...,(0,...,0,1)\}\subseteq\{0,1\}^{K}.$ In the training stage, let $\hat{y}=f_{\mathbf{c}}(\mathbf{c})$ $\epsilon\in[0,1]^{\mathcal{Y}\times\mathcal{Y}}$ with
+We ”one-hot encode” these classes, so that $y$ takes values in the set $\mathcal{Y}=\{(1,0,...,0),(0,1,0,...,0),...,(0,...,0,1)\}\subseteq\{0,1\}^{K}.$ In the training stage, let $\hat{y}=f_{\mathbf{c}}(\mathbf{c})$ $\epsilon\in[0,1]^{\mathcal{Y}\times\mathcal{Y}}$ with
 $\epsilon_{y_{i},y_{j}}=p[\hat{y}=y_{i}\mid y=y_{j}]$
@@
 $=\sum_{y_{i}\in\mathcal{Y}}E[\hat{y}\mid y=y_{i},\mathbf{b}]p[y=y_{i}\mid\mathbf{b}]$
-$\hat{y}\textbf{ only depends on }y$
+y^ only depends on $\hat{y}\textbf${ only depends on }y
 $=\sum_{y_{i}\in\mathcal{Y}}E[\hat{y}\mid y=y_{i}]p[y=y_{i}\mid\mathbf{b}]$
@@
 When $\epsilon$ is non-singular, this has the unique solution $E[y\mid\mathbf{b}]=\epsilon^{-1}E[\hat{y}\mid\mathbf{b}]$, giving a multiclass equivalent.
-However, it is numerically more stable to estimate $E[y\mid\mathbf{b}]$ by the least-squares solution
+However, it is numerically more stable to estimate $E[y\mid\mathbf{b}]$ by the least-squares solution.
 $\operatorname{argmin}_{p\in\Delta^{\mathcal{Y}}}\left\|\epsilon p-E[\hat{y}\mid\mathbf{b}]\right\|_{2}.$
-Appendix D Proof of section 2
-D.1 Decomposition and invariant
-Assume by causal representation learning, we have obtained $\mathbf{c}$ (stable latent variables) and $\mathbf{b}$ (unstable latent variables). Our goal is to estimate $p(y\mid\mathbf{c},\mathbf{b})$.
-$\frac{p(y=1\mid\mathbf{c},\mathbf{b})}{p(y=0\mid\mathbf{c},\mathbf{b})}$ (22)
-$=\frac{p(\mathbf{c},\mathbf{b},y=1)}{p(\mathbf{c},\mathbf{b},y=0)}$
-$=\frac{p(\mathbf{b}\mid y=1)p(\mathbf{c}\mid y=1)p(y=1)}{p(\mathbf{b}\mid y=0)p(\mathbf{c}\mid y=0)p(y=0)}$
-$=\frac{p(y=1\mid\mathbf{b})p(\mathbf{c}\mid y=1)}{p(y=0\mid\mathbf{b})p(\mathbf{c}\mid y=0)}$
-$=\frac{p(y=1\mid\mathbf{b})\frac{p(y=1\mid\mathbf{c})}{p(y=1)}}{p(y=0\mid\mathbf{b})\frac{p(y=0\mid\mathbf{c})}{p(y=0)}}$
-Assume we have $e_{all}=\{e_{0},e_{1},...,e_{n}\}$, representing different domains.
-We assume the source domains are $e_{s}=\{e_{0},e_{1},e_{s}\}$, and target domains are $e_{t}=\{e_{s+1},...,e_{n}\}$.
-$\forall e_{i},e_{j}\in e_{all}$ (23)
-[…]
-${p(\mathbf{c}\mid y=1,e=e_{i})}=p(\mathbf{c}\mid y=1,e=e_{j})$
-   Use Bayes rule
-$\frac{p(y=1,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}=\frac{p(y=1,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
-$p(y=1\mid\mathbf{c})=\sum_{e\in e_{s}}{\frac{p(y=1,e=e\mid\mathbf{c})}{p(y=1,e=e)}p(e\mid e_{s})}$
-This result allows us to estimate $p(y=1\mid\mathbf{c})$ using data from source domains and apply it in the target domain.
-D.2 Proof of theorem 2.5
-In this section, we would like to prove $h_{0}+h_{1}=1$ if and only if $\mathbf{c}\perp\!\!\!\perp y$.
-Forward Implication: Suppose $h_{0}+h_{1}=1$.
-$h_{0}+h_{1}=1\quad\text{if and only if}\quad\mathbf{c}\perp y.$
-Note that the entire result also holds, with almost identical proof, in the multi-environment setting of Sections 3 and 5, conditioned on a particular environment $E$.
-Proof. We first prove the forward implication.
-Suppose $h_{0}+h_{1}=1$.
-If $\Pr[y=1]\in\{0,1\}$, then $\mathbf{c}$ and $y$ are trivially independent, because the value of $y$ is constant. In this case, the result holds trivially. So, we assume $\Pr[y=1]\in(0,1)$.
-We now compute the expected value of the pseudo-label $\hat{y}$ using the Law of Total Expectation:
-$E[\hat{y}]=\Pr[\hat{y}=1\mid y=1]\cdot\Pr[y=1]+\Pr[\hat{y}=0\mid y=0]\cdot\Pr[y=0].$
-That is,
-$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p),$
-where $p=\Pr[y=1]$.
-Next, we simplify this expression:
-$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p)=(h_{0}+h_{1}-1)\cdot p+1-h_{0}.$
-Using the assumption $h_{0}+h_{1}=1$, we get:
-$E[\hat{y}]=1-h_{0}.$
-Thus, $E[\hat{y}]=1-h_{0}$.
-Now, since the expected value of $\hat{y}$ is independent of the actual value of $y$, this suggests that $\hat{y}$ provides no information about $y$, implying that $\hat{y}$ and $y$ are independent.
-Since $\hat{y}$ is a deterministic function of $\mathbf{c}$, this implies that $\mathbf{c}$ and $y$ must also be independent.
-Therefore, we have shown that if $h_{0}+h_{1}=1$, then $\mathbf{c}\perp y$.
-To prove the reverse implication:
-Suppose $\mathbf{c}\perp y$.
-If $\mathbf{c}$ and $y$ are independent, then the pseudo-label $\hat{y}$ is also independent of $y$. Hence, we have:
-$\Pr[\hat{y}=1\mid y=1]=\Pr[\hat{y}=1\mid y=0].$
-This implies that:
-$h_{1}=1-h_{0}.$
-Therefore, we have:
-$h_{0}+h_{1}=h_{0}+(1-h_{0})=1.$
-Thus, we have shown that if $\mathbf{c}\perp y$, then $h_{0}+h_{1}=1$.
-Putting the forward and reverse implications together, we conclude that:
-$h_{0}+h_{1}=1\quad\Longleftrightarrow\quad\mathbf{c}\perp y.$
-This completes the proof.
-D.3 The derivation of the KL optimization
-We want to solve the following problem:
-$q^{*}(y\mid{\mathbf{b}})\;=\;\underset{q(\cdot\mid{\mathbf{b}})}{\arg\min}\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}\Bigr{]}.$
-Let
-$\mathcal{L}[q(\cdot\mid{\mathbf{b}})]\;=\;\mathbb{E}_{e\sim p(e\mid{\mathbf{b}})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},E)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}\Bigr{]}.$
-By the definition of KL divergence, for discrete $y$:
-$D_{\mathrm{KL}}\bigl{(}P(y)\,\|\,Q(y)\bigr{)}=\sum_{y}P(y)\,\log\frac{P(y)}{Q(y)}.$
-Hence, for each $e$,
-$D_{\mathrm{KL}}\bigl{(}p(y\mid{\mathbf{b}},e)\,\|\,q(y\mid{\mathbf{b}})\bigr{)}=\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
-Therefore,
-$\mathcal{L}[q]=\sum_{e}p(e\mid{\mathbf{b}})\,\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
-We want to choose $q(y\mid{\mathbf{b}})$ to minimize $\mathcal{L}[q]$, which needs $q(y\mid{\mathbf{b}})$ to satisfy $\sum_{y}q(y\mid{\mathbf{b}})\;=\;1,\text{for each fixed }{\mathbf{b}}.$ To incorporate this constraint, we introduce a Lagrange multiplier $\lambda$ and consider the Lagrangian:
-$\mathcal{J}[q,\lambda]=\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e)\,\log\!\biggl{[}\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\biggr{]}\;+\;\lambda\Bigl{(}\sum_{y}q(y\mid{\mathbf{b}})-1\Bigr{)}.$
-To find the minimizing $q$, we set
-[…]
-First, note
-$\log\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}\;=\;\log p(y\mid{\mathbf{b}},e)-\log q(y\mid{\mathbf{b}}).$
-Only the term $-\log q(y\mid{\mathbf{b}})$ depends on $q$. For a specific $y$,
-$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\bigl{[}-\log q(y^{\prime}\mid{\mathbf{b}})\bigr{]}=\begin{cases}-\frac{1}{q(y\mid{\mathbf{b}})},&\text{if }y^{\prime}=y,\\ 0,&\text{if }y^{\prime}\neq y.\end{cases}$
-Hence, for each $e$:
-$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\biggl{[}\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}\biggr{]}=-\frac{p(y\mid{\mathbf{b}},e)}{q(y\mid{\mathbf{b}})}.$
-Summing over $e$,
-$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y^{\prime}}p(y^{\prime}\mid{\mathbf{b}},e)\,\log\frac{p(y^{\prime}\mid{\mathbf{b}},e)}{q(y^{\prime}\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
-Also,
-$\frac{\partial}{\partial q(y\mid{\mathbf{b}})}\Bigl{[}\lambda\bigl{(}\sum_{y^{\prime}}q(y^{\prime}\mid{\mathbf{b}})-1\bigr{)}\Bigr{]}=\lambda.$
-Hence,
-$\frac{\partial\mathcal{J}}{\partial q(y\mid{\mathbf{b}})}=-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda.$
-Setting this to zero:
-$-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda=0,$
-which gives
-$\frac{1}{q(y\mid{\mathbf{b}})}\,\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\lambda.$
-Thus
-$q(y\mid{\mathbf{b}})=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
-We require $\sum_{y}q(y\mid{\mathbf{b}})=1$. Then:
-$1=\sum_{y}q(y\mid{\mathbf{b}})=\sum_{y}\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)=\frac{1}{\lambda}\sum_{e}p(e\mid{\mathbf{b}})\sum_{y}p(y\mid{\mathbf{b}},e).$
-Since $\sum_{y}p(y\mid{\mathbf{b}},e)=1$ for each $e$, we get
-$\sum_{e}p(e\mid{\mathbf{b}})\,\underbrace{\sum_{y}p(y\mid{\mathbf{b}},e)}_{=1}=\sum_{e}p(e\mid{\mathbf{b}})=1.$
-Hence
-$11=\frac{1}{\lambda}\cdot 1\quad\Longrightarrow\quad\lambda=$ 1 .
-Therefore,
-$q(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
-Appendix E Algorithms
-E.1 Binary Bias-Aware Generalization
+Appendix D Algorithms
+D.1 Binary Bias-Aware Generalization
 Algorithm 1 Bias-Aware Generalization
 1:Input: Source data $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})_{i=1}^{M}$; hyper-parameters $\lambda_{0},\lambda_{1}$,
-2:  : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding, $Pr$ means logits prior, $\operatorname{logit}(p(y))$.
+2: : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding, PrPr means logits prior, $\operatorname{logit}(p(y))$.
 3:Stage 1: Source Training
-4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
+4: Obtain $\mathbf{c},\mathbf{b}]$ via feeding $\mathbf{x}$} into encoder.
 5: Reconstruct $\hat{{\mathbf{x}}}$.
 6: Compute VAE loss $\mathcal{L}_{\text{vae}}$.
-7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
+7: Calculate the results of the invariant predictor $f_{c}{({\mathbf{c}})}$ .
 8: Calculate the bias predictor, obtain the final prediction.
@@
 15: Generate pseudo-label as $\hat{y}^{\mathcal{T}}=f_{c}(\mathbf{c}^{\mathcal{T}})$.
-16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
+[…]
 17: Correct the the results of bias predictor as $\phi(\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}}))=\operatorname{logit}((\frac{\sigma((\tilde{f}_{b}(\mathbf{b}^{\mathcal{T}})))+h_{0}-1}{h_{0}+h_{1}-1}))$ ,
 18: Make the final prediction.
-E.2 Multi-Class Bias-Aware Generalization
+D.2 Multi-Class Bias-Aware Generalization
 Algorithm 2 Multi-Class Bias-Aware Generalization
@@
 3:Stage 1: Source Training
-4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
+4: Obtain $\mathbf{c},\mathbf{b}]$ via feeding $\mathbf{x}$} into encoder.
 5: Reconstruct $\hat{{\mathbf{x}}}$.
 6: Compute VAE loss $\mathcal{L}_{\text{vae}}$.
-7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
+7: Calculate the results of the invariant predictor $f_{c}{({\mathbf{c}})}$.
 8: Calculate the bias predictor, obtain the final prediction as equation 21.
@@
 15: Generate pseudo-label as $\hat{y}^{\mathcal{T}}=f_{c}(\mathbf{c}^{\mathcal{T}})$.
-16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
+[…]
 17: Correct the the results of bias predictor by least squares using $\epsilon$ according Appendix C
 18: Make the final prediction as equation 21.
-Appendix F Experiment details and ablation Study
-F.1 Experiment details
+Appendix E Experiment details and ablation Study
+E.1 Experiment details
 PACS
@@
 We train with the Adam optimizer for 70 epochs and select the final model based on its performance under test-time augmentation on the held-out domain.
-We used the code based on [11].
+We used the code based on (Jiang and Veitch, 2022).
 Office-Home
-We evaluate our BAG model on the Office-Home dataset [13], which comprises four domains (Art, Clipart, Product, Real-World) and 65 object categories.
+We evaluate our BAG model on the Office-Home dataset (Venkateswara et al., 2017), which comprises four domains (Art, Clipart, Product, Real-World) and 65 object categories.
 Following the leave-one-domain-out protocol, we train on three domains and test on the held-out one.
-Hyperparameters for each held-out domain are optimized via Bayesian optimization; the selected values for learning rate, regularization weight, number of epochs for the $z_{s}$ classifier, and VAE weight are given in Table F.1.
+Hyperparameters for each held-out domain are optimized via Bayesian optimization; the selected values for learning rate, regularization weight, number of epochs for the $z_{s}$ classifier, and VAE weight are given in Table 3.
 The full code are provided in the supplementary material.
-Table F.1: Optimal hyperparameters on Office-Home
-Domain Learning Rate $\lambda_{1}$ epochs (test time) $\lambda_{0}$
+Table 3: Optimal hyperparameters on Office-Home
+Domain Learning Rate $\lambda_{1}$ # epochs (test time) $\lambda_{0}$
 Art (Ar) $5.7\times 10^{-5}$ 0.030 3 $1.2\times 10^{-4}$
@@
 Real-World (Rw) $5.0\times 10^{-5}$ 0.046 6 $2.5\times 10^{-5}$
-F.2 Ablation Study
-Table F.2 presents the performance of various BAG variants on the PACS dataset and in simulation experiments.
-All variants outperform ERM, demonstrating the effectiveness of our two bias-utilization methods.
-Notably, BAG-RE and BAG-TTA each exhibit their own strengths and weaknesses, which may be due to data-specific effects.
-BAG-VAE performs somewhat worse than BAG, underscoring the importance of accurately identifying and disentangling $\mathbf{b}$ and $\mathbf{c}$.
-Table F.2: Accuracy of different variants including BAG, BAG-VAE, BAG-RE, and BAG-TTA on four PACS domains and Synthetic data.
+Computing resources
+We evaluate the BAG method on DomainNet under the leave-one-domain-out protocol, using clipart as the target domain and infograph, painting, quickdraw, real, and sketch as source domains, trained for 30 epochs with a fixed random seed and batch size 64.
+The experiment is conducted on a single AMD Instinct MI210 GPU (ROCm).
+Peak GPU memory usage reported by PyTorch is 26.43 GB, and the total wall-clock time is 21,547.08 s (approximately 5 h 59 m 7 s).
+For SFB under the same protocol and settings (clipart as the target domain, 30 epochs, batch size 64, fixed seed, single AMD Instinct MI210 GPU), the training time is 21,583.46 s (approximately 5 h 59 m 43 s) and the total time is 21,584.40 s (approximately 5 h 59 m 44 s), with a peak PyTorch-allocated memory of 26.35 GB.
+Compared with the most relevant baseline, BAG achieves lower runtime under similar memory usage; the peak memory is dominated by the ResNet-50 backbone and its training-time activations and optimizer states, which are shared across both methods, so the overall memory footprints are nearly identical.
+E.2 Ablation Study
+Table 4 presents a systematic ablation of BAG on PACS and in our synthetic simulations.
+The full method comprises two optimization stages and a VAE module.
+In the ablations, we remove exactly one component at a time, as indicated by a hyphen in the variant name: BAG-VAE drops the VAE in favor of a linear projection; BAG–TTA removes the Test Time Adaptation that uses pseudo labels to guide the bias branch; and BAG–RE removes the Reweight Experents mechanism.
+For transparency on the predictive roles of the two branches, Table 5 further reports the behavior of the bias predictor and the content predictor when used in isolation.
+Empirically, all BAG variants outperform ERM across datasets, confirming the effectiveness of our bias–utilization design.
+BAG–RE and BAG–TTA show complementary strengths and weaknesses—consistent with data–specific factors such as domain composition and the severity of distribution shift—whereas BAG–VAE lags behind the full BAG, underscoring the value of accurately identifying and disentangling $\mathbf{b}$ (bias–aligned factors) and $\mathbf{c}$ (content–aligned factors).
+Using only the bias predictor or only the content predictor yields average accuracies of roughly $78\%$ and $80\%$, respectively; each alone lacks sufficient information and, without a learnable class prior, is vulnerable to label shift.
+Once the two predictors are combined under a learnable label prior, performance rises to state–of–the–art levels among the baselines considered.
+In summary, the ablations show that each component contributes meaningfully: (i) explicit disentanglement of bias and content, (ii) expert reweighting and test–time adaptation to stabilize routing, and (iii) a compact generative prior via the VAE.
+Together, these elements enable BAG to deliver robust and consistently strong domain–generalization performance, while simplified variants remain competitive and illuminate the role of each design choice.
+Table 4: Accuracy of different variants including BAG, BAG-VAE, BAG-RE, and BAG-TTA on four PACS domains and Synthetic data.
 Algorithm PACS Synthetic
@@
 BAG-TTA 96.3 80.0 78.8 73.0 82.0 85.5
-Appendix G Statement
-This work contributes to the advancement of machine learning by challenging conventional out-of-distribution (OOD) approaches and introducing a framework that strategically leverages bias rather than eliminating it.
+Table 5: Different parts performance results on PACS. Values are accuracies (%).
+Method P A C S Avg
+Content predictor Only 95.57 81.10 75.00 67.83 79.88
+Bias predictor Only 94.01 73.58 78.67 66.23 78.12
+BAG 96.29 83.69 80.29 76.05 84.08
+E.3 Sensitivity of experts and embeddings
+Table 6: Sensitivity with fixed experts E=3E{=}3 and varying embedding dimension $d$. Column-wise best in bold. Values are accuracies (%).
+DsDs $P$ AA $C$ $S$ average
+5 95.81 84.03 78.71 70.73 82.32
+10 96.29 83.69 80.29 76.05 84.08
+15 96.53 84.08 69.24 71.54 80.35
+20 96.23 83.30 75.85 76.00 82.85
+Table 7: Sensitivity with fixed embedding dimension d=10d{=}10 and varying number of experts $E$. Column-wise best in bold. Values are accuracies (%).
+$E$ $P$ AA $C$ $S$ average
+1 95.51 83.11 75.00 74.65 82.07
+3 96.29 83.69 80.29 76.05 84.08
+4 95.27 83.59 77.52 68.92 81.33
+5 95.51 84.28 75.13 73.15 82.02
+6 96.23 85.06 73.98 73.91 82.29
+7 95.45 84.08 74.62 73.63 81.94
+10 96.35 83.98 78.92 73.63 83.22
+Is the environmental predictor necessary, or could one directly predict the label with similar performance?
+Yes, the environmental predictor is necessary.
+Directly predicting $y$ from bias–aligned features $b$ effectively fits $p(y\mid b)$, which is brittle under label shift: when $p(y)$ changes across domains, the conditional $p(y\mid b)$ drifts, and target performance degrades.
+Our approach explicitly estimates a latent environment and routes each example to domain–conditioned experts, separating environment–specific variation from task–relevant signals and improving robustness.
+As shown in Table 4, removing the environmental predictor (method BAG–RE) reduces the mean accuracy to $81.8\%$, below that of the full model.
+In our hyperparameter sweeps (Tables 6 and 7), most configurations that include the predictor exceed $81.8\%$ on average, and the best setting with E=3,d=10E{=}3,d{=}10 reaches $84.08\%$ with strong performance on the challenging domains $C{=}80.29\%$ and $S{=}76.05\%$.
+In summary, conditioning the classifier on an explicitly predicted environment addresses distribution shift and yields measurable, domain–wise robust gains over direct prediction.
+How sensitive is performance to the number of experts and the embedding dimensionality?
+Sensitivity is non–monotonic in both factors.
+With experts fixed at E=3E{=}3 (Table 6), a medium environment embedding performs best: d=10d{=}10 attains the highest average $84.08\%$, whereas a smaller embedding underfits ($d{=}5:\ 82.32\%$).
+With embedding fixed at d=10d{=}10 (Table 7), increasing experts from E=1E{=}1 to E=3E{=}3 improves the average from $82.07\%$ to $\textbf{84.08}\%$; further increases provide no consistent gains (e.g., $E{=}10:\ 83.22\%$).
+A few well–specialized experts ($E{\approx}3$) together with a compact environment embedding ($d{\approx}10$) strike a reliable balance between expressivity and out–of–domain generalization.
+To summarize: BAG is not sensitive to these two parameters, and the performance is within a controllable range; even suboptimal parameter choices do not push it below baselines such as ERM, IRM and ACTIR.
+E.4 Disentanglement of Latent Variables
+We qualitatively examine the learned disentanglement on PACS by visualizing attributions for the content branch $\mathbf{c}$} and the bias branch $\mathbf{b}$} with Grad-CAM.
+Concretely, we reuse the trained checkpoint and the identical backbone as training.
+Given an input image $x$, we route it through the original BAG model but expose two thin wrappers that return the logits of the $\mathbf{c}$} head ($f_{\beta}$) and the $\mathbf{b}$} head ($f_{\eta}$), respectively.
+For Grad-CAM, we target the last residual block of layer4 in the backbone.
+We then compute two grayscale attribution maps: one for $\mathbf{c}$} by backpropagating through $f_{\beta}$, and one for $\mathbf{b}$} by backpropagating through $f_{\eta}$, each with ClassifierOutputTarget set to the ground-truth class.
+To make the overlays visually comparable across images, we (i) bilinearly resize each Grad-CAM map to the input resolution; (ii) min–max normalize it to $[0,1]$; and (iii) render red intensity for $\mathbf{c}$} and blue intensity for $\mathbf{b}$} before alpha-blending with the RGB image (default $\alpha=0.5$).
+For each sample we compose a four-tile horizontal panel in the fixed order: original image, $\mathbf{b}$} (blue), $\mathbf{c}$} (red), and the composite overlay; the titles b/c are boldfaced to reduce ambiguity.
+Across the shown examples, $\mathbf{c}$} consistently concentrates on class-relevant object regions, whereas $\mathbf{b}$} emphasizes background textures, styles, or peripheral cues.
+These observations align with the intended decomposition and support the use of $\mathbf{c}$} as content and $\mathbf{b}$} as environment/style bias in subsequent modules.
+[…]
+Appendix F Statement
+This work contributes to the advancement of machine learning by challenging conventional out‐of‐distribution (OOD) approaches and introducing a framework that strategically leverages bias rather than eliminating it.
 On the positive side, our methods can improve model robustness when deployed in naturally biased environments—e.g., medical imaging systems that must generalize across hospitals with different demographic distributions.
-By revealing how bias can be harnessed as a constructive signal, we hope to inspire new lines of research into principled bias-aware generalization across diverse domains.
+By revealing how bias can be harnessed as a constructive signal, we hope to inspire new lines of research into principled bias‐aware generalization across diverse domains.
 Because this is foundational research evaluated on controlled benchmarks and synthetic datasets, we do not identify any immediate negative societal impacts.